### v2

# Sharp Bounds for Multiple Models in Matrix Completion

Abstract
In this paper, we demonstrate how a class of advanced matrix concentration inequalities, introduced in Brailovskaya and van Handel [2024], can be used to eliminate the dimensional factor in the convergence rate of matrix completion. This dimensional factor represents a significant gap between the upper bound and the minimax lower bound, especially in high dimension. Through a more precise spectral norm analysis, we remove the dimensional factors for five different estimators in various settings, thereby establishing their minimax rate optimality.
1 Introduction
Matrix recovery from a small subset of entries, often referred to as matrix completion, is a classical problem in high-dimensional statistics with wide-ranging applications across various fields. Numerous frameworks and methods have been developed for matrix completion, with a fundamental understanding that a large matrix cannot be recovered unless it exhibits certain low-rank properties. As a convex relaxation of rank penalization, nuclear norm penalized trace regression has become a popular approach and corresponding theoretical research has been fruitful. See, for example, Candès and Tao [2010], Recht [2011], Koltchinskii et al. [2011], Negahban and Wainwright [2012], Klopp [2014], Klopp et al. [2017], Minsker [2018], Yu et al. [2024].
In particular, many studies focus on the convergence rate of the nuclear norm penalization methods [Koltchinskii et al., 2011, Negahban and Wainwright, 2012, Klopp, 2014, Yu et al., 2024]. However, the minimax lower bound, obtained in Koltchinskii et al. [2011] and Negahban and Wainwright [2012], does not include a logarithmic dimension factor that always appears in the upper bound of the convergence rate. This discrepancy between the two bounds becomes more pronounced in high-dimensional settings, highlighting a potential gap in the existing theoretical understanding. Due to this issue, previous works have had to qualify their results by stating “our estimator is minimax optimal, up to a logarithmic factor”. In this paper, we use a class of sharp matrix concentration inequalities to remove the logarithmic dimension factor. After our work, the qualification will no longer be necessary.
Problem setup
Let $A_{0}\in\mathbb{R}^{m_{1}\times m_{2}}$ be the unknown matrix we aim to recover. We observe independent samples $\{(X_{i},Y_{i})\}_{i=1}^{\nu}$ from the model
$Y_{i}=\langle X_{i},A_{0}\rangle+\xi_{i},\quad i=1,\ldots,\nu,$ (1)
where $X_{i}\in\mathbb{R}^{m_{1}\times m_{2}}$, $\langle X_{i},A_{0}\rangle=\text{tr}(X_{i}^{T}A_{0})$, and $\text{tr}(B)$ denotes the trace of matrix $B$. The noise variables $\xi_{i}$ are independent and identically distributed. The sampling matrices $X_{i}\in\mathbb{R}^{m_{1}\times m_{2}}$ take values from $\{e_{l}(m_{1})e_{k}^{T}(m_{2}),1\leq l\leq m_{1},1\leq k\leq m_{2}\}$, where $e_{j}(m_{1}),e_{k}(m_{2})$ denote standard basis vectors in $\mathbb{R}^{m_{1}},\mathbb{R}^{m_{2}}$, respectively. This implies that each $X_{i}$ only has one non-zero entry, indicating the location of the sampled entry.
The distribution of the noise variables $\xi_{i}$ will be specified in different settings. The sampling distribution of $X_{i}$ is denoted by
$P_{jk}=P(X_{i}=e_{j}(m_{1})e_{k}^{T}(m_{2})).$
Additionally, $X_{i}$ and $\xi_{i}$ are assumed to be independent. In this paper, we consider that $\text{rank}(A_{0})\leq r$ or $A_{0}$ is the sum of a low rank matrix and a sparse matrix, and $\|A_{0}\|_{\infty}\leq a$ where $\|A\|_{\infty}=\max_{1\leq i\leq m_{1},1\leq j\leq m_{2}}|a_{ij}|$ is the entrywise maximum norm of $A=(a_{ij})$. The total sample size $\nu=n{\rm\leavevmode\nobreak\ or\leavevmode\nobreak\ }N$ depending on the settings.
The logarithmic dimension factor issue
The matrix completion problem is often addressed by a penalized regression. Assuming $A_{0}$ is of low rank, a natural choice for the penalty is the rank of a matrix. However, the rank penalty is non-convex and makes the corresponding program difficult to solve. The nuclear norm $\|\cdot\|_{1}$, namely the sum of all the singular values of a matrix, is usually used as a convex relaxation for the rank. There are many classical works on nuclear norm penalization methods under different assumptions. Assuming the noise has exponential decay, Koltchinskii et al. [2011] consider the case that the sampling distribution is known; Negahban and Wainwright [2012] study the situation where the sampling distribution is unknown and the row indices and the column indices of the observations are sampled independently; Klopp [2014] explores the problem for general sampling distributions. Consider the classical nuclear norm penalized least squares estimator
$\hat{A}=\operatorname*{arg\,min}_{\|A\|_{\infty}\leq a}\left\{\dfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}+\lambda\|A\|_{1}\right\}.$
Assuming $\text{rank}(A_{0})\leq r$, $X_{i}$ has a general non-uniform distribution and the noise has exponential decay, Klopp [2014] obtains the following bound that holds with high probability:
$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\apprle\log(m_{1}+m_{2})\dfrac{r\max(m_{1},m_{2})}{n},$
[…]
$\inf_{\hat{A}}\sup_{\mathcal{A}}P\left(\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\apprge\dfrac{r\max(m_{1},m_{2})}{n}\right)\geq C,$
where $C\in(0,1)$ is a constant, and the infimum and supremum are taken over all the possible estimators and the parameter space $\{A_{0}\in\mathbb{R}^{m_{1}\times m_{2}}:\text{rank}(A_{0})\leq r,\|A_{0}\|_{\infty}\leq a\}$ respectively. The logarithmic dimension factor $\log(m_{1}+m_{2})$, as the gap between the upper and lower bounds, exists in other matrix completion settings as well. See, for instance, Klopp et al. [2017] for corrupted matrix completion, Yu et al. [2024] for matrix completion with heavy tailed noise.
Our contribution
Recently, Bandeira et al. [2023] and Brailovskaya and van Handel [2024] introduce new matrix concentration inequalities, which establish sharper bounds for certain important random matrices. We showcase how the new bounds can be leveraged in the matrix completion problem to drop the logarithmic dimension factor. Specifically, We consider five estimators across three different settings as follows:
  • First setting: the unknown matrix is of low rank and the noise is sub-Gaussian. This is the most classical setting. Two estimators are studied in Klopp [2014], one for the case of known variance and another for the case of unknown variance.
  • Second setting: the unknown matrix is of low rank and the noise is heavy tailed. The estimator in Yu et al. [2024] is designed for the case that the noise is only assumed to have finite second moment.
  • Third setting: the unknown matrix is corrupted and the noise is sub-Gaussian. Two estimators are studied under different corruption models in Klopp et al. [2017]: entrywise corruption and columnwise corruption.
All the previous results on the convergence rates of the above five estimators involve the logarithmic dimension factor. We remove the dimensional factor for all the five estimators, thereby proving their minimax rate optimality.
The remainder of the paper is organized as follows. Continuing in this section, we introduce the notation. In Section 2, we present our main results. We begin by discussing a classical matrix completion model and demonstrating how to eliminate the dimension factor. We then introduce additional examples and results. We give some concluding remarks in Section 3. Section 4 is dedicated to the proofs.
Notation
To present our results, we summarize the notations used throughout this paper.
  • $M=m_{1}\vee m_{2}$, $m=m_{1}\wedge m_{2}$, $d=m_{1}+m_{2}$.
  • Let $A=(a_{ij})$, for $1\leq i\leq m_{1},1\leq j\leq m_{2}$. Let $\{\sigma_{l}\}_{l=1}^{m}$ denote ordered singular values of $A$: $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{m}$. $\|A\|_{F}=\sqrt{\sum_{i=1}^{m_{1}}\sum_{j=1}^{m_{2}}a_{ij}^{2}}$, $\|A\|_{1}=\sum_{l=1}^{m}\sigma_{l}$, $\|A\|=\sigma_{1}$, $\|A\|_{\infty}=\max_{i\leq m_{1},j\leq m_{2}}|a_{ij}|$, $\|A\|_{1,*}=\sum_{i,j}|a_{ij}|$. With $A^{k}$ denoting the $k^{th}$ column of A, define $\|A\|_{2,1}=\sum_{k=1}^{m_{2}}\|A^{k}\|_{2}$, where $\|\cdot\|_{2}$ is the usual $l_{2}$-norm. $\|A\|_{w(F)}=\sqrt{\sum_{i=1}^{m_{1}}\sum_{i=1}^{m_{2}}a^{2}_{ij}P_{ij}}$ is a weighted Frobenius norm associated with the distribution of the sampling matrices $X_{i}$.
  • $C,C_{1},C_{2},\ldots$ are positive constants whose value may vary at each occurrence.
  • For a matrix $A$, denote its singular value decomposition as $A=U\Sigma V^{T}$. We define the projection operators
    $P_{A}(B)=UU^{T}BVV^{T}$
    and
    $P_{A}^{\perp}(B)=B-P_{A}(B).$
  • $\{\epsilon_{i}\}_{i=1}^{n}$ are i.i.d. symmetric Bernoulli random variables, namely
    $P(\epsilon_{i}=1)=P(\epsilon_{i}=-1)=\frac{1}{2}.$
2 Main Results
In this section, we present new upper bounds for the five estimators mentioned in Section 1. In particular, compared to the existing bounds, the new bounds do not involve the dimension factor $\log d$ and hence match with the minimax lower bounds (under mild conditions).
2.1 A warming up example
We start with the most classical setting, where the noise in model (1) has exponentially decaying tails [Koltchinskii et al., 2011, Negahban and Wainwright, 2012, Klopp, 2014].
Assumption 1.
The noise variables $\xi_{i}$ satisfy $E[\xi_{i}]=0$, $E[\xi^{2}_{i}]=\sigma^{2}$ and
$E\exp(t\xi_{i}/\sigma)\leq\exp(t^{2}/L_{1}),\forall t\in\mathbb{R},$
where $L_{1}\geq 0$ is a constant.
Regarding the distribution of $X_{i}$, we consider a non-uniform sampling distribution.
Assumption 2.
$X_{i}$’s distribution satisfies
$(1)\leavevmode\nobreak$ $\max_{j}\sum_{k=1}^{m_{2}}P_{jk}\leq\frac{L_{2}}{m},$
$(2)\leavevmode\nobreak$ $\max_{k}\sum_{j=1}^{m_{1}}P_{jk}\leq\frac{L_{2}}{m},$
$(3)\leavevmode\nobreak$ $\min_{j,k}P_{jk}\geq\frac{1}{\mu m_{1}m_{2}},$
$(4)\leavevmode\nobreak$ $\max_{j,k}P_{jk}\leq\frac{L_{3}}{m\log^{3}d},$
where $L_{2},L_{3}\geq 0$ are constants, and $\mu\geq 1$ is allowed to change with the sample size or dimension.
Conditions $(1)$-$(3)$ of Assumption 2 are the same as in Klopp [2014]. $L_{2}$ is introduced to control the row and column sampling probabilities, preventing them from being too large. Condition (4) requires that the sampling doe not extremely concentrate on a few entries. Note that Condition (1) or (2) already implies $\max_{j,k}P_{jk}\apprle\frac{1}{m}$ which only differs from Condition (4) by a logarithmic factor. Hence, (4) is a very mild condition.
Under the assumption of sub-Gaussian tailed noise, a classical estimator for $A_{0}$ is:
$\hat{A}=\operatorname*{arg\,min}_{\|A\|_{\infty}\leq a}\left\{\dfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}+\lambda\|A\|_{1}\right\}.$ (2)
We now present our new result:
Theorem 1.
Let Assumptions 1 and 2 hold. Assume that $n\geq C_{1}m\log^{6}d\cdot\log^{2}n$. There exists a constant $C_{2}$ such that if we take
$\lambda=C_{2}\sigma\sqrt{\frac{1}{nm}},$
with probability at least $1-\frac{2}{d}$, the following bound holds
$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq\dfrac{C_{3}\mu^{2}\max(\sigma^{2},a^{2})rM}{n}.$ (3)
Here, $C_{1}$ is a universal constant, and $C_{2},C_{3}$ depend on the constants $L_{1},L_{2},L_{3}$ from Assumptions 1-2.
To compare with existing bounds, we mention one representative result from Klopp [2014]:
Theorem 2 (Theorem 7 in Klopp [2014]).
Let Assumption 1 and Conditions (1)-(3) in Assumption 2 hold. Assume that $n\geq C_{1}m\log^{2}d$. If we take
$\lambda=C_{2}\sigma\sqrt{\dfrac{\log d}{nm}},$
for a suitable constant $C_{2}$, the following bound holds with probability $1-\frac{3}{d}$:
$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{3}\max\left\{\dfrac{\mu^{2}\max(a^{2},\sigma^{2})rM\log d}{n},\mu a^{2}\sqrt{\frac{\log d}{n}}\right\}.$
Here, $C_{1}$ is a universal constant, and $C_{2},C_{3}$ depend on the constants $L_{1},L_{2}$ from Assumptions 1-2.
Comparing Theorems 1 and 2, we make several remarks:
  1. 1.
    Our new result eliminates the $\log d$ factor in the convergence rate. Koltchinskii et al. [2011] provides the following minimax lower bound in the case of gaussian noise and uniform sampling distribution:
    $\inf_{\hat{A}}\sup_{A_{0}}P\left(\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\geq C_{1}\dfrac{\min(\sigma^{2},a^{2})rM}{n}\right)\geq C_{2},$ (4)
    where $C_{1},C_{2}$ are universal constants, and the infimum and supremum are taken over all possible estimators and rank-$r$ matrices respectively. Combining the new bound (3) with the minimax lower bound (4) shows that the removal of the dimension factor $\log d$ establishes the minimax rate optimality of the estimator (2).
  2. 2.
    Our analysis reveals a more precise choice for the tuning parameter $\lambda$: the optimal choice for $\lambda$ should be of the order $O(\sqrt{1/(nm)})$ instead of $O(\sqrt{(\log d)/(nm)})$.
  3. 3.
    The new bound requires a slightly stronger condition on $n$, as $n\geq Cm\log^{6}d\log^{2}n$ is needed rather than $n\geq Cm\log^{2}d$. The $\log n$ term arises from the truncation of the sub-Gaussian noise, in order to apply the matrix concentration inequalities from Brailovskaya and van Handel [2024].
  4. 4.
    Compared to the sampling distribution assumption in Klopp [2014], Condition (4) of Assumption 2 is an additional condition we impose. But as we explained after Assumption 2, this is a very mild condition.
  5. 5.
    The convergence rate in Theorem 2 takes a form of the maximum between two terms. In the high-dimensional regime where $n\leq m_{1}m_{2}$, it is straightforward to verify that the first term–the desirable term–is the larger one. However, in less challenging scenarios in which $n$ is sufficiently large, the second term dominates. This limitation arises from a Frobenius norm based peeling argument that Klopp [2014] used to prove restricted strong convexity. In contrast, we have adopted a different peeling scheme introduced in Wainwright [2019], which peels certain matrix space by the infinity norm and the nuclear norm. As a result, our new bound does not include the nuisance term $O(\sqrt{(\log d)/n})$.
2.1.1 How to remove the $\log d$ factor
In this section, we discuss where the dimensional factor $\log d$ arises and how to eliminate it. Denote $L(A)=\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}$. The high-dimensional analysis of (2) includes two basic steps [Negahban et al., 2012]. The first step involves selecting $\lambda$ such that
$\lambda\geq C\|\nabla L(A_{0})\|=2C\Big{\|}\dfrac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\Big{\|},$ (5)
to ensure that $\hat{A}-A_{0}$ lies in a nearly low rank space $\Omega$. In the second step, we establish strong convexity of $L(A)$ over $\{A:A-A_{0}\in\Omega\}$, namely,
$L(\hat{A})-L(A_{0})-\langle\nabla L(A_{0}),\hat{A}-A_{0}\rangle=\frac{1}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}-A_{o}\rangle^{2}\geq C\|\hat{A}-A_{0}\|^{2}_{w(F)}-\mathcal{E},$
where $\mathcal{E}$ represents a tolerance term and $\|A\|^{2}_{w(F)}=E\langle X_{i},A\rangle^{2}$. To achieve this, we shall obtain bounds on the following empirical processes
$Z=\sup_{A-A_{0}\in\Omega}\left|\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},A-A_{0}\rangle^{2}-\|A-A_{0}\|^{2}_{w(F)}\right|.$
Bounding this type of empirical processes, after applying Talagrand concentration and symmetrization & contraction, invokes a bound on
$E\left\|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|.$ (6)
With the two steps completed, we can use some standard non-asymptotic arguments to obtain upper bounds for the convergence rate in averaged Frobenius norm. Specifically, Lemma 14 reveals how (5) and (6) influence the error rate: with probability at least $1-\frac{1}{d}$, it holds
$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\apprle\mu^{2}\left(\lambda^{2}+a^{2}E^{2}\left\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|\right)rm_{1}m_{2}.$ (7)
Traditional matrix concentration inequalities (e.g. Theorem 6.17 and Example 6.18 in Wainwright [2019]; Lemmas 5 and 6 in Klopp [2014]), yield
$\left\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\leq C_{1}\sigma\sqrt{\frac{\log d}{nm}}$
holds with high probability and
$E\left\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|\leq C_{2}\sqrt{\frac{\log d}{nm}}.$
Plugging $\lambda=C_{1}\sigma\sqrt{\frac{\log d}{nm}}$ and $E\left\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|\leq C_{2}\sqrt{\frac{\log d}{nm}}$ into (7) leads to the dimension factor dependent result in Theorem 2.
To remove the $\log d$ factor, we refine the evaluation of the two key quantities $\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\|$ and $E\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\|$. Recently, Bandeira et al. [2023] introduce new matrix concentration inequalities for random matrices with Gaussian entries using novel mathematical tools. Later, Brailovskaya and van Handel [2024] extend the results for more general distributions. This type of concentration inequalities fully take the non-commutativity of random matrices into account and provide sharper bounds. For instance, their results on the expectation of the spectral norm of a random matrix $X\in\mathbb{R}^{m_{1}\times m_{2}}$ typically take the form:
$E\|X\|\apprle\max(\|EX^{T}X\|,\|EXX^{T}\|)^{1/2}+v(X)\log^{\beta}d,$
where $\max(\|EX^{T}X\|,\|EXX^{T}\|)^{1/2}$ is the desirable part and $\beta$ is a positive constant. The term $v(X)$ is a statistic that captures the non-commutativity of the random matrix $X$ and is often smaller than $\max(\|EX^{T}X\|,\|EXX^{T}\|)^{1/2}$ in many situations. If we are given that
$v(X)\log^{\beta}d\leq\max(\|EX^{T}X\|,\|EXX^{T}\|)^{1/2},$
the bound is free of the dimensional factor $\log^{\beta}d$. Our computations confirm that this holds for $\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\|$ and $E\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\|$. Assuming some mild conditions, we can obtain
$P\left(\left\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\geq C_{1}\sigma\sqrt{\dfrac{1}{nm}}\right)\leq\dfrac{C_{2}}{d},$
and
$E\left\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|\leq C\sqrt{\dfrac{1}{nm}}.$
As a result, the $\log d$ factor is successfully removed. See Lemmas 12 and 13 for more details.
A similar issue appears in the existing results for the other four estimators. The advanced inequalities in Brailovskaya and van Handel [2024] can also be adapted to resolve the issue. We present sharper bounds for the remaining estimators in Sections 2.2-2.4.
2.2 Matrix completion with unknown variance of the noise
The choice of $\lambda$ in Theorem 2 is dependent on the variance of the noise. In practice, $\sigma$ is often unknown. To deal with this issue, Klopp [2014] proposes the following square-root lasso type estimator and studies its convergence rate.
$\hat{A}_{S}=\operatorname*{arg\,min}_{\|A\|_{\infty}\leq a}\left\{\sqrt{\dfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}}+\lambda\|A\|_{1}\right\}.$ (8)
We obtain a new result for (8).
Theorem 3.
Let Assumptions 1 and 2 hold. Let $n\geq C_{1}m\log^{6}d\log^{2}n$. If we take
$\lambda=C_{2}\sqrt{\frac{1}{nm}},$
with probability at least $1-\frac{2}{d}-2\exp(-C_{3}n)$, the following bound holds
$\dfrac{\|\hat{A}_{S}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq\dfrac{C_{4}\mu^{2}\max(\sigma^{2},a^{2})rM}{n},$ (9)
provided that $n\geq\frac{C_{2}}{10}\mu rM$. Here, $C_{1}$ is a universal constant, $C_{3}$ depends on $L_{1}$, and $C_{2},C_{4}$ depend on $L_{1},L_{2},L_{3}$, where $L_{i}$’s are constants from Assumptions 1-2.
We can compare the above bound with the one obtained in Klopp [2014].
Theorem 4 (Theorem 10 in Klopp [2014]).
Let Assumption 1 and Conditions (1)-(3) in Assumption 2 hold. If we take
$\lambda=C_{1}\sqrt{\frac{\log d}{nm}},$
the following upper bound holds with probability $1-\frac{3}{d}-2\exp(-C_{2}n)$,
$\dfrac{\|\hat{A}_{S}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{3}\left\{\dfrac{\mu^{2}\max(a^{2},\sigma^{2})rM\log d}{n},a^{2}\mu\sqrt{\frac{\log d}{n}}\right\},$
provided that $n\geq 8C_{2}\mu rM\log d$. Here $C_{1},C_{3}$ depend on $L_{1},L_{2}$ and $C_{2}$ depends on $L_{1}$, where $L_{1},L_{2}$ are constants from Assumptions 1-2.
Comparing the convergence rate in Theorem 4 and the lower bound in (4), they are not matched due to the existence of the $\log d$ factor. Our new bound in (9) resolves the issue and shows that the estimator (8) is indeed minimax rate optimal. Regarding more comparisons between Theorems 3 and 4, we refer to remarks after Theorem 2 for similar comments.
2.3 Matrix completion with heavy tailed noise
Heavy-tailed data is ubiquitous in many fields such as finance, microeconomics and biology. To address heavy-tailed noise in the context of matrix completion, several estimators have been proposed. Assuming that the second moment of the noise is finite, Minsker [2018] proposes a two-step procedure, where a truncated-type estimator $\tilde{A}$ is obtained first and then a nuclear norm penalized regression is performed on $\tilde{A}$. Fan et al. [2021] also proposes a two-step procedure while a stronger moment condition $E\xi_{i}^{2k}<\infty$ for $k>1$ is required. Elsener and van de Geer [2018] considers a class of loss functions including Huber loss function. They assume the distribution of the noise is symmetric and some additional conditions hold. In particular, Yu et al. [2024] only assumes that the noise variables have finite second moments and still derives strong properties for the estimator under study. The estimator proposed in their paper is based on the Huber loss function, which is one of the fundamental tools for addressing heavy tailed noise. We first define
$l_{\tau}(x)=\begin{cases}\frac{x^{2}}{2},\leavevmode\nobreak\ \leavevmode\nobreak\ |x|\leq\tau,\\ \tau|x|-\frac{\tau^{2}}{2},\leavevmode\nobreak\ \leavevmode\nobreak\ |x|>\tau,\\ \end{cases}$
and Huber loss is defined as $\frac{1}{n}\sum_{i=1}^{n}l_{\tau}(Y_{i}-\langle X_{i},A\rangle)$. The derivative of $l_{\tau}(\cdot)$ is denoted by
$\psi_{\tau}(x)=\begin{cases}x,|x|\leq\tau,\\ \operatorname{sign}(x)\tau,|x|>\tau.\end{cases}$
The estimator to be studied is
$\hat{A}_{H}=\operatorname*{arg\,min}_{\|A\|_{\infty}\leq a}\left\{\dfrac{1}{n}\sum_{i=1}^{n}l_{\tau}(Y_{i}-\langle X_{i},A\rangle)+\lambda\|A\|_{1}\right\}.$ (10)
We formally propose the following assumption.
Assumption 3.
$\xi_{i}$ have finite second moments, namely
$E[\xi_{i}^{2}]=\sigma^{2}.$
We obtain the following new convergence rate for $\hat{A}_{H}$.
Theorem 5.
Let Assumptions 2 and 3 hold. We additionally assume $n\geq 160m\log^{4}d$ and $\xi_{i}$ are symmetric. If we take
$\lambda=C_{1}\max(\sigma,a)\sqrt{\frac{1}{nm}},$
$\tau=\frac{\max(\sigma,a)}{\log^{2}d}\sqrt{\frac{n}{m}},$
with probability at least $1-\frac{2}{d}$, the following bound holds
$\dfrac{\|\hat{A}_{H}-A_{0}\|^{2}_{F}}{m_{1}m_{2}}\leq C_{2}\frac{\max(\sigma^{2},a^{2})rM}{n}.$ (11)
Here, $C_{1}$ and $C_{2}$ depend on the constants $L_{2},L_{3}$ from Assumption 2.
Yu et al. [2024] provides the following theorem.
Theorem 6.
Let Assumption 3 hold and set $P_{jk}=\frac{1}{m_{1}m_{2}}$. If we take
$\lambda=C_{1}\max(a,\sigma)\sqrt{\frac{\log d}{nm}},$
$\tau=C_{2}\max(a,\sigma)\sqrt{\frac{n}{m\log d}},$
then with probability at least $1-\frac{3}{d}$, we have
$\dfrac{\|\hat{A}_{H}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{3}\left\{\max(a^{2},\sigma^{2})\dfrac{rM\log d}{n},a^{2}\sqrt{\frac{\log d}{n}}\right\}$
provided that $n\geq C_{4}m\log d$. Here, $C_{1},\ldots,C_{3}$ are universal constants.
We have the following remarks for the comparison of Theorems 5 and 6.
  1. 1.
    Theorem 5 improves over Theorem 6 in several aspects: (a) Theorem 5 allows for a more general sampling distribution; (b) It drops the nuisance term $O(\sqrt{(\log d)/n})$ from the convergence rate, via the peeling by infinity and nuclear norm to the Huber loss function; (c) Like in Theorems 1 and 3, the $\log d$ factor is removed by utilizing the advanced matrix concentration inequalities in Brailovskaya and van Handel [2024].
  2. 2.
    The lower bound in (4) is derived assuming that the noises are Gaussian. Since the Gaussian distribution is of finite second moments, the lower bound also applies to the heavy tailed case. Thus, our new bound (11) together with (4) proves the minimax optimality for the estimator (10).
  3. 3.
    Theorem 5 requires symmetry of the noise, while Theorem 6 does not. If the noise is not symmetric and the sampling distribution is uniform, the bound (11) continues to hold under the additional condition $M\geq m\log^{5}d$. While this condition is very mild, this seems to be a pure technical condition and is expected to be unnecessary under future investigation. See our analysis in the proof of (11).
2.4 Matrix completion under corruption
In real applications, the true matrix $A_{0}$ may be corrupted. We consider the situation where the observed noisy entry is from a corrupted matrix $A_{0}=L_{0}+S_{0}$ where $L_{0}$ is of rank $r$ and $S_{0}$ is a sparse corruption matrix. Klopp et al. [2017] considers two corruption models: entrywise corruption and columnwise corruption. We adopt the notation of Klopp et al. [2017]. The total number of observations is denoted by $N$. The set of uncorrupted observations is denoted by $\Omega$, namely $Y_{i}=\langle X_{i},L_{0}\rangle+\xi_{i}$ for $i\in\Omega$. The set of corrupted observations is denoted by $\tilde{\Omega}$, namely $Y_{i}=\langle X_{i},L_{0}+S_{0}\rangle+\xi_{i}$ for $i\in\tilde{\Omega}$. We denote $n=|\Omega|$, $N-n=|\tilde{\Omega}|$, and $\alpha=\frac{N}{n}$.
2.4.1 Entrywise corruption
For entrywise corruption, the set of corrupted entries is denoted by
$\mathcal{I}_{1}=\{(j,k):S_{0}(j,k)\neq 0\},$
where $S_{0}(j,k)$’s are entries of the corruption matrix $S_{0}$. We assume that $S_{0}$ has at most $s$ non-zero entries, i.e. $|\mathcal{I}_{1}|\leq s$. We also need the following assumption.
Assumption 4.
For $i\in\Omega$ and $i\in\tilde{\Omega}$, the sampling is uniform:
$P(X_{i}\in e_{j}(m_{1})e_{k}^{T}(m_{2}))=\dfrac{1}{|\mathcal{I}_{1}|},\quad{\rm for\leavevmode\nobreak\ }(j,k)\in\mathcal{I}_{1},i\in\tilde{\Omega}$
and
$P(X_{i}\in e_{j}(m_{1})e_{k}^{T}(m_{2}))=\dfrac{1}{m_{1}m_{2}-|\mathcal{I}_{1}|},\quad{\rm for\leavevmode\nobreak\ }(j,k)\in\mathcal{I}_{1}^{c},i\in\Omega.$
Klopp et al. [2017] considers general sampling distributions. To ease the notation and demonstrate more clearly the ideas of our method, we focus on the uniform sampling distribution in Assumption 4. Our results can be easily extended to the general distribution case. For this corruption model, Klopp et al. [2017] considers the following estimator:
$(\hat{L}_{E},\hat{S}_{E})=\operatorname*{arg\,min}_{\begin{subarray}{c}\|L\|_{\infty}\leq a\\ \|S\|_{\infty}\leq a\end{subarray}}\left\{\dfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},L+S\rangle)^{2}+\lambda_{1}\|L\|_{1}+\lambda_{2}\|S\|_{1,*}\right\}.$ (12)
We obtain the following theorem.
Theorem 7.
Let Assumptions 1 and 4 hold. Let $n\geq C_{1}m\log^{6}d\log^{2}n$ and $N\leq m_{1}m_{2}$. Taking
$\lambda_{1}=C_{2}\max(a,\sigma)\sqrt{\frac{1}{nm}},$
$\lambda_{2}=C_{3}\max(a,\sigma)\frac{\log d}{N},$
the following bound holds with probability at least $1-\frac{6}{d}$,
$\dfrac{\|\hat{L}_{E}-L_{0}\|_{F}^{2}}{m_{1}m_{2}}+\dfrac{\|\hat{S}_{E}-S_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{4}\Bigg{(}\alpha\max(\sigma^{2},a^{2})\left(\dfrac{rM}{n}+\dfrac{|\tilde{\Omega}|\log d}{n}\right)+\dfrac{a^{2}s}{m_{1}m_{2}}\Bigg{)}.$ (13)
Here, $C_{1}$ is a universal constant, and $C_{2},C_{3},C_{4}$ depend on the constant $L_{1}$ from Assumption 1.
Klopp et al. [2017] gives the following result.
Theorem 8 (Corollary 11 of Klopp et al. [2017]).
Let Assumptions 1 and 4 hold. Assume that $n\geq C_{1}m\log d\log^{2}m$ and $N\leq m_{1}m_{2}$. Taking
$\lambda_{1}=C_{2}\max(a,\sigma)\sqrt{\frac{\log d}{nm}},$
$\lambda_{2}=C_{3}\max(a,\sigma)\frac{\log d}{N},$
the following bound holds with probability at least $1-\frac{6}{d}$,
$\dfrac{\|\hat{L}_{E}-L_{0}\|_{F}^{2}}{m_{1}m_{2}}+\dfrac{\|\hat{S}_{E}-S_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{4}\left(\alpha\max(\sigma^{2},a^{2})\left(\dfrac{rM\log d}{n}+\dfrac{|\tilde{\Omega}|\log d}{n}\right)+\dfrac{a^{2}s}{m_{1}m_{2}}\right).$ (14)
Here, $C_{1}$ is a universal constant, and $C_{2},C_{3},C_{4}$ depend on the constant $L_{1}$ from Assumption 1.
Klopp et al. [2017] also obtains the minimax lower bound under this setting assuming $\xi_{i}\sim\mathcal{N}(0,\sigma^{2})$:
$\inf_{(\hat{L},\hat{S})}\sup_{\mathcal{A}}P\left(\dfrac{\|\hat{L}-L_{0}\|_{F}^{2}}{m_{1}m_{2}}+\dfrac{\|\hat{S}-S_{0}\|_{F}^{2}}{m_{1}m_{2}}\apprge\min(\sigma^{2},a^{2})\left(\dfrac{rM+|\tilde{\Omega}|}{n}+\dfrac{s}{m_{1}m_{2}}\right)\right)\geq C.$ (15)
Here, $C\in(0,1)$, the infimum is taken over all possible estimators for $(L_{0},S_{0})$, and $\mathcal{A}$ is the parameter space such that $\text{rank}(L_{0})\leq r$ and $|\mathcal{I}_{1}|\leq s$. Our new result removes the $\log d$ factor in the term $\frac{rM\log d}{n}$ in (14). When $rM\geq|\tilde{\Omega}|\log d$, namely the number of corrupted observations is not too large, (13) and (15) together show that the estimator (12) is minimax rate optimal.
2.4.2 Columnwise corruption
For columnwise corruption model, we assume there are at most $s$ corrupted columns, namely $|\mathcal{I}_{2}|\leq s$, where $\mathcal{I}_{2}=\{1\leq i\leq m_{2};S_{0}^{i}\neq\textbf{0}\in\mathbb{R}^{m_{1}}\}$ includes the indices of non-zero columns of $S_{0}$. And we make the following assumption.
Assumption 5.
For $i\in\Omega$, the sampling distribution is uniform:
$P(X_{i}=e_{j}(m_{1})e_{k}^{T}(m_{2}))=\dfrac{1}{m_{1}(m_{2}-|\mathcal{I}_{2}|)}.$
For $i\in\tilde{\Omega}$, the sampling is uniform:
$P(X_{i}=e_{j}(m_{1})e_{k}^{T}(m_{2}))=\dfrac{1}{m_{1}|\mathcal{I}_{2}|}.$
Klopp et al. [2017] proposes the following estimator:
$(\hat{L}_{C},\hat{S}_{C})=\operatorname*{arg\,min}_{\begin{subarray}{c}\|L\|_{\infty}\leq a\\ \|S\|_{\infty}\leq a\end{subarray}}\left\{\dfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},L+S\rangle)^{2}+\lambda_{1}\|L\|_{1}+\lambda_{2}\|S\|_{2,1}\right\}.$ (16)
We obtain the following theorem.
Theorem 9.
Let Assumptions 1 and 5 hold. Assume $n\geq 2\log d\max(m_{2},m\log^{2}m)$, $N\leq m_{1}m_{2}$, and $n\geq C_{1}m\log^{6}d\log^{2}n$. Taking
$\lambda_{1}=C_{2}\max(a,\sigma)\sqrt{\frac{1}{Nm}},$
$\lambda_{2}=C_{3}\max(a,\sigma)\sqrt{\frac{\log d}{Nm_{2}}},$
with probability at least $1-\frac{6}{d}$, we have
$\dfrac{\|\hat{L}_{C}-L_{0}\|_{F}^{2}}{m_{1}m_{2}}+\dfrac{\|\hat{S}_{C}-S_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{4}\left(\alpha\max(\sigma^{2},a^{2})\left(\dfrac{rM}{n}+\dfrac{|\tilde{\Omega}|\log d}{n}\right)+\dfrac{a^{2}s}{m_{2}}\right).$ (17)
Here, $C_{1}$ is a universal constant, and $C_{2},C_{3},C_{4}$ depend on the constant $L_{1}$ from Assumption 1.
Klopp et al. [2017] obtains the following result.
Theorem 10 (Corollary 7 in Klopp et al. [2017]).
Let Assumptions 1 and 5 hold. Assume $N\leq m_{1}m_{2}$, and $n\geq 2\log d\max(m_{2},m\log^{2}m)$. Taking
$\lambda_{1}=C_{2}\max(a,\sigma)\sqrt{\frac{\log d}{Nm}},$
$\lambda_{2}=C_{3}\max(a,\sigma)\sqrt{\frac{\log d}{Nm_{2}}},$
the following bound holds with probability at least $1-\frac{6}{d}$,
$\dfrac{\|\hat{L}_{C}-L_{0}\|_{F}^{2}}{m_{1}m_{2}}+\dfrac{\|\hat{S}_{C}-S_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{4}\left(\alpha\max(\sigma^{2},a^{2})\left(\dfrac{rM\log d}{n}+\dfrac{|\tilde{\Omega}|\log d}{n}\right)+\dfrac{a^{2}s}{m_{2}}\right),$ (18)
where $C_{1}$ is a universal constant, and $C_{2},C_{3},C_{4}$ depend on the constant $L_{1}$ from Assumption 1.
According to Theorem 3 in Klopp et al. [2017], we have the following minimax lower bound for columnwise corruption assuming $\xi_{i}\sim\mathcal{N}(0,\sigma^{2})$
$\inf_{(\hat{L},\hat{S})}\sup_{\mathcal{A}}P_{A_{0}}\left(\dfrac{\|\hat{L}-L_{0}\|_{F}^{2}}{m_{1}m_{2}}+\dfrac{\|\hat{S}-S_{0}\|_{F}^{2}}{m_{1}m_{2}}\apprge\min(\sigma^{2},a^{2})\left(\dfrac{rM+|\tilde{\Omega}|}{n}+\dfrac{s}{m_{2}}\right)\right)\geq C,$ (19)
where $C\in(0,1)$ is a universal constant, the infimum is taken over all the estimators, and $\mathcal{A}$ is the parameter space such that $\text{rank}(L_{0})\leq r$ and $|\mathcal{I}_{2}|\leq s$. We remove the $\log d$ factor attached to $\frac{rM}{n}$ in (18). Comparing (19) and (17), we can conclude that the estimator (16) is minimax rate optimal provided that $rM\geq|\tilde{\Omega}|\log d$, i.e. the set of corrupted observations is not too large.
3 Conclusion
In this paper, we considered three popular matrix completion settings: (1) the unknown matrix is of low rank and the noise is sub-Gaussian; (2) the unknown matrix is of low rank and the noise is heavy tailed; (3) the unknown matrix is a low rank matrix corrupted by a sparse matrix in the entrywise or columnwise sense, and the noise is sub-Gaussian. We revisited five existing estimators in these settings, and developed sharper upper bounds to establish their minimax rate optimality. The key technical contribution lies in employing powerful matrix concentration inequalities introduced in Brailovskaya and van Handel [2024] to successfully remove the dimension factor $\log d$ that has remained in the upper bounds prior to our work. The $\log d$ factor problem not only exists for the fives estimators studied in this paper, but is ubiquitous throughout the literature. We believe that many other existing results on matrix completion can be possibly improved by applying these new advanced inequalities.
4 Proof
4.1 Proof of sharp bounds for spectral norms
In this section, we present advanced inequalities introduced in [Brailovskaya and van Handel, 2024] and how to use them. Let $Q=\sum_{i=1}^{n}Q_{i}\in\mathbb{R}^{m_{1}\times m_{2}}$ be a random matrix where $Q_{i}$ are independent random matrices with mean zero. We define
$\gamma(Q)$ $=\left(\max(\|EQQ^{T}\|,\|EQ^{T}Q\|)\right)^{1/2},$
$\gamma_{*}(Q)$ $=\sup_{\|y\|_{2}=\|z\|_{2}=1}\left(E(y^{T}Qz)^{2}\right)^{1/2},$
$g(Q)$ $=\|\operatorname{Cov}(Q)\|^{1/2},$
$R(Q)$ $=\max_{1\leq i\leq n}\|Q_{i}\|,$
where $\operatorname{Cov}(Q)$ is the covariance matrix of all the entries of $Q$ and its dimension is $m_{1}m_{2}\times m_{1}m_{2}$.
Theorem 11 (Corrollary 2.17 in [Brailovskaya and van Handel, 2024]).
With the above notations, there exist a universal constant $C$ such that
$P(\|Q\|\geq 2\gamma(Q)+C(g(Q)^{1/2}\gamma(Q)^{1/2}\log^{3/4}d+\gamma_{*}(Q)t^{1/2}+R(Q)^{1/3}\gamma(Q)^{2/3}t^{2/3}+R(Q)t))\leq de^{-t}$ (20)
and a universal constant $C$ such that
$E\|Q\|\leq 2\gamma(Q)+C(g(Q)^{1/2}\gamma(Q)^{1/2}\log^{3/4}d+R(Q)^{1/3}\gamma(Q)^{2/3}\log^{2/3}d+R(Q)\log d).$ (21)
The matrix inequalities in [Brailovskaya and van Handel, 2024] are formulated for $Q_{i}\in\mathbb{R}^{d\times d}$ (not necessarily self-adjoint), and can be extended to rectangular matrices in the above form. See Remark 2.1 in [Brailovskaya and van Handel, 2024]. Applying the above inequalities, we have the following results.
Lemma 12.
Assume the sampling matrices $X_{i}$ satisfy Assumption 2 and the noise random variables $\xi_{i}$ satisfy Assumption 1. The following inequality holds with probability at least $1-\frac{1}{d}$,
$\left\|\dfrac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\leq C\sigma\sqrt{\dfrac{1}{nm}},$ (22)
given $n\geq m\log^{6}d\cdot\log^{2}n$. Here $C$ is a constant dependent on $L_{1}\sim L_{3}$. Assume $X_{i}$ satisfy Assumption 2 , $\xi_{i}$ satisfy Assumption 3 and are symmetric. Taking $\tau=\frac{\max(\sigma,a)}{\log^{2}d}\sqrt{\frac{n}{m}}$, the following inequality holds with probability at least $1-\frac{1}{d}$,
$\left\|\dfrac{1}{n}\sum_{i=1}^{n}\psi_{\tau}(\xi_{i})X_{i}\right\|\leq C\max(\sigma,a)\sqrt{\dfrac{1}{nm}}.$ (23)
where $C$ is dependent on $L_{2}$ and $L_{3}$.
Lemma 13.
If $n\geq m\log^{4}d$, we have the following result:
$E\left\|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|\leq C\sqrt{\frac{1}{nm}},$ (24)
where $C$ is dependent on $L_{2}$ and $L_{3}$.
Proof of Lemma 12.
Under the notation of Theorem 11, $Q_{i}=\frac{\xi_{i}X_{i}}{n}$. Note $\|Q_{i}\|=\frac{1}{n}|\xi_{i}|$. Since $R(Q)$ requires $\|Q_{i}\|$ to be bounded, we need to truncate $\xi_{i}$ to $\xi_{i}\textbf{1}(|\xi_{i}|\leq\tau)$ for some $\tau\geq 0$. Define $\bar{\xi}_{i}=\xi_{i}\textbf{1}(|\xi_{i}|\leq\tau)-E[\xi_{i}\textbf{1}(|\xi_{i}|\leq\tau)]$. Note $E[\bar{\xi}_{i}]=0$ and $E[\bar{\xi}_{i}^{2}]\leq\sigma^{2}$. We now compute the four parameters for $Q=\frac{1}{n}\sum_{i=1}^{n}\bar{\xi}_{i}X_{i}$.
1. $\gamma(Q)\leq\sqrt{\frac{L_{2}\sigma^{2}}{nm}}$. Note $E[X_{i}X_{i}^{T}]$ is a diagonal matrix whose elements on the diagonal line are $\sum_{k=1}^{m_{2}}P_{jk},j=1,2,\ldots,m_{1}$, which are less than $\frac{L_{2}}{m}$. The same bound holds for $E[X_{i}^{T}X_{i}]$. Since
$E[QQ^{T}]=\frac{1}{n^{2}}\sum_{i=1}^{n}E[\bar{\xi}_{i}^{2}X_{i}X_{i}^{T}]=\frac{1}{n}E[\bar{\xi}_{1}^{2}X_{1}X_{1}^{T}],$
we have $\|E[QQ^{T}]\|\leq\frac{L_{2}\sigma^{2}}{nm}$. Similarly, $\|E[Q^{T}Q]\|\leq\frac{L_{2}\sigma^{2}}{nm}$. Thus $\gamma(Q)\leq\sqrt{\frac{L_{2}\sigma^{2}}{nm}}$.
2. $\gamma_{*}(Q)\leq\sqrt{\frac{L_{3}\sigma^{2}}{nm\log^{3}d}}$. For $y\in\mathbb{R}^{m_{1}},z\in\mathbb{R}^{m_{2}}$ with $\|y\|_{2}=\|z\|_{2}=1$, we have the following result:
$E(y^{T}\bar{\xi}_{i}X_{i}z)^{2}$ $=E(y^{T}X_{i}z)^{2}E[\bar{\xi}_{i}^{2}]$
$=E[\bar{\xi}_{i}^{2}]\sum_{j,k}P_{jk}y^{2}(j)z^{2}(k)$
$\leq\sigma^{2}\sum_{j,k}P_{jk}y^{2}(j)z^{2}(k)$
$\leq\dfrac{L_{3}\sigma^{2}}{m\log^{3}d}.$
where we use $\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}y^{2}(j)z^{2}(k)=\|y\|_{2}^{2}\cdot\|z\|_{2}^{2}=1$ and $\max P_{jk}\leq\frac{L_{3}}{m\log^{3}d}$. Since $\bar{\xi}_{i}X_{i}$ are independent and mean-zero, $E(y^{T}Qz)^{2}=\dfrac{1}{n^{2}}\sum_{i=1}^{n}E\left(y^{T}\bar{\xi}_{i}X_{i}z\right)^{2}\leq\frac{L_{3}\sigma^{2}}{nm\log^{3}d}$. Thus $\gamma_{*}(Q)\leq\sqrt{\frac{L_{3}\sigma^{2}}{nm\log^{3}d}}$.
(3) $g(Q)\leq\sqrt{\frac{L_{3}\sigma^{2}}{nm\log^{3}d}}$. Since $\bar{\xi}_{i}X_{i}$’s are i.i.d., $\operatorname{Cov}(Q)=\dfrac{1}{n^{2}}\sum_{i=1}^{n}\operatorname{Cov}(\bar{\xi}_{i}X_{i})=\frac{1}{n}\operatorname{Cov}(\bar{\xi}_{1}X_{1})$. Define $\textbf{1}_{(j,k)}=\textbf{1}(X_{1}=e_{j}(m_{1})e_{k}(m_{2})^{T})$ and denote $I(j,k)=\bar{\xi}_{i}\textbf{1}_{(j,k)}$. Note the following facts
$E[I(j,k)]=0,$
$I(j,k)I(l,h)=0\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ (j,k)\neq(l,h).$
Thus $\text{Cov}\big{(}I(j,k),I(l,h)\big{)}=E[I(j,k)I(l,h)]-E[I(j,k)]E[I(l,h)]=0$ for $(j,k)\neq(l,h)$. Then we have
$\text{Cov}(\bar{\xi}_{i}X_{i})\Big{(}(j,k),(l,h)\Big{)}=\begin{cases}E[\bar{\xi}^{2}_{i}]P_{jk}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ \leavevmode\nobreak\ (j,k)=(l,h),\\ 0\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ (j,k)\neq(l,h).\end{cases}$
Since $P_{jk}\leq\frac{L_{3}}{m\log^{3}d}$, $\|\operatorname{Cov}(\bar{\xi}_{i}X_{1})\|\leq\frac{L_{3}\sigma^{2}}{m\log^{3}d}$. Thus, $g(Q)\leq\sqrt{\frac{L_{3}\sigma^{2}}{nm\log^{3}d}}$ holds.
4.$R(Q)\leq\frac{2\tau}{n}$. Obviously, $\|\bar{\xi}_{i}X_{i}\|=|\bar{\xi}_{i}|$, and $\max_{1\leq i\leq n}|\bar{\xi}_{i}|\leq 2\tau$.
Now, we can plug the four estimates in (20). Taking $t=3\log d$, we get that with probability at least $1-\frac{1}{d^{2}}$, the following holds
$\|Q\|$ $\leq 2\sqrt{\frac{\sigma^{2}}{nm}}$
$+C\Bigg{(}\left(\frac{\sigma^{2}}{nm}\right)^{1/2}\cdot\left(\dfrac{\log^{3/4}d}{\log^{3/4}d}\right)+\left(\frac{\sigma^{2}}{nm}\right)^{1/2}\cdot\left(\dfrac{\log^{1/2}d}{\log^{3/2}d}\right)$
$+\left(\frac{\log^{2}d\cdot\tau}{n}\right)^{1/3}\cdot\left(\frac{\sigma^{2}}{nm}\right)^{1/3}+\dfrac{\log d\cdot\tau}{n}\Bigg{)},$ (25)
where $C$ is dependent on $L_{2}\sim L_{3}$. We can see the quantities in the first line and the second line of RHS are less than $C\sqrt{\frac{\sigma^{2}}{nm}}$.
To prove the dimension free bound for $\frac{1}{n}\sum_{i=1}^{n}\psi_{\tau}(\xi_{i})X_{i}$, we consider $Q_{i}=\frac{\psi_{\tau}(\xi_{i})}{n}X_{i}$. Since $E[\psi_{\tau}(\xi_{i})]=0$, $E[\psi(\xi_{i})^{2}]\leq\sigma^{2}$, $|\psi_{\tau}(\xi_{i})|\leq\tau$, the estimates for the four parameters still apply. Then, with $\tau=\frac{\max(\sigma,a)}{\log^{2}d}\sqrt{\frac{n}{m}}$, we obtain
$\left(\frac{\log^{2}d}{n}\cdot\frac{\max(\sigma,a)}{\log^{2}d}\sqrt{\frac{n}{m}}\right)^{1/3}\cdot\left(\frac{\sigma^{2}}{nm}\right)^{1/3}\leq\max(\sigma,a)\sqrt{\dfrac{1}{nm}},$
$\dfrac{\log d}{n}\frac{\max(\sigma,a)}{\log^{2}d}\sqrt{\frac{n}{m}}\leq\max(\sigma,a)\sqrt{\dfrac{1}{nm}}.$
Then the upper bound of $\|\frac{1}{n}\sum_{i=1}^{n}\psi_{\tau}(\xi_{i})X_{i}\|$ is free of the $\log d$ factor. For the unbounded noise case, we take $\tau=C_{1}\sigma\log n\cdot\log d$. To make the quantities in the last line be less than $C\sqrt{\frac{\sigma^{2}}{nm}}$, we need
$\left(\frac{\sigma\log^{3}d\cdot\log n}{n}\right)^{1/3}\cdot\left(\frac{\sigma^{2}}{nm}\right)^{1/3}\leq\sqrt{\dfrac{\sigma^{2}}{nm}},$
$\dfrac{\sigma\log^{2}d\cdot\log n}{n}\leq\sqrt{\dfrac{\sigma^{2}}{nm}}.$
Those imply
$n\geq m\log^{6}d\cdot\log^{2}n,$
$n\geq m\log^{4}d\log^{2}n.$
On $\{\max_{1\leq i\leq n}|\xi_{i}|\leq\tau\}$,
$\left\|\dfrac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|$ $\leq\|Q\|+\left\|\dfrac{1}{n}\sum_{i=1}^{n}E[\xi_{i}\textbf{1}(|\xi_{i}|\leq\tau)]X_{i}\right\|$
$\leq\|Q\|+|E[\xi_{1}\textbf{1}(|\xi_{1}|\leq\tau)]|.$
Since $E[\xi_{i}]=0$, by Cauchy-Schwarz inequality, for $d$ sufficiently large,
$|E\xi_{i}\textbf{1}(|\xi_{i}|\leq\tau)|$ $=|E\xi_{i}\textbf{1}(|\xi_{i}|>\tau)|$
$\leq E|\xi_{i}|\textbf{1}(|\xi_{i}|>\tau)$
$=\sigma E[|\xi_{i}|/\sigma]\textbf{1}(|\xi_{i}|/\sigma>C_{1}\log d\cdot\log n)$
$\leq\sigma E^{1/2}[|\xi_{i}|^{2}/\sigma^{2}]\cdot P^{1/2}(|\xi_{i}|/\sigma>C_{1}\log d\cdot\log n)$
$\lesssim\dfrac{L_{1}\sigma}{n^{C\log d}}$
$\lesssim\frac{L_{1}\sigma}{n},$
for a suitably chosen constant $C_{1}$ and sufficiently large $d$. $\frac{\sigma}{n}\leq\sqrt{\frac{\sigma^{2}}{nm}}$ holds under $n\geq m\log^{6}d\cdot\log^{2}n$. By the concentration inequality for the maximum of sub-Gaussian random variables (see Theorem 2.2.27 in [Talagrand, 2014]), $P(\max_{1\leq i\leq n}|\xi_{i}|\leq\tau)\geq 1-\frac{1}{d^{2}}$ for some suitably chosen $C_{1}$. Thus, combining the pieces, $\left\|\dfrac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\leq C\sigma\sqrt{\frac{1}{nm}}$ holds with probability at least $1-\frac{1}{d}$, where $C$ is dependent on $L_{1}\sim L_{3}$. ∎
Proof of Lemma 13.
Let $Q=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}$. Since $\epsilon_{i}\leq 1$, we have $R(Q)=\frac{1}{n}$. And we have estimated the four parameters in the above. In the last line of (25), we need
$\left(\frac{\log^{2}d}{n}\right)^{1/3}\cdot\left(\frac{1}{nm}\right)^{1/3}\leq C\sqrt{\dfrac{1}{nm}},$
$\dfrac{\log d}{n}\leq C\sqrt{\dfrac{1}{nm}},$
which requires $n\geq m\log^{4}d.$
4.2 Proof of Theorem 1
We first derive a general bound in Lemma 14, which reveals the role of the two key quantities $\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\|$ and $E\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\|$ in the error rate. Then combining it with the sharp spectral norm results for $\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\|$ and $E\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\|$ from Section 4.1 will complete the proof.
Lemma 14.
Let Assumption 1 and 2 hold. Denote $R=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}$. We take
$\lambda\geq 3\left\|\frac{2}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|.$
Then, the following bound holds with probability at least $1-\frac{1}{d}$,
$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C\left(\mu^{2}(\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu^{2}a^{2}\log d}{n}\right),$ (26)
where $C$ is a universal constant.
Now we give the proof of Theorem 1.
Proof of Theorem 1.
By (22), we know that with probability at least $1-\frac{1}{d}$,
$\left\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\leq C_{1}\sigma\sqrt{\dfrac{1}{nm}}.$
where $C_{1}$ is dependent on $L_{1}\sim L_{3}$. Thus we can take $\lambda=6C_{1}\sigma\sqrt{\frac{1}{nm}}$. And by (24),
$\left\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|\leq C_{2}\sqrt{\dfrac{1}{nm}},$
where $C_{2}$ depends on $L_{2}\sim L_{3}$. Plugging the two above estimates in (26), we obtain, with probability at least $1-\frac{2}{d}$,
$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\lesssim\dfrac{\mu^{2}\max(a^{2},\sigma^{2})rM}{n}+\frac{\mu^{2}a^{2}\log d}{n},$
where the second term on RHS can be covered by the first term. ∎
Proof of Lemma 14.
Since $\hat{A}$ is the solution of (2), we have
$\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},\hat{A}\rangle)^{2}+\|\hat{A}\|_{1}\leq\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A_{0}\rangle)^{2}+\|A_{0}\|_{1},$
which implies
$\frac{1}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}-A_{0}\rangle^{2}$ $\leq 2\left\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\|\hat{A}-A_{0}\|_{1}+\lambda(\|A_{0}\|_{1}-\|\hat{A}\|_{1})$
$\leq\frac{4}{3}\lambda\|\hat{A}-A_{0}\|_{1}.$
By Lemma 22, we know that
$\|\hat{A}-A_{0}\|_{1}\leq 3\|P_{A_{0}}(\hat{A}-A_{0})\|_{1}\leq 3\sqrt{r}\|\hat{A}-A_{0}\|_{F},$ (27)
which implies
$\frac{1}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}-A_{0}\rangle^{2}\leq 5\lambda\sqrt{r}\|\hat{A}-A_{0}\|_{F}.$
By Lemma 18 and $\|\hat{A}-A_{0}\|^{2}_{w(F)}\geq\frac{\|\hat{A}-A_{0}\|^{2}_{F}}{\mu m_{1}m_{2}}$, we obtain that with probability at least $1-\frac{1}{d}$,
$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{\mu m_{1}m_{2}}\leq C\left(\dfrac{\mu\|\hat{A}-A_{0}\|_{\infty}^{2}\log d}{n}+\|\hat{A}-A_{0}\|_{\infty}\|\hat{A}-A_{0}\|_{1}E\|R\|+\lambda\sqrt{r}\|\hat{A}-A_{0}\|_{F}\right).$
By $\|\hat{A}-A_{0}\|_{\infty}\leq 2a$, we have
$\dfrac{\mu\|\hat{A}-A_{0}\|_{\infty}^{2}\log d}{n}\leq\frac{4\mu a^{2}\log d}{n}.$
By $\|\hat{A}-A_{0}\|_{\infty}\leq 2a$ and (27), we obtain
$\quad\|\hat{A}-A_{0}\|_{\infty}\|\hat{A}-A_{0}\|_{1}E\|R\|+\lambda\sqrt{r}\|\hat{A}-A_{0}\|_{F}$
$\leq 2a\sqrt{r}\|\hat{A}-A_{0}\|_{F}E\|R\|+\frac{1}{10C}\frac{\|\hat{A}-A_{0}\|_{F}^{2}}{\mu m_{1}m_{2}}+10C\mu\lambda^{2}rm_{1}m_{2}$
$\leq\frac{1}{5C}\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{\mu m_{1}m_{2}}+40C\mu a^{2}E^{2}\|R\|rm_{1}m_{2}+10C\mu\lambda^{2}rm_{1}m_{2}.$
Combining the pieces above, we obtain the following bound holds with probability at least $1-\frac{1}{d}$
$\frac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\lesssim\mu^{2}(\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\dfrac{a^{2}\mu^{2}\log d}{n}.$
4.3 Proof of Theorem 3
In this section, denote $R=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}$ and $L(A)=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}}$. We use the following general result to prove Theorem 3.
Lemma 15.
Let Assumptions 1 and 2 hold. Assume $\mu\lambda^{2}rm_{1}m_{2}\leq\frac{1}{10}$. We take
$\lambda\geq\frac{3\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\|}{L(A_{0})}.$
Then the following bound holds with probability at least $1-\frac{1}{d}$,
$\frac{\|\hat{A}_{S}-A_{0}\|^{2}_{F}}{m_{1}m_{2}}\lesssim\mu^{2}(L^{2}(A_{0})\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu^{2}a^{2}\log d}{n}.$ (28)
Proof of Theorem 3.
Note $L(A_{0})=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\xi_{i}^{2}}$. By Theorem 3.1.1. in [Vershynin, 2018], there exists a universal constant $C$ such that
$\left|\sqrt{\frac{1}{n}\sum_{i=1}^{n}\xi^{2}_{i}}-\sigma\right|\leq C\sigma\sqrt{\frac{L_{1}t}{n}}$
holds with probability at least $1-2\exp(-t)$. Then, there exist a constant $C_{1}$ dependent on $L_{1}$ such that
$0.9\sigma\leq\sqrt{\frac{1}{n}\sum_{i=1}^{n}\xi_{i}}\leq 1.1\sigma$
holds with probability at least $1-2\exp(-C_{1}n)$. By (22), with probability at least $1-\frac{1}{d}$,
$\left\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\leq C_{2}\sigma\sqrt{\dfrac{1}{nm}}.$
Thus, we can take $\lambda=4C_{2}\sqrt{\frac{1}{nm}}$. And by (24),
$E\left\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|\leq C_{3}\sqrt{\dfrac{1}{nm}}.$
Here $C_{2}$ depends on $L_{1}\sim L_{3}$ and $C_{3}$ depends on $L_{2}\sim L_{3}$. Plugging in the above estimates, we obtain that
$\dfrac{\|\hat{A}_{S}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\lesssim\dfrac{\mu^{2}\max(a^{2},\sigma^{2})rM}{n}+\frac{\mu^{2}a^{2}\log d}{n},$
holds with probability at least $1-\frac{2}{d}-\exp(-C_{1}n)$. And the second term on RHS can be covered by the first term. ∎
Proof of Lemma 15.
The basic steps follow the proof in [Klopp, 2014]. Note
$\frac{1}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}_{S}-A_{0}\rangle^{2}=L^{2}(\hat{A}_{S})-L^{2}(A_{0})+\frac{2}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}_{S}-A_{0}\rangle.$
Since $L(\hat{A}_{S})$ is the solution of (8),
$L(\hat{A}_{S})-L(A_{0})\leq\lambda(\|A_{0}\|_{1}-\|\hat{A}_{S}\|_{1}).$
Immediately, we obtain
$L^{2}(\hat{A}_{S})-L^{2}(A_{0})\leq(L(\hat{A}_{S})+L(A_{0}))(\lambda\|\hat{A}_{s}-A_{0}\|_{1}).$
Again, by the definition of $\hat{A}_{S}$, we obtain
$L(\hat{A})+L(A_{0})$ $\leq 2L(A_{0})+\lambda(\|A_{0}\|_{1}-\|\hat{A}_{S}\|_{1})\leq 2L(A_{0})+\lambda\|\hat{A}-A_{0}\|_{1}.$
Gathering the pieces, we get
$\frac{1}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}_{S}-A_{0}\rangle^{2}$ $\leq\frac{2}{n}\sum_{i=1}^{n}\langle\xi_{i}X_{i},\hat{A}_{S}-A_{0}\rangle+(2L(A_{0})+\lambda\|\hat{A}_{S}-A_{0}\|_{1})(\lambda\|\hat{A}-A_{0}\|_{1})$
$\leq\frac{8L(A_{0})}{3}\lambda\|\hat{A}_{S}-A_{0}\|_{1}+\lambda^{2}\|\hat{A}_{S}-A_{0}\|_{1}^{2}.$
By Lemma 18, we get with probability at least $1-\frac{1}{d}$,
$\|\hat{A}_{S}-A_{0}\|^{2}_{w(F)}$ $\leq C\left(\|\hat{A}-A_{0}\|^{2}_{\infty}\frac{\mu\log d}{n}+\|\hat{A}_{S}-A_{0}\|_{\infty}\|\hat{A}_{S}-A_{0}\|_{1}E\|R\|\right)$
$+\frac{8L(A_{0})}{3}\lambda\|\hat{A}_{S}-A_{0}\|_{1}+\lambda^{2}\|\hat{A}_{S}-A_{0}\|_{1}^{2}.$
Since $\|\hat{A}_{S}-A_{0}\|_{\infty}\leq 2a$,
$\|\hat{A}_{S}-A_{0}\|^{2}_{\infty}\frac{\mu\log d}{n}\leq\frac{4\mu a^{2}\log d}{n}.$
By lemma (22), $\|\hat{A}_{S}-A_{0}\|_{1}\leq 3\|P_{A_{0}}(\hat{A}_{S}-A_{0})\|_{1}\leq 3\sqrt{r}\|\hat{A}_{S}-A_{0}\|_{F}$. Then we have
$\|\hat{A}_{S}-A_{0}\|_{\infty}\|\hat{A}_{S}-A_{0}\|_{1}E\|R\|$ $\leq 6a\sqrt{r}E\|R\|\|\hat{A}_{S}-A_{0}\|_{F}$
$\leq 360C\mu a^{2}E^{2}\|R\|rm_{1}m_{2}+\frac{1}{10C}\frac{\|\hat{A}_{S}-A_{0}\|_{F}^{2}}{\mu m_{1}m_{2}}.$
Similarly,
$\quad\left\|\frac{2}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\|\hat{A}_{S}-A_{0}\|_{1}+L(A_{0})\lambda\|\hat{A}_{S}-A_{0}\|_{1}$ $\leq\frac{5}{3}L(A_{0})\lambda\sqrt{r}\|\hat{A}_{S}-A_{0}\|_{F}$
$\leq 40\mu L^{2}(A_{0})\lambda^{2}rm_{1}m_{2}+\frac{1}{10}\frac{\|\hat{A}_{S}-A_{0}\|_{F}^{2}}{\mu m_{1}m_{2}}.$
Given $\mu\lambda^{2}rm_{1}m_{2}\leq\frac{1}{10}$,
$\lambda^{2}\|\hat{A}_{S}-A_{0}\|_{1}^{2}\leq\frac{1}{10}\frac{\|\hat{A}_{S}-A_{0}\|_{F}^{2}}{\mu m_{1}m_{2}}.$
Combining the pieces, we get
$\|\hat{A}_{S}-A_{0}\|^{2}_{w(F)}\leq C\left(\mu^{2}(L^{2}(A_{0})\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu^{2}a^{2}\log d}{n}\right)+\frac{3}{10}\frac{\|\hat{A}_{S}-A_{0}\|_{F}^{2}}{\mu m_{1}m_{2}}.$
Since $\|\hat{A}_{S}-A_{0}\|^{2}_{w(F)}\geq\frac{\|\hat{A}_{S}-A_{0}\|^{2}_{F}}{\mu m_{1}m_{2}}$, we obtain the following bound holds with probability at least $1-\frac{1}{d}$,
$\frac{\|\hat{A}_{S}-A_{0}\|^{2}_{F}}{m_{1}m_{2}}\lesssim\mu^{2}(L^{2}(A_{0})\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu^{2}a^{2}\log d}{n}.$
4.4 Proof of Theorem 5
The proof is inspired by [Yu et al., 2024]. Throughout this section, we set $L(A)=\frac{1}{n}\sum_{i=1}^{n}l_{\tau}(Y_{i}-\langle X_{i},A\rangle)$. In this section, $c_{3}$ and $c_{4}$ are specific constants those values do not change at each occurrence. We have the following general result.
Lemma 16.
Let Assumptions 2 and 3 hold. Let $n\geq 160m\log^{4}d$ and $\tau=\frac{\max(\sigma,a)}{\log^{2}d}\sqrt{\frac{n}{m}}$. If we take
$\lambda\geq 3\left\|\frac{1}{n}\sum_{i=1}^{n}\psi_{\tau}(\xi_{i})X_{i}\right\|,$
then the following bound holds with probability at least $1-\frac{1}{d}$,
$\frac{\|\hat{A}_{H}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\lesssim\mu^{2}(\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu^{2}\log d}{n}.$
Here $R$ is defined as $R=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}},$ where $\chi_{i}=\{|\xi_{i}|\leq\tau/2\}$.
Proof of Theorem 5.
We let $\tau=\frac{\max(a,\sigma)}{\log^{2}d}\sqrt{\frac{n}{m}}$. By (23), with probability at least $1-\frac{1}{d}$,
$\left\|\frac{1}{n}\sum_{i=1}^{n}\psi_{\tau}(\xi_{i})X_{i}\right\|\leq C_{1}\max(\sigma,a)\sqrt{\dfrac{1}{nm}},$ (29)
where $C_{1}$ depends on $L_{2}\sim L_{3}$. Since $E[\epsilon_{i}\textbf{1}_{\chi_{i}}]=0$ and $|\epsilon_{i}\textbf{1}_{\chi_{i}}|\leq 1$,by (24),
$E\left\|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}\right\|\leq C_{2}\sqrt{\frac{1}{nm}},$ (30)
where $C_{2}$ depends on $L_{2}\sim L_{3}$. Then all we need to do is to plug in the above two estimates in the bound in Lemma 16.
In the above, if we do not assume $\xi_{i}$ are symmetric, we need to estimate
$\left\|\frac{1}{n}\sum_{i=1}^{n}(E[\psi_{\tau}(\xi_{i})])X_{i}\right\|$ $=|E[\psi_{\tau}(\xi_{1})]|\left\|\dfrac{1}{n}\sum_{i=1}^{n}X_{i}\right\|$
$\leq|E[\psi_{\tau}(\xi_{1})]|\left(\left\|\dfrac{1}{n}\sum_{i=1}^{n}(X_{i}-E[X_{i}])\right\|+\|E[X_{1}]\|\right).$
By Cauchy-Schwarz inequality, $|E[\psi_{\tau}(\xi_{i})]|\leq\sqrt{E[4\xi_{i}^{2}]}\sqrt{P(|\xi_{i}|\geq\tau)}\leq 2\sigma\sqrt{\frac{m\log^{4}d}{n}}$. Assuming $P_{jk}=\frac{1}{m_{1}m_{2}}$, we apply Theorem 7.3.1 in[Tropp et al., 2015] to $\frac{1}{n}\sum_{i=1}^{n}(X_{i}-E[X_{i}])$. Since $\frac{1}{n}\|(X_{i}-E[X_{i}])\|$ is bounded by $\frac{2}{n}$, we obtain the following holds with probability at least $1-\frac{1}{d}$,
$\left\|\frac{1}{n}\sum_{i=1}^{n}(X_{i}-E[X_{i}])\right\|\lesssim\sqrt{\frac{\log d}{nm}}+\frac{\log d}{n},$
where we use easy estimates $\|E(X_{i}-E[X_{i}])(X_{i}-E[X_{i}])^{T}\|\lesssim\frac{L_{2}}{m}$ and $\|E(X_{i}-E[X_{i}])^{T}(X_{i}-E[X_{i}])\|\lesssim\frac{L_{2}}{m}$. Given $n\geq m\log^{5}d$, we have
$|E[\psi_{\tau}(\xi_{1})]|\left\|\dfrac{1}{n}\sum_{i=1}^{n}(X_{i}-E[X_{i}])\right\|\lesssim\sigma\sqrt{\frac{1}{nm}}.$
All the entries of $E[X_{i}]$ are $\frac{1}{m_{1}m_{2}}$, then we have $\|E[X_{i}]\|\leq\frac{1}{\sqrt{m_{1}m_{2}}}$. Note $\sqrt{\frac{m\log^{4}d}{n}}\cdot\frac{1}{\sqrt{m_{1}m_{2}}}\leq\sqrt{\frac{1}{nm}}$ holds only when $M\geq m\log^{4}d$. ∎
Proof Lemma 16.
As a solution of the convex program (10), by (41), $\hat{A}_{H}$ satisfies:
$\langle\nabla L(\hat{A}_{H})-L(A_{0}),\hat{A}_{H}-A_{0}\rangle\leq\frac{4}{3}\|\hat{A}_{H}-A_{0}\|_{1}.$
By Lemma 22,
$\|\hat{A}_{H}-A_{0}\|_{1}\leq 3\|P_{A_{0}}(\hat{A}_{H}-A_{0})\|_{1}\leq 3\sqrt{r}\|\hat{A}_{H}-A_{0}\|_{F}.$
By Lemma 20, assuming $\|\hat{A}_{H}-A_{0}\|^{2}_{w(F)}\leq s^{2}$, there is a constant $c_{3}$ such that the following holds with probability at least $1-\frac{1}{d}$,
$\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\left(\frac{9}{10}-\dfrac{4\sigma^{2}}{\tau^{2}}\right)\leq 4\sqrt{r}\lambda\|\hat{A}_{H}-A_{0}\|_{F}+\dfrac{64a^{2}s^{2}}{\tau^{2}}$
$+c_{3}\Bigg{(}\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{1}E\|R\|+\frac{\mu a^{2}\log d}{n}\Bigg{)},$ (31)
where we use $\|\hat{A}_{H}-A_{0}\|_{\infty}\leq 2a$. On the RHS of (31), we have
$\lambda\sqrt{r}\|\hat{A}_{H}-A_{0}\|_{F}\leq 40\mu\lambda^{2}rm_{1}m_{2}+\frac{1}{40}\frac{\|\hat{A}_{H}-A_{0}\|^{2}_{F}}{\mu m_{1}m_{2}}.$
By $\|\hat{A}_{H}-A_{0}\|_{\infty}\leq 2a$, we get
$\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{1}E\|R\|$ $\leq 2a\sqrt{r}\|\hat{A}_{H}-A_{0}\|_{F}E\|R\|$
$\leq 40c_{3}\mu a^{2}E^{2}\|R\|rm_{1}m_{2}+\frac{1}{10c_{3}}\frac{\|\hat{A}_{H}-A_{0}\|^{2}_{F}}{\mu m_{1}m_{2}}.$
Then we obtain
$\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\left(\frac{9}{10}-\dfrac{4\sigma^{2}}{\tau^{2}}\right)\leq 160\mu\lambda^{2}rm_{1}m_{2}+\frac{1}{5}\frac{\|\hat{A}_{H}-A_{0}\|^{2}_{F}}{\mu m_{1}m_{2}}+\dfrac{64a^{2}s^{2}}{\tau^{2}}$
$+c_{3}\Bigg{(}40c_{3}\mu a^{2}E^{2}\|R\|rm_{1}m_{2}+\frac{4\mu a^{2}\log d}{n}\Bigg{)},$
If $n\geq 160m\log^{4}d$, we have $\frac{4\sigma^{2}}{\tau^{2}}\leq\frac{1}{10}$ and $\frac{64a^{2}}{\tau^{2}}\leq\frac{2}{5}$. Then $\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\left(\frac{9}{10}-\frac{4\sigma^{2}}{\tau^{2}}\right)\geq\frac{4}{5}\|\hat{A}_{H}-A_{0}\|^{2}_{w(F)}$. Then we obtain that there is a constant $c_{4}$ such that the following holds with probability at least $1-\frac{1}{d}$,
$\frac{3}{5}\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\leq c_{4}\left(\mu(\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu\log d}{n}\right)+\frac{2}{5}s^{2}.$
This implies
$\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\leq\frac{5c_{4}}{3}\left(\mu(\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu\log d}{n}\right)+\frac{2}{3}s^{2}.$
Denote $\delta=\frac{5c_{4}}{3}\left(\mu(\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu\log d}{n}\right)$. We take $s^{2}=4\delta$. If $\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}>s^{2}$, we can take $\tilde{A}_{H}=A_{0}+\eta(\hat{A}_{H}-A_{0})$ such that $\|\tilde{A}_{H}-A_{0}\|_{w(F)}=\|\eta(\hat{A}_{H}-A_{0})\|_{w(F)}=s$. Note that $\tilde{A}_{H}$ also satisfies
$\|\tilde{A}_{H}-A_{0}\|_{1}\leq 3\sqrt{r}\|\tilde{A}_{H}-A_{0}\|_{F}.$
By Lemma F.2 in [Fan et al., 2018], we have
$\langle\nabla L(\tilde{A}_{H})-\nabla L(A_{0}),\tilde{A}_{H}-A_{0}\rangle$ $\leq\eta\langle\nabla L(\hat{A}_{H})-\nabla L(A_{0}),\hat{A}_{H}-A_{0}\rangle$
$\leq 4\lambda\|P_{A_{0}}(\tilde{A}-A_{0})\|_{1}$
Then, we can repeat the previous steps and Lemma 20 also applies to $\tilde{A}_{H}-A_{0}$. Finally we get that
$\|\tilde{A}_{H}-A_{0}\|_{w(F)}^{2}\leq\delta+\frac{2}{3}s^{2}\leq\frac{11}{12}s^{2}.$
This contradicts with that $\|\tilde{A}_{H}-A_{0}\|_{w(F)}^{2}=s^{2}$. Thus, we get
$\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\leq\frac{20c_{4}}{3}\left(\mu(\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu\log d}{n}\right)$
and
$\frac{\|\hat{A}_{H}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\lesssim\mu^{2}(\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu^{2}\log d}{n}.$
4.5 Proof of Theorem 7 and 9
Based on [Klopp et al., 2017], (12) and (16) can be summarized in the same form
$(\hat{L},\hat{S})\in\underset{\begin{subarray}{c}\|L\|_{\infty}\leq\mathrm{a}\\ \|S\|_{\infty}\leq\mathrm{a}\end{subarray}}{\arg\min}\left\{\frac{1}{N}\sum_{i=1}^{N}\left(Y_{i}-\left\langle X_{i},L+S\right\rangle\right)^{2}+\lambda_{1}\|L\|_{1}+\lambda_{2}\mathcal{R}(S)\right\}.$
where $\mathcal{R}$ is certain matrix norm for $S$. Denote $\mathcal{R}^{*}(B)=\sup_{\mathcal{R}(A)\leq 1}\langle A,B\rangle$. [Klopp et al., 2017] offers the following general results for robust matrix completion.
Theorem 17.
Let Assumptions 1 and 4 be satisfied. Let $\lambda_{2}\geq 4\left(\mathcal{R}^{*}(\frac{1}{N}\sum_{i\in\Omega}\xi_{i}X_{i})+2a\mathcal{R}^{*}(\frac{1}{N}\sum_{i\in\Omega}X_{i})\right)$, and $\lambda_{1}>4\|\frac{1}{N}\sum_{i\in\Omega}\xi_{i}X_{i}\|$. Then, with probability at least $1-\frac{4.5}{d}$,
$\frac{\left\|\hat{L}-L_{0}\right\|_{2}^{2}}{m_{1}m_{2}}+\frac{\left\|\hat{S}-S_{0}\right\|_{2}^{2}}{m_{1}m_{2}}\leq C\left\{\Psi_{1}+\Psi_{2}+\Psi_{3}\right\},$
where $C$ is an absolute constant, and $\Psi_{1}$ is defined as
$\Psi_{1}=\mu^{2}m_{1}m_{2}r\left(\alpha^{2}\lambda_{1}^{2}+\mathbf{a}^{2}\left(E\left\|\frac{1}{n}\sum_{i\in\Omega}\epsilon_{i}X_{i}\right\|\right)^{2}\right)+\mathbf{a}^{2}\mu\sqrt{\frac{\log(d)}{n}}.$
Here, we do not need to care about the definitions of $\Psi_{2}$ and $\Psi_{3}$ since they are unrelated to $O(\frac{rM}{n})$ part in the final bounds. The term $\frac{rM\log d}{n}$ only results from $\Psi_{1}$. Note $\frac{1}{N}\sum_{i\in\Omega}\xi_{i}X_{i}=\frac{1}{\alpha n}\sum_{i\in\Omega}\xi_{i}X_{i}$. By (22) and (24), we can obtain
$\left\|\dfrac{1}{N}\sum_{i\in\Omega}\xi_{i}X_{i}\right\|\leq C\frac{\sigma}{\alpha}\sqrt{\dfrac{1}{nm}}=C\sigma\sqrt{\frac{1}{\alpha Nm}}$
with probability at least $1-\frac{1}{d}$ and
$E\left\|\dfrac{1}{n}\sum_{i\in\Omega}\epsilon_{i}X_{i}\right\|\leq C\sqrt{\dfrac{1}{nm}}.$
Plug the above two bounds in $\Psi_{1}$, then we remove the $\log d$ factor.
4.6 Proof of restricted strong convexity for $L^{2}$ loss function
In this section, $c_{0}\sim c_{4}$ are positive constants whose values do not change at each occurrence.
Lemma 18.
Recall $\|A\|^{2}_{w(F)}=\sum_{i,j}a_{ij}^{2}P_{ij}$. Assume $n\geq m\log^{6}d\log^{2}n$. For a non-zero matrix $A$, the following inequality holds, uniformly with probability at least $1-\dfrac{1}{d}$,
$\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},A\rangle^{2}\geq\dfrac{3}{4}\|A\|_{w(F)}^{2}-c_{3}\left(\|A\|_{\infty}^{2}\dfrac{\mu\log d}{n}+\|A\|_{\infty}\|A\|_{1}E\|R\|\right),$ (32)
where $c_{3}$ is a universal positive constant and $R$ is defined as
$R=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}.$
Proof of Lemma 18.
At first, we introduce a concentration inequality for empirical processes.
Lemma 19 (Theorem 3.27 in [Wainwright, 2019]).
Let $Z_{i},1\leq i\leq n$, be independent but not necessarily identically distributed random variables taking values in measurable spaces $\mathcal{Z},1\leq i\leq n$. $\mathscr{F}$ is a function class. Consider the random variable
$Z=\sup_{f\in\mathscr{F}}\frac{1}{n}\sum_{i=1}^{n}f(Z_{i}).$
If $\|f\|_{\infty}\leq b$ for all $f\in\mathscr{F}$, we have the following inequality, for $t>0$,
$P(Z\geq c_{0}E[Z]+c_{1}\tilde{\sigma}\sqrt{t}+c_{2}bt)\leq\exp\left(-nt\right)$ (33)
where $\tilde{\sigma}^{2}=\sup_{f\in\mathscr{F}}\frac{1}{n}\sum_{i=1}^{n}Ef^{2}(Z_{i})$ and $c_{0},c_{1},c_{2}$ are some universal constants.
Without loss of generality, we can assume $\|A\|_{w(F)}=1$. For a pair of fixed parameters $(\alpha,\rho)$, define:
$S(\alpha,\rho)=\{A\in\mathbb{R}^{m_{1}\times m_{2}};\|A\|_{w(F)}=1,\|A\|_{\infty}\leq\alpha,\|A\|_{1}\leq\rho\}.$
Introducing the convenient short hand notation $F_{A}(X)=\langle X,A\rangle^{2}$, we define
$Z(\alpha,\rho)=\sup_{A\in S(\alpha,\rho)}\left|\dfrac{1}{n}\sum_{i=1}^{n}F_{A}(X_{i})-\|A\|_{w(F)}^{2}\right|.$
For all $A\in S(\alpha,\rho)$, $|F_{A}(X_{i})|\leq\alpha^{2}$ for all $1\leq i\leq n$. Thus, we obtain
$\operatorname{Var}\left(F_{A}(X_{i})\right)$ $=E[F_{A}(X_{i})^{2}]-\left(EF_{A}(X_{i})\right)^{2}$
$\leq E[F_{A}(X_{i})^{2}]$
$\leq\alpha^{2}\|A\|^{2}_{w(F)}.$
Then we have $\tilde{\sigma}^{2}\leq\alpha^{2}$. Applying the concentration inequality (33) for the empirical processes with $t=\frac{2\mu\log d}{N}$, we get
$P\left(Z(\alpha,\rho)\geq c_{0}E[Z(\alpha,\rho)]+\sqrt{2}c_{1}\alpha\sqrt{\frac{\mu\log d}{n}}+2c_{2}\frac{\alpha^{2}\mu\log d}{n}\right)$ $\leq\exp\left(-2\mu\log d\right).$
Now we estimate $E[Z(\alpha,\rho)]$. By Rademacher symmetrization and Ledoux-Talagrand contraction (see [Wainwright, 2019] Proposition 4.11) , we get
$E[Z(\alpha,\rho)]\leq 2E\sup_{A\in S(\alpha,\rho)}\Big{|}\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\langle X_{i},A\rangle^{2}\Big{|}\leq 4\alpha E\sup_{A\in S(\alpha,\rho)}\Big{|}\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\langle X_{i},A\rangle\Big{|}.$
Since $\|A\|_{1}\leq\rho$,
$E\sup_{A\in S(\alpha,\rho)}\Big{|}\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\langle X_{i},A\rangle\Big{|}\leq\rho E\left\|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|=\rho E\|R\|,$
where we use $|\langle A,B\rangle|\leq\|A\|\|B\|_{1}$. Thus, we have
$P\left(Z(\alpha,\rho)\geq 4c_{0}\alpha\rho E\|R\|+\sqrt{2}c_{1}\alpha\sqrt{\frac{\mu\log d}{n}}+2c_{2}\frac{\mu\alpha^{2}\log d}{n}\right)\leq\frac{1}{d^{2\mu}}.$ (34)
Then we use the peeling argument to prove that the following:
$\left|\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},A\rangle^{2}-\|A\|_{w(F)}^{2}\right|\leq 16c_{0}\|A\|_{\infty}\|A\|_{1}E\|R\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\dfrac{\mu\log d}{n}}+8c_{2}\frac{\mu\|A\|^{2}_{\infty}\log d}{n}$ (35)
holds uniformly with probability at least $1-\frac{1}{d}$. Let $\mathcal{E}$ be the event that the bound (35) is violated for some $A$ with $\|A\|_{w(F)}=1$. Note that under Assumption 2,
$\|A\|_{F}\leq\sqrt{\mu m_{1}m_{2}}\|A\|_{w(F)}=\sqrt{\mu m_{1}m_{2}}.$
and
$\|A\|_{F}\geq\|A\|_{w(F)}=1.$
Then, we have
$\|A\|_{1}\leq\sqrt{m}\|A\|_{F}\leq\sqrt{\mu d^{3}}$
and
$1\leq\|A\|_{F}\leq\|A\|_{1}.$
Without loss of generality, we can assume $\|A\|\in[1,\sqrt{\mu d^{3}}]$. Similarly, we have
$\|A\|_{\infty}\leq\|A\|_{F}\leq\sqrt{\mu m_{1}m_{2}}$
and
$\|A\|_{\infty}\geq\frac{\|A\|_{F}}{\sqrt{m_{1}m_{2}}}\geq\frac{1}{d}.$
For an integer pair $(k,l)$, define
$\tilde{S}(k,l)=\{A;\|A\|_{w(F)}=1,2^{k-1}\leq\|A\|_{\infty}\leq 2^{k},\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ 2^{l-1}\leq\|A\|_{1}\leq 2^{l}\}$
Every matrix $A$ with $\|A\|_{w(F)}=1$ belongs to one $\tilde{S}(k,l)$ with $k=\lfloor\log_{2}\frac{1}{d}\rfloor,\ldots,\lceil\frac{1}{2}\log_{2}(\mu d)\rceil$ and $l=1,\ldots,\lceil\frac{1}{2}\log_{2}(\mu d^{3})\rceil$. Thus the total number of the pairs $(k,l)$ is less than $H^{2}=9\log_{2}^{2}(\mu d)$. Let $\mathcal{E}_{k,l}$ be the event that the bound (35) is violated for some $A\in\tilde{S}_{k,l}$. It can be seen that
$\mathcal{E}\subseteq\bigcup_{1\leq k\leq H\atop 1\leq l\leq H}\mathcal{E}_{kl}.$
Thus, if there exists a matrix $A$ with $\|A\|_{w(F)}=1$ that violates the bound (35), it must belong to some set $\tilde{S}_{k,l}$. Next, for $\rho=2^{l},\alpha=2^{k}$, define the event
~$\widetilde{\mathcal{E}}_{k,l}:=\left\{Z(\alpha,\rho)\geq 4c_{0}\alpha\rho E\|R\|+\sqrt{2}c_{1}\alpha\sqrt{\dfrac{\mu\log d}{n}}+4c_{2}\frac{\mu\alpha^{2}\log d}{n}\right\}.$
We claim that $\mathcal{E}_{k,l}\subseteq\widetilde{\mathcal{E}}_{k,l}$. Indeed, if event $\mathcal{E}_{k,l}$ occurs, then there must exist some $A\in\tilde{S}_{k,l}$ such that
$\left|\frac{1}{n}\sum_{i=1}^{n}\langle X_{i},A\rangle^{2}-\|A\|_{w(F)}^{2}\right|$ $\geq 16c_{0}\|A\|_{\infty}\|A\|_{1}E\|R\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\dfrac{\mu\log d}{n}}+16c_{2}\dfrac{\|A\|^{2}_{\infty}\mu\log d}{n}$
$\geq 16c_{0}2^{k-1}2^{l-1}E\|R\|+2\sqrt{2}c_{1}2^{k-1}\sqrt{\dfrac{\mu\log d}{n}}+16c_{2}2^{2(k-1)}\frac{\mu\log d}{n}$
$\geq\tilde{4}c_{0}2^{k}2^{l}E\|R\|+\sqrt{2}c_{1}2^{k}\sqrt{\dfrac{\mu\log d}{n}}+4c_{2}2^{2k}\frac{\mu\log d}{n},$
showing that ~$\widetilde{\mathcal{E}}_{k,l}$ occurs. Putting together the pieces, we have, for $d$ sufficiently large,
$P(\mathcal{E})\stackrel{{\scriptstyle\text{ (i) }}}{{\leq}}\sum_{1\leq k\leq H\atop 1\leq l\leq H}P\left(\widetilde{\mathcal{E}}_{k,l}\right)\stackrel{{\scriptstyle\text{ (ii) }}}{{\leq}}\frac{H^{2}}{d^{2\mu}}\leq\frac{9\log^{2}\mu d}{d^{\mu}}\cdot\frac{1}{d^{\mu}}\leq\dfrac{1}{d},$
where inequality (i) follows from the union bound applied to the inclusion $\mathcal{E}\subseteq\bigcup_{k,\ell=1}^{H}\widetilde{\mathcal{E}}_{k,l}$; inequality (ii) is a consequence of the tail bound (34). So far, we have proved (35).
In the above argument, we normalize the matrix by its $\|\cdot\|_{w(F)}$ norm. Namely we consider $\frac{A}{\|A\|_{w(F)}}$, whose infinity norm and nuclear norm in (35) are $\dfrac{\|A\|_{\infty}}{\|A\|_{w(F)}}$ and $\dfrac{\|A\|_{1}}{\|A\|_{w(F)}}$. Multiplying $\|A\|_{w(F)}$ on both sides of formula (35), we obtain that with probability at least $1-\frac{1}{d}$,
$\left|\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},A\rangle^{2}-\|A\|_{w(F)}^{2}\right|$ $\leq 16c_{0}\|A\|_{\infty}\|A\|_{1}\sqrt{\dfrac{1}{nm}}+2\sqrt{2}c_{1}\|A\|_{\infty}\|A\|_{w(F)}\sqrt{\dfrac{\mu\log d}{n}}$
$+16c_{2}\|A\|^{2}_{\infty}\frac{\mu\log d}{n}.$
which implies
$\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},A\rangle^{2}$ $\geq\|A\|_{w(F)}^{2}-16c_{0}\|A\|_{\infty}\|A\|_{1}E\|R\|-\frac{\|A\|_{w(F)}^{2}}{4}-(8c_{1}^{2}+4c_{2})\|A\|^{2}_{\infty}\frac{\mu\log d}{n}$
$=\frac{3}{4}\|A\|_{w(F)}^{2}-16c_{0}\|A\|_{\infty}\|A\|_{1}E\|R\|-(8c_{1}^{2}+4c_{2})\|A\|^{2}_{\infty}\frac{\mu\log d}{n}.$
4.7 Proof of restricted strong convexity for Huber loss function
In this section, $c_{0}\sim c_{3}$ are positive constants whose values do not change at each occurrence, $L(A)=\frac{1}{n}\sum_{i=1}^{n}l_{\tau}(Y_{i}-\langle X_{i},A\rangle)$.
Lemma 20.
Assume $n\geq m\log^{4}d$ and $\|\hat{A}_{H}-A_{0}\|_{w(F)}\leq s$. With probability at least $1-\frac{1}{d}$, we have
$\langle\nabla L(\hat{A}_{H})-\nabla L(A_{0}),\hat{A}_{H}-A_{0}\rangle$ $\geq\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\left(\frac{9}{10}-\dfrac{4\sigma^{2}}{\tau^{2}}\right)-\dfrac{16\|\hat{A}_{H}-A_{0}\|^{2}_{\infty}s^{2}}{\tau^{2}}$
$-c_{3}\Bigg{(}\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{1}E\|R\|+\frac{\mu\|\hat{A}_{H}-A_{0}\|_{\infty}^{2}\log d}{n}\Bigg{)},$ (36)
where $c_{3}$ is a universal constant.
Proof of Lemma 20.
Consider events $H_{i}=\{|\langle X_{i},\hat{A}_{H}-A_{0}\rangle|\leq\tau/2\}\cap\{|\xi_{i}|\leq\tau/2\}$. We have the following inequality:
$\quad\langle\nabla L(\hat{A}_{H})-\nabla L(A_{0}),\hat{A}_{H}-A_{0}\rangle$
$=\frac{1}{n}\sum_{i=1}^{n}(\psi_{\tau}(Y_{i}-\langle X_{i},A_{0}\rangle)-\psi_{\tau}(Y_{i}-\langle X_{i},\hat{A}_{H}\rangle))\langle X_{i},\hat{A}_{H}-A_{0}\rangle$
$\geq\frac{1}{n}\sum_{i=1}^{n}(\psi_{\tau}(Y_{i}-\langle X_{i},A_{0}\rangle)-\psi_{\tau}(Y_{i}-\langle X_{i},\hat{A}_{H}\rangle))\langle X_{i},\hat{A}_{H}-A_{0}\rangle\textbf{1}_{H_{i}}$
$=\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}_{H}-A_{0}\rangle^{2}\textbf{1}_{H_{i}}.$
Consider function $G_{R}(x):\mathbb{R}\rightarrow\mathbb{R}$,
$G_{R}(x)=\begin{cases}x^{2},|x|\leq R,\\ (x-2R\cdot\operatorname{sign}(x))^{2},R\leq|x|\leq 2R,\\ 0,\text{otherwise}.\end{cases}$
$G_{R}(x)$ has the following properties:
  • 1.
    $x^{2}\textbf{1}_{|x|\leq 2R}\geq G_{R}(x)\geq x^{2}\textbf{1}_{|x|\leq R}$.
  • 2.
    $G_{R}(x)$ is Lipschitz with $|G_{R}(x)-G_{R}(y)|\leq 2R|x-y|$ since $|G_{R}^{\prime}(x)|\leq 2R$. Moreover, if we let $|x|,|y|\leq a$, $|G_{R}(x)-G_{R}(y)|\leq 2a|x-y|$.
Assume $\|\hat{A}_{H}-A_{0}\|_{w(F)}\leq s$. Then by Property 1 of $G_{R}(\cdot)$,
$\quad\leavevmode\nobreak\ \left\langle X_{i},\dfrac{\hat{A}_{H}-A_{0}}{\|\hat{A}_{H}-A_{0}\|_{w(F)}}\right\rangle^{2}\textbf{1}_{H_{i}}$
$=\left\langle X_{i},\dfrac{\hat{A}_{H}-A_{0}}{\|\hat{A}_{H}-A_{0}\|_{w(F)}}\right\rangle^{2}\textbf{1}\left(\dfrac{\langle X_{i},\hat{A}_{H}-A_{0}\rangle}{\|\hat{A}_{H}-A_{0}\|_{w(F)}}\leq\dfrac{\tau}{2\|\hat{A}_{H}-A_{0}\|_{w(F)}}\right)\textbf{1}\left(|\xi_{i}|\leq\tau/2\right)$
$\geq\left\langle X_{i},\dfrac{\hat{A}_{H}-A_{0}}{\|\hat{A}_{H}-A_{0}\|_{w(F)}}\right\rangle^{2}\textbf{1}\left(\dfrac{\langle X_{i},\hat{A}_{H}-A_{0}\rangle}{\|\hat{A}_{H}-A_{0}\|_{w(F)}}\leq\dfrac{\tau}{2s}\right)\textbf{1}_{\chi_{i}}$
$\geq G_{\frac{\tau}{4s}}\left(\left\langle X_{i},\frac{\hat{A}_{H}-A_{0}}{\|\hat{A}_{H}-A_{0}\|_{w(F)}}\right\rangle\right)\textbf{1}_{\chi_{i}}.$
where $\chi_{i}=\{|\xi_{i}|\leq\tau/2\}$. Hence, we have
$\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}_{H}-A_{0}\rangle^{2}\textbf{1}_{H_{i}}$ $=\dfrac{\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}}{n}\sum_{i=1}^{n}\left\langle X_{i},\dfrac{\hat{A}_{H}-A_{0}}{\|\hat{A}_{H}-A_{0}\|_{w(F)}}\right\rangle^{2}\textbf{1}_{H_{i}}$
$\geq\|\hat{A}_{H}-A_{0}\|^{2}_{w(F)}\cdot\left(\dfrac{1}{n}\sum_{i=1}^{n}G_{\frac{\tau}{4s}}\left(\left\langle X_{i},\dfrac{\hat{A}_{H}-A_{0}}{\|\hat{A}_{H}-A_{0}\|_{w(F)}}\right\rangle\right)\textbf{1}_{\chi_{i}}\right).$
For a matrix $A\in\mathbb{R}^{m_{1}\times m_{2}}$, we define
$f(A):=E_{X_{i},\xi_{i}}[G_{\frac{\tau}{4s}}(X_{i},A)\textbf{1}_{\chi_{i}}].$
We have,
$f(A)$ $\geq E[\langle X_{i},A\rangle^{2}]\textbf{1}\{|\langle X_{i},A\rangle|\leq\tau/4s\}\textbf{1}_{\chi_{i}}$
$\geq E\langle X_{i},A\rangle^{2}-E\langle X_{i},A\rangle^{2}\textbf{1}\{|\langle X_{i},A\rangle|>\tau/4s\}-E\langle X_{i},A\rangle^{2}\textbf{1}\{|\xi_{i}|>\tau/2\}$
$\geq\|A\|_{w(F)}^{2}-\dfrac{16s^{2}\|A\|^{2}_{\infty}}{\tau^{2}}\|A\|_{w(F)}^{2}-\dfrac{4\sigma^{2}}{\tau^{2}}\|A\|_{w(F)}^{2}$
$=\|A\|_{w(F)}^{2}\left(1-\dfrac{16s^{2}\|A\|_{\infty}^{2}}{\tau^{2}}-\dfrac{4\sigma^{2}}{\tau^{2}}\right).$
Thus
$\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}f\left(\dfrac{\hat{A}_{H}-A_{0}}{\|\hat{A}_{H}-A_{0}\|_{w(F)}}\right)\geq\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}(1-\dfrac{4\sigma^{2}}{\tau^{2}})-\dfrac{16\|\hat{A}_{H}-A_{0}\|^{2}_{\infty}s^{2}}{\tau^{2}}.$ (37)
Now we introduce a useful lemma.
Lemma 21.
Assume $\|A\|_{w(F)}^{2}=1$. Then there are universal constants $c_{0},c_{1},c_{2}$ such that
$\left|\dfrac{1}{n}\sum_{i=1}^{n}G_{\frac{\tau}{4s}}(\langle X_{i},A\rangle)-f(A)\right|\leq 8c_{0}\|A\|_{\infty}\|A\|_{1}E\|R\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\frac{\mu\log d}{n}}+8c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}$ (38)
for all such $A$, uniformly with probability at least $1-\frac{1}{d}$. Here $R=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}$ and $\chi_{i}=\{|\xi_{i}|\leq\tau/2\}$.
By (37) and (38), we get
$\quad\langle\nabla L(\hat{A}_{H})-\nabla L(A_{0}),\hat{A}_{H}-A_{0}\rangle\geq\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}(1-\dfrac{4\sigma^{2}}{\tau^{2}})-\dfrac{16a^{2}s^{2}}{\tau^{2}}$
$-\Bigg{(}8c_{0}\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{1}E\|R\|+2\sqrt{2}c_{1}\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{w(F)}\sqrt{\frac{\mu\log d}{n}}$
$+8c_{2}\frac{\mu\|\hat{A}_{H}-A_{0}\|_{\infty}^{2}\log d}{n}\Bigg{)}$
$\geq\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}(1-\dfrac{4\sigma^{2}}{\tau^{2}})-\dfrac{16\|\hat{A}_{H}-A_{0}\|^{2}_{\infty}s^{2}}{\tau^{2}}$
$-c_{3}\Bigg{(}\|\hat{A}_{H}-A_{0}\|_{1}\|\hat{A}_{H}-A_{0}\|_{\infty}+\frac{\mu\|\hat{A}_{H}-A_{0}\|^{2}_{\infty}\log d}{n}\Bigg{)}-\frac{\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}}{10}$ (39)
where in the last step, we apply the basic inequality ($2ab\leq\frac{a^{2}}{c^{2}}+c^{2}b^{2}$) and choose a suitable constant $c_{3}$. By (39), we get (36).
Proof of Lemma 21.
We follow a similar procedure as in the proof of Lemma 18. For fixed $(\alpha,\rho)\in\mathbb{R}^{2}$, define the space:
$S(\alpha,\rho)=\{A:\|A\|_{w(F)}^{2}=1,\|A\|_{1}\leq\rho,\|A\|_{\infty}\leq\alpha\}.$
Denote
$F_{A}(X_{i},\xi_{i})=G_{\frac{\tau}{4s}}(\langle X_{i},A\rangle)\textbf{1}_{\chi_{i}}.$
for $A\in S(\alpha,\rho)$. Recall $f(A)=E_{X_{i},\xi_{i}}[F_{A}(X_{i},\xi_{i})]$. Since $x^{2}\geq x^{2}\textbf{1}_{|x|\leq 2R}\geq G_{R}(x)$, for $A\in S(\alpha,\rho)$,
$\quad\leavevmode\nobreak\ E[F^{2}_{A}(X_{i},\xi_{i})]-f^{2}(A)$
$\leq E[F^{2}_{A}(X_{i},\xi_{i})]$
$\leq E[G^{2}_{\tau/4s}(\langle X_{i},A\rangle)]$
$\leq\alpha^{2}\|A\|_{w(F)}^{2}$
$\leq\alpha^{2},$
which implies $\tilde{\sigma}\leq\alpha$ in (33). Consider
$Z(\alpha,\rho)=\sup_{A\in S(\alpha,\rho)}\left|\dfrac{1}{n}\sum_{i=1}^{n}F_{A}(X_{i},\xi_{i})-f(A)\right|$
We have
$E[Z(\alpha,\rho)]$ $\stackrel{{\scriptstyle\text{i}}}{{\leq}}E\sup_{A\in\Omega(\alpha,\rho)}\left|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}F_{A}(X_{i},\xi_{i})\right|$
$\stackrel{{\scriptstyle\text{ii}}}{{\leq}}2\alpha E\sup_{A\in\Omega(\alpha,\rho)}\left|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\langle X_{i}\textbf{1}_{\chi_{i}},A\rangle\right|$
$\stackrel{{\scriptstyle\text{iii}}}{{\leq}}4\alpha\rho E\left\|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}\right\|.$
In step (i) we use symmetrization technique. By Property (ii) of $G_{R}(\cdot)$ and $\|A\|_{\infty}\leq\alpha$, $G_{\frac{\tau}{4s}}(\cdot)$ is Lipschitz with Lipschitz constant less than $2\alpha$. Then, in step (ii) we apply Talagrand-Ledoux contraction inequality. Step (iii) is due to $|\langle A,B\rangle|\leq\|A\|\|B\|_{1}$.
We take $t=\frac{2\mu\log d}{n}$ in (33). Since $\tilde{\sigma}\leq\alpha$ and $\|F_{A}(X_{i},\xi_{i})\|_{\infty}\leq\alpha^{2}$, with probability at least $1-\exp(-2\mu\log d)$,
$Z(\alpha,\rho)\leq 4c_{0}\alpha\rho E\|R\|+\sqrt{2}c_{1}\alpha\sqrt{\frac{\mu\log d}{n}}+4c_{2}\frac{\mu\alpha^{2}\log d}{n}.$ (40)
Note that under Assumption 2,
$\|A\|_{F}\leq\sqrt{\mu m_{1}m_{2}}\|A\|_{w(F)}=\sqrt{\mu m_{1}m_{2}}.$
and
$\|A\|_{F}\geq\|A\|_{w(F)}=1.$
Then, we have
$\|A\|_{1}\leq\sqrt{m}\|A\|_{F}\leq\sqrt{\mu d^{3}}$
and
$1\leq\|A\|_{F}\leq\|A\|_{1}.$
Without loss of generality, we can assume $\|A\|_{1}\in[1,\sqrt{\mu d^{3}}]$. Similarly, we have
$\|A\|_{\infty}\leq\|A\|_{F}\leq\sqrt{\mu m_{1}m_{2}}$
and
$\|A\|_{\infty}\geq\frac{\|A\|_{F}}{\sqrt{m_{1}m_{2}}}\geq\frac{1}{d}.$
Consider
$\tilde{S}(k,l)=\{A;\|A\|_{w(F)}=1,2^{k-1}\leq\|A\|_{\infty}\leq 2^{k},\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ 2^{l-1}\leq\|A\|_{1}\leq 2^{l}\}$
Every matrix $A$ with $\|A\|_{w(F)}=1$ belongs to some $\tilde{S}(k,l)$ with $k=\lfloor\log_{2}\frac{1}{d}\rfloor,\ldots,\lceil\frac{1}{2}\log_{2}(\mu d^{2})\rceil$ and $l=1,\ldots,\lceil\log_{2}\frac{1}{2}(\mu d^{3})\rceil$. Thus the total number of the pairs $(k,l)$ is less than $H^{2}=9\log_{2}^{2}(\mu d)$.
Denote the events $\mathcal{E}$ that a matrix $A$ violates the following relation
$\left|\dfrac{1}{n}\sum_{i=1}^{n}F_{A}(X_{i},\xi_{i})-f(A)\right|\leq 16c_{0}\|A\|_{\infty}\|A\|_{1}E\|R\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\frac{\mu\log d}{n}}+16c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}$
and denote $\mathcal{E}_{k,l}$ the event that there exists a matrix in $\tilde{S}(k,l)$ violating the following relation
$\left|\dfrac{1}{n}\sum_{i=1}^{n}F_{A}(X_{i},\xi_{i})-f(A)\right|\leq 4c_{0}2^{k}2^{l}E\|R\|+\sqrt{2}c_{1}2^{k}\sqrt{\frac{\mu\log d}{n}}+4c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}.$
$\mathcal{E}$ is contained in the union $\cup_{k,l}\mathcal{E}_{k,l}$. Indeed, letting $\alpha=2^{k}$ and $\rho=2^{l}$, for a matrix $A$ in $\tilde{S}(k,l)$, we have
$\left|\dfrac{1}{n}\sum_{i=1}^{n}F_{A}(X_{i},\xi_{i})-f(A)\right|$ $\geq 16c_{0}\|A\|_{\infty}\|A\|_{1}E\|R\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\frac{\mu\log d}{n}}+16c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}$
$\geq 16c_{0}2^{k-1}2^{l-1}E\|R\|+2\sqrt{2}c_{1}2^{k-1}\sqrt{\frac{\mu\log d}{n}}+16c_{2}2^{2(k-1)}\frac{\mu\log d}{n}$
$=4c_{0}\alpha\rho+\sqrt{2}\alpha\sqrt{\frac{\mu\log d}{n}}+4c_{2}\alpha^{2}\frac{\mu\log d}{n}.$
Thus, by (40),
$P(\mathcal{E})\leq P(\cup_{k,l}\mathcal{E}_{k,l})\leq H^{2}\exp(-2\mu\log d)\leq\frac{9\log_{2}^{2}\mu d}{d^{\mu}}\cdot\frac{1}{d^{\mu}}\leq\frac{1}{d}$
for sufficiently large $d$. ∎
4.8 Other technical lemmas
Lemma 22.
Recall the definitions of $P_{A_{0}}(\cdot)$ and $P_{A_{0}}^{\perp}(\cdot)$ defined in Section 1. Consider the following estimator:
$\bar{A}=\operatorname*{arg\,min}_{A\in\Omega}\left\{L(A)+\lambda\|A\|_{1}\right\},$
where $\Omega$ is the domain of $A$ and $L(A)$ is a convex loss function. If we take $\lambda\geq 3\|\nabla L(A_{0})\|$, then $\bar{A}-A_{0}$ satisfies
$\|P_{A_{0}}^{\perp}(\bar{A}-A_{0})\|_{1}\leq 2\|P_{A_{0}}(\bar{A}-A_{0})\|_{1}.$
As special cases, we have the following inequalities. Take $\lambda\geq 3\|\frac{2}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\|$ in (2), then
$\|P_{A_{0}}^{\perp}(\hat{A}-A_{0})\|_{1}\leq 2\|P_{A_{0}}(\hat{A}-A_{0})\|_{1}.$
Take $\lambda\geq 3\frac{\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\|}{\sqrt{\frac{1}{n}\sum_{i=1}^{n}\xi_{i}^{2}}}$ in (8) , then
$\|P_{A_{0}}^{\perp}(\hat{A}_{S}-A_{0})\|_{1}\leq 2\|P_{A_{0}}(\hat{A}_{S}-A_{0})\|_{1}.$
Take $\lambda\geq 3\|\frac{1}{n}\sum_{i=1}^{n}\psi_{\tau}(\xi_{i})X_{i}\|$ in (10) , then
$\|P_{A_{0}}^{\perp}(\hat{A}_{H}-A_{0})\|_{1}\leq 2\|P_{A_{0}}(\hat{A}_{H}-A_{0})\|_{1}.$
Proof.
As a solution of a convex program, by Proposition 1.3 in [Bubeck, 2014] , $\bar{A}$ satisfies:
$\langle\nabla L(\bar{A})+\lambda Z,\bar{A}-A_{0}\rangle\leq 0.$
where $Z\in\partial\|\bar{A}\|_{1}$. Since $\|\cdot\|_{1}$ is convex,
$\langle Z,A_{0}-\bar{A}\rangle\leq\|A_{0}\|_{1}-\|\bar{A}\|_{1}.$
By the convexity of $L(\cdot)$,
$\langle\nabla L(\bar{A})-\nabla L(A_{0}),\bar{A}-A_{0}\rangle\geq 0.$
Combining the results above, we obtain:
$0\leq\langle\nabla L(\bar{A})-\nabla L(A_{0}),\bar{A}-A_{0}\rangle\leq\lambda(\|A_{0}\|_{1}-\|\bar{A}\|_{1})+\|\nabla L(A_{0})\|\|\bar{A}-A_{0}\|_{1}.$ (41)
Note
$\|\bar{A}\|_{1}$ $=\|P_{A_{0}}(\bar{A}-A_{0})+P_{A_{0}}^{\perp}(\bar{A}-A_{0})+P_{A_{0}}(A_{0})\|_{1}$
$\geq\|P_{A_{0}}^{\perp}(\bar{A}-A_{0})\|_{1}+\|P_{A_{0}}(A_{0})\|_{1}-\|P_{A_{0}}(\bar{A}-A_{0})\|_{1}.$
Then, we get
$0\leq\frac{4}{3}\|P_{A_{0}}(\bar{A}-A_{0})\|_{1}-\frac{2}{3}\|P_{A_{0}}^{\perp}(\bar{A}-A_{0})\|_{1}.$
For (2), $L(A)=\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}$ and $\nabla L(A_{0})=\frac{2}{n}\sum_{i=1}^{n}\xi_{i}X_{i}$. For (8), $L(A)=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}}$ and $\nabla L(A_{0})=\frac{\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}}{L(A_{0})}=\frac{\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}}{\sqrt{\frac{1}{n}\sum_{i=1}^{n}\xi_{i}^{2}}}$. For (10), $L(A)=\frac{1}{n}\sum_{i=1}^{n}l_{\tau}(Y_{i}-\langle X_{i},A\rangle)$, and $\nabla L(A_{0})=\frac{1}{n}\sum_{i=1}^{n}\psi_{\tau}(\xi_{i})X_{i}$. ∎

### v3

@@
 This dimensional factor represents a significant gap between the upper bound and the minimax lower bound, especially in high dimension.
-Through a more precise spectral norm analysis, we remove the dimensional factors for five different estimators in various settings, thereby establishing their minimax rate optimality.
+Through a more precise spectral norm analysis, we remove the dimensional factors for three popular matrix completion estimators, thereby establishing their minimax rate optimality.
 1 Introduction
@@
 [2024].
-In particular, many studies focus on the convergence rate of the nuclear norm penalization methods [Koltchinskii et al., 2011, Negahban and Wainwright, 2012, Klopp, 2014, Yu et al., 2024].
+In particular, many studies focus on the convergence rate of the nuclear norm penalization methods Koltchinskii et al.
+[2011], Negahban and Wainwright [2012], Klopp [2014], Yu et al.
+[2024].
 However, the minimax lower bound, obtained in Koltchinskii et al.
@@
 This discrepancy between the two bounds becomes more pronounced in high-dimensional settings, highlighting a potential gap in the existing theoretical understanding.
-Due to this issue, previous works have had to qualify their results by stating “our estimator is minimax optimal, up to a logarithmic factor”.
-In this paper, we use a class of sharp matrix concentration inequalities to remove the logarithmic dimension factor.
+Due to the ubiquitousness of this issue, all the previous works have had to qualify their results by stating “our estimator is minimax optimal, up to a logarithmic factor”.
+In this paper, we leverage a class of sharp matrix concentration inequalities to remove the logarithmic dimension factor.
 After our work, the qualification will no longer be necessary.
 Problem setup
-Let $A_{0}\in\mathbb{R}^{m_{1}\times m_{2}}$ be the unknown matrix we aim to recover. We observe independent samples $\{(X_{i},Y_{i})\}_{i=1}^{\nu}$ from the model
-$Y_{i}=\langle X_{i},A_{0}\rangle+\xi_{i},\quad i=1,\ldots,\nu,$ (1)
+Let $A_{0}\in\mathbb{R}^{m_{1}\times m_{2}}$ be the unknown matrix we aim to recover. We observe independent samples $\{(X_{i},Y_{i})\}_{i=1}^{n}$ from the model
+$Y_{i}=\langle X_{i},A_{0}\rangle+\xi_{i},\quad i=1,\ldots,n,$ (1)
 where $X_{i}\in\mathbb{R}^{m_{1}\times m_{2}}$, $\langle X_{i},A_{0}\rangle=\text{tr}(X_{i}^{T}A_{0})$, and $\text{tr}(B)$ denotes the trace of matrix $B$.
-The noise variables $\xi_{i}$ are independent and identically distributed.
-The sampling matrices $X_{i}\in\mathbb{R}^{m_{1}\times m_{2}}$ take values from $\{e_{l}(m_{1})e_{k}^{T}(m_{2}),1\leq l\leq m_{1},1\leq k\leq m_{2}\}$, where $e_{j}(m_{1}),e_{k}(m_{2})$ denote standard basis vectors in $\mathbb{R}^{m_{1}},\mathbb{R}^{m_{2}}$, respectively.
+The sampling matrices $X_{i}\in\mathbb{R}^{m_{1}\times m_{2}}$ take values from $\{e_{j}(m_{1})e_{k}^{T}(m_{2}),1\leq j\leq m_{1},1\leq k\leq m_{2}\}$, where $\{e_{j}(m_{1})\}_{j=1}^{m_{1}},\{e_{k}(m_{2})\}_{k=1}^{m_{2}}$ denote standard basis vectors in $\mathbb{R}^{m_{1}},\mathbb{R}^{m_{2}}$, respectively.
 This implies that each $X_{i}$ only has one non-zero entry, indicating the location of the sampled entry.
 The distribution of the noise variables $\xi_{i}$ will be specified in different settings. The sampling distribution of $X_{i}$ is denoted by
-$P_{jk}=P(X_{i}=e_{j}(m_{1})e_{k}^{T}(m_{2})).$
-Additionally, $X_{i}$ and $\xi_{i}$ are assumed to be independent.
-In this paper, we consider that $\text{rank}(A_{0})\leq r$ or $A_{0}$ is the sum of a low rank matrix and a sparse matrix, and $\|A_{0}\|_{\infty}\leq a$ where $\|A\|_{\infty}=\max_{1\leq i\leq m_{1},1\leq j\leq m_{2}}|a_{ij}|$ is the entrywise maximum norm of $A=(a_{ij})$.
-The total sample size $\nu=n{\rm\leavevmode\nobreak\ or\leavevmode\nobreak\ }N$ depending on the settings.
+$P_{jk}=\mathbb{P}(X_{i}=e_{j}(m_{1})e_{k}^{T}(m_{2})),\quad 1\leq j\leq m_{1},1\leq k\leq m_{2}.$
+$X_{i}$ and $\xi_{i}$ can be dependent.
+Throughout the paper, we consider that $\text{rank}(A_{0})\leq r$ and $\|A_{0}\|_{\infty}\leq a$, where $\|A\|_{\infty}=\max_{1\leq j\leq m_{1},1\leq k\leq m_{2}}|a_{jk}|$ is the entrywise maximum norm of $A=(a_{jk})$.
 The logarithmic dimension factor issue
@@
 However, the rank penalty is non-convex and makes the corresponding program difficult to solve.
-The nuclear norm $\|\cdot\|_{1}$, namely the sum of all the singular values of a matrix, is usually used as a convex relaxation for the rank.
+The nuclear norm $\|\cdot\|_{*}$, namely the sum of all the singular values of a matrix, is usually used as a convex relaxation for the rank.
 There are many classical works on nuclear norm penalization methods under different assumptions.
 Assuming the noise has exponential decay, Koltchinskii et al.
-[2011] consider the case that the sampling distribution is known; Negahban and Wainwright [2012] study the situation where the sampling distribution is unknown and the row indices and the column indices of the observations are sampled independently; Klopp [2014] explores the problem for general sampling distributions.
+[2011] consider the case that the sampling distribution is known; Negahban and Wainwright [2012] study the situation where the row indices and the column indices of the observations are sampled independently and the penalty is a weighte nuclear norm; Klopp [2014] explores the problem for general sampling distributions.
 Consider the classical nuclear norm penalized least squares estimator
-$\hat{A}=\operatorname*{arg\,min}_{\|A\|_{\infty}\leq a}\left\{\dfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}+\lambda\|A\|_{1}\right\}.$
+$\hat{A}=\operatorname*{argmin}_{\|A\|_{\infty}\leq a}\left\{\dfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}+\lambda\|A\|_{*}\right\}.$
 Assuming $\text{rank}(A_{0})\leq r$, $X_{i}$ has a general non-uniform distribution and the noise has exponential decay, Klopp [2014] obtains the following bound that holds with high probability:
-$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\apprle\log(m_{1}+m_{2})\dfrac{r\max(m_{1},m_{2})}{n},$
+$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\apprle\log(m_{1}+m_{2})\dfrac{r\max(m_{1},m_{2})}{n},$ (2)
 […]
-$\inf_{\hat{A}}\sup_{\mathcal{A}}P\left(\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\apprge\dfrac{r\max(m_{1},m_{2})}{n}\right)\geq C,$
+$\inf_{\hat{A}}\sup_{\mathcal{A}}\mathbb{P}\left(\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\apprge\dfrac{r\max(m_{1},m_{2})}{n}\right)\geq C,$
 where $C\in(0,1)$ is a constant, and the infimum and supremum are taken over all the possible estimators and the parameter space $\{A_{0}\in\mathbb{R}^{m_{1}\times m_{2}}:\text{rank}(A_{0})\leq r,\|A_{0}\|_{\infty}\leq a\}$ respectively.
@@
 [2024] for matrix completion with heavy tailed noise.
+By our analysis, this gap arises essentially from the spectral norm analysis of some random matrices.
+In matrix completion, we frequently need to control the spectral norms of random matrices of the form $\frac{1}{n}\sum_{i=1}^{n}\zeta_{i}X_{i}$, where $\zeta_{i}$ are bounded or light-tailed random variables.
+The use of traditional concentration inequalities in this context typically introduces a logarithmic dependence on the dimension.
+This limitation is not unique to matrix completion—it arises in a variety of statistical and mathematical problems and has garnered significant attention in recent years.
+Notably, Bandeira et al.
+[2023] and Brailovskaya and van Handel [2024] proposed sharp new concentration inequalities that can successfully remove the logarithmic dependence in various settings.
+These results will serve as key tools in our paper.
+Another point worth mentioning is that there is an alternative sampling model in matrix completion.
+In this model, there are Bernoulli random variables $\eta_{ij}$ indicating whether the entry of $A_{0}$ at location $(i,j)$ is observed.
+Under this assumption, the observations can be written as:
+[…]
+where $\xi_{ij}$ represents the noise at location $(i,j)$ and $A_{0}=(a_{0,ij})$.
+This sampling model is often referred to as sampling without replacement because each entry can be selected at most once.
+A rich body of literature has investigated this model [Candes and Plan, 2010, Chen and Wainwright, 2015, Candes and Recht, 2012, Keshavan et al., 2009, Chen and Wainwright, 2015, Chen et al., 2020].
+As a contrast, the sampling strategy considered in our paper is known as sampling with replacement, as it allows repeated observations of the same entry.
+Since $n$ is often assumed to be much smaller than the matrix dimension $m_{1}\times m_{2}$, the number of repeated observations could be very small.
+Despite different sampling strategies, the theoretical results derived in both models are often compared side by side [Chen and Wainwright, 2015, Chen et al., 2020].
+For instance, some works derive bounds of the following form [Chen and Wainwright, 2015]:
+$\quad\frac{\|\widehat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\lesssim\frac{r\max(m_{1},m_{2})}{m_{1}m_{2}p},$ (3)
+under the non-replacement model with $\eta_{ij}\sim\text{Bernoulli}(p)$.
+Here, as the expected sample size, $m_{1}m_{2}p$ plays a role as $n$ in the with-replacement model.
+Compared to (3), the extra $\log(m_{1}+m_{2})$ factor in (2) is a common critique of algorithms developed under the with-replacement model[Chen and Wainwright, 2015, Chen et al., 2020].
+Therefore, removing this logarithmic factor can help significantly defend the theoretical validity of the algorithms [Koltchinskii et al., 2011, Negahban and Wainwright, 2012, Klopp, 2014].
 Our contribution
-Recently, Bandeira et al.
-[2023] and Brailovskaya and van Handel [2024] introduce new matrix concentration inequalities, which establish sharper bounds for certain important random matrices.
-We showcase how the new bounds can be leveraged in the matrix completion problem to drop the logarithmic dimension factor.
-Specifically, We consider five estimators across three different settings as follows:
-  • First setting: the unknown matrix is of low rank and the noise is sub-Gaussian.
-This is the most classical setting.
-Two estimators are studied in Klopp [2014], one for the case of known variance and another for the case of unknown variance.
-  • Second setting: the unknown matrix is of low rank and the noise is heavy tailed.
-The estimator in Yu et al.
-[2024] is designed for the case that the noise is only assumed to have finite second moment.
-  • Third setting: the unknown matrix is corrupted and the noise is sub-Gaussian.
-Two estimators are studied under different corruption models in Klopp et al.
-[2017]: entrywise corruption and columnwise corruption.
-All the previous results on the convergence rates of the above five estimators involve the logarithmic dimension factor.
-We remove the dimensional factor for all the five estimators, thereby proving their minimax rate optimality.
+In this paper, we revisit three estimators that have been studied in other research works:
+  • The unknown matrix is of low rank and the noise is heavy tailed. The estimator in Yu et al. [2024] is designed for the case that the noise is only assumed to have finite second moment.
+  • The unknown matrix is of low rank and the noise is sub-Gaussian.
+Nuclear norm penalized least squares is a common approach under this setting [Koltchinskii et al., 2011, Negahban and Wainwright, 2012, Klopp, 2014].
+We consider the estimator proposed in Klopp [2014] to illustrate our new results.
+  • The unknown matrix is of low rank and the noise is sub-Gaussian with unknown variance.
+The parameter tuning in matrix completion often requires prior knowledge of the noise variance.
+Klopp [2014] proposes an estimator to solve such a problem.
+All the previous results on the convergence rates of the above three estimators involve the logarithmic dimension factor.
+We remove the dimensional factor for all the three estimators, thereby proving their minimax rate optimality.
 The remainder of the paper is organized as follows.
@@
 In Section 2, we present our main results.
-We begin by discussing a classical matrix completion model and demonstrating how to eliminate the dimension factor.
-We then introduce additional examples and results.
 We give some concluding remarks in Section 3.
@@
 Notation
-To present our results, we summarize the notations used throughout this paper.
+To present our results, we collect the notations used throughout this paper for reader’s convenience.
   • $M=m_{1}\vee m_{2}$, $m=m_{1}\wedge m_{2}$, $d=m_{1}+m_{2}$.
@@
 Let $\{\sigma_{l}\}_{l=1}^{m}$ denote ordered singular values of $A$: $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{m}$.
-$\|A\|_{F}=\sqrt{\sum_{i=1}^{m_{1}}\sum_{j=1}^{m_{2}}a_{ij}^{2}}$, $\|A\|_{1}=\sum_{l=1}^{m}\sigma_{l}$, $\|A\|=\sigma_{1}$, $\|A\|_{\infty}=\max_{i\leq m_{1},j\leq m_{2}}|a_{ij}|$, $\|A\|_{1,*}=\sum_{i,j}|a_{ij}|$.
-With $A^{k}$ denoting the $k^{th}$ column of A, define $\|A\|_{2,1}=\sum_{k=1}^{m_{2}}\|A^{k}\|_{2}$, where $\|\cdot\|_{2}$ is the usual $l_{2}$-norm.
+$\|A\|_{F}=\sqrt{\sum_{i=1}^{m_{1}}\sum_{j=1}^{m_{2}}a_{ij}^{2}}$, $\|A\|_{*}=\sum_{l=1}^{m}\sigma_{l}$, $\|A\|_{\infty}=\max_{i\leq m_{1},j\leq m_{2}}|a_{ij}|$, $\|A\|=\sigma_{1}$.
 $\|A\|_{w(F)}=\sqrt{\sum_{i=1}^{m_{1}}\sum_{i=1}^{m_{2}}a^{2}_{ij}P_{ij}}$ is a weighted Frobenius norm associated with the distribution of the sampling matrices $X_{i}$.
@@
   • $\{\epsilon_{i}\}_{i=1}^{n}$ are i.i.d. symmetric Bernoulli random variables, namely
-    $P(\epsilon_{i}=1)=P(\epsilon_{i}=-1)=\frac{1}{2}.$
+    $\mathbb{P}(\epsilon_{i}=1)=\mathbb{P}(\epsilon_{i}=-1)=\frac{1}{2}.$
+    They are independent of the observations $\{(X_{i},Y_{i})\}_{i=1}^{n}$.
+  • We write $(\log d)^{\alpha}$ as $\log^{\alpha}d$ and $(\mathbb{E}[X])^{\alpha}$ as $\mathbb{E}^{\alpha}[X]$ for aesthetic reasons.
+  • A random variable $X$ is called $\psi_{\alpha}$ random variable if it has finite $\psi_{\alpha}$ norm,
+    $\|X\|_{\psi_{\alpha}}=\inf\left\{C>0;\mathbb{E}\exp\left(\left(\frac{|X|}{C}\right)^{\alpha}\right)\leq 2\right\}.$
+    In particular, when $\alpha=2$, $\|\cdot\|_{\psi_{2}}$ is called sub-gaussian norm and X is called sub-gaussian random variable.
 2 Main Results
-In this section, we present new upper bounds for the five estimators mentioned in Section 1.
-In particular, compared to the existing bounds, the new bounds do not involve the dimension factor $\log d$ and hence match with the minimax lower bounds (under mild conditions).
-2.1 A warming up example
-We start with the most classical setting, where the noise in model (1) has exponentially decaying tails [Koltchinskii et al., 2011, Negahban and Wainwright, 2012, Klopp, 2014].
+In this section, we present new upper bounds for the three estimators mentioned in Section 1.
+In particular, compared to the existing bounds, the new bounds do not involve the dimension factor $\log d$ and hence match with the minimax lower bounds.
+At first, we introduce assumptions used in our paper.
+Regarding the noise $\xi_{i}$, they can be heavy tailed or sub-Gaussian in different contexts and their distributions can depend on the sampled entry.
+We introduce the following assumptions.
 Assumption 1.
-The noise variables $\xi_{i}$ satisfy $E[\xi_{i}]=0$, $E[\xi^{2}_{i}]=\sigma^{2}$ and
-$E\exp(t\xi_{i}/\sigma)\leq\exp(t^{2}/L_{1}),\forall t\in\mathbb{R},$
-where $L_{1}\geq 0$ is a constant.
+Given $X_{i}$, the noise variables satisfy
+$E[\xi_{i}|X_{i}]=0$
+and their second moments are uniformly bounded
+$E[\xi_{i}^{2}|X_{i}]\leq\sigma^{2}.$
+Assumption 2.
+The noise variables $\xi_{i}$ satisfy $E[\xi_{i}|X_{i}]=0$, $E[\xi^{2}_{i}]=\sigma^{2}$ and given $X_{i}$, it holds uniformly
+$E\exp(t\xi_{i}/\sigma)\leq\exp(L_{1}^{2}t^{2}),\forall t\in\mathbb{R},$
+where $L_{1}>0$ is a constant.
 Regarding the distribution of $X_{i}$, we consider a non-uniform sampling distribution.
@@
 Hence, (4) is a very mild condition.
-Under the assumption of sub-Gaussian tailed noise, a classical estimator for $A_{0}$ is:
-$\hat{A}=\operatorname*{arg\,min}_{\|A\|_{\infty}\leq a}\left\{\dfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}+\lambda\|A\|_{1}\right\}.$ (2)
-We now present our new result:
-Theorem 1.
-Let Assumptions 1 and 2 hold. Assume that $n\geq C_{1}m\log^{6}d\cdot\log^{2}n$. There exists a constant $C_{2}$ such that if we take
-$\lambda=C_{2}\sigma\sqrt{\frac{1}{nm}},$
-with probability at least $1-\frac{2}{d}$, the following bound holds
-$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq\dfrac{C_{3}\mu^{2}\max(\sigma^{2},a^{2})rM}{n}.$ (3)
-Here, $C_{1}$ is a universal constant, and $C_{2},C_{3}$ depend on the constants $L_{1},L_{2},L_{3}$ from Assumptions 1-2.
-To compare with existing bounds, we mention one representative result from Klopp [2014]:
-Theorem 2 (Theorem 7 in Klopp [2014]).
-Let Assumption 1 and Conditions (1)-(3) in Assumption 2 hold. Assume that $n\geq C_{1}m\log^{2}d$. If we take
-$\lambda=C_{2}\sigma\sqrt{\dfrac{\log d}{nm}},$
-for a suitable constant $C_{2}$, the following bound holds with probability $1-\frac{3}{d}$:
-$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{3}\max\left\{\dfrac{\mu^{2}\max(a^{2},\sigma^{2})rM\log d}{n},\mu a^{2}\sqrt{\frac{\log d}{n}}\right\}.$
-Here, $C_{1}$ is a universal constant, and $C_{2},C_{3}$ depend on the constants $L_{1},L_{2}$ from Assumptions 1-2.
-Comparing Theorems 1 and 2, we make several remarks:
-  1. 1.
-    Our new result eliminates the $\log d$ factor in the convergence rate.
-Koltchinskii et al.
-[2011] provides the following minimax lower bound in the case of gaussian noise and uniform sampling distribution:
-    $\inf_{\hat{A}}\sup_{A_{0}}P\left(\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\geq C_{1}\dfrac{\min(\sigma^{2},a^{2})rM}{n}\right)\geq C_{2},$ (4)
-    where $C_{1},C_{2}$ are universal constants, and the infimum and supremum are taken over all possible estimators and rank-$r$ matrices respectively.
-Combining the new bound (3) with the minimax lower bound (4) shows that the removal of the dimension factor $\log d$ establishes the minimax rate optimality of the estimator (2).
-  2. 2.
-    Our analysis reveals a more precise choice for the tuning parameter $\lambda$: the optimal choice for $\lambda$ should be of the order $O(\sqrt{1/(nm)})$ instead of $O(\sqrt{(\log d)/(nm)})$.
-  3. 3.
-    The new bound requires a slightly stronger condition on $n$, as $n\geq Cm\log^{6}d\log^{2}n$ is needed rather than $n\geq Cm\log^{2}d$.
-The $\log n$ term arises from the truncation of the sub-Gaussian noise, in order to apply the matrix concentration inequalities from Brailovskaya and van Handel [2024].
-  4. 4.
-    Compared to the sampling distribution assumption in Klopp [2014], Condition (4) of Assumption 2 is an additional condition we impose.
-But as we explained after Assumption 2, this is a very mild condition.
-  5. 5.
-    The convergence rate in Theorem 2 takes a form of the maximum between two terms.
-In the high-dimensional regime where $n\leq m_{1}m_{2}$, it is straightforward to verify that the first term–the desirable term–is the larger one.
-However, in less challenging scenarios in which $n$ is sufficiently large, the second term dominates.
-This limitation arises from a Frobenius norm based peeling argument that Klopp [2014] used to prove restricted strong convexity.
-In contrast, we have adopted a different peeling scheme introduced in Wainwright [2019], which peels certain matrix space by the infinity norm and the nuclear norm.
-As a result, our new bound does not include the nuisance term $O(\sqrt{(\log d)/n})$.
-2.1.1 How to remove the $\log d$ factor
-In this section, we discuss where the dimensional factor $\log d$ arises and how to eliminate it.
-Denote $L(A)=\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}$.
-The high-dimensional analysis of (2) includes two basic steps [Negahban et al., 2012].
-The first step involves selecting $\lambda$ such that
-$\lambda\geq C\|\nabla L(A_{0})\|=2C\Big{\|}\dfrac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\Big{\|},$ (5)
-to ensure that $\hat{A}-A_{0}$ lies in a nearly low rank space $\Omega$. In the second step, we establish strong convexity of $L(A)$ over $\{A:A-A_{0}\in\Omega\}$, namely,
-$L(\hat{A})-L(A_{0})-\langle\nabla L(A_{0}),\hat{A}-A_{0}\rangle=\frac{1}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}-A_{o}\rangle^{2}\geq C\|\hat{A}-A_{0}\|^{2}_{w(F)}-\mathcal{E},$
-where $\mathcal{E}$ represents a tolerance term and $\|A\|^{2}_{w(F)}=E\langle X_{i},A\rangle^{2}$. To achieve this, we shall obtain bounds on the following empirical processes
-$Z=\sup_{A-A_{0}\in\Omega}\left|\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},A-A_{0}\rangle^{2}-\|A-A_{0}\|^{2}_{w(F)}\right|.$
-Bounding this type of empirical processes, after applying Talagrand concentration and symmetrization & contraction, invokes a bound on
-$E\left\|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|.$ (6)
-With the two steps completed, we can use some standard non-asymptotic arguments to obtain upper bounds for the convergence rate in averaged Frobenius norm.
-Specifically, Lemma 14 reveals how (5) and (6) influence the error rate: with probability at least $1-\frac{1}{d}$, it holds
-$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\apprle\mu^{2}\left(\lambda^{2}+a^{2}E^{2}\left\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|\right)rm_{1}m_{2}.$ (7)
-Traditional matrix concentration inequalities (e.g. Theorem 6.17 and Example 6.18 in Wainwright [2019]; Lemmas 5 and 6 in Klopp [2014]), yield
-$\left\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\leq C_{1}\sigma\sqrt{\frac{\log d}{nm}}$
-holds with high probability and
-$E\left\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|\leq C_{2}\sqrt{\frac{\log d}{nm}}.$
-Plugging $\lambda=C_{1}\sigma\sqrt{\frac{\log d}{nm}}$ and $E\left\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|\leq C_{2}\sqrt{\frac{\log d}{nm}}$ into (7) leads to the dimension factor dependent result in Theorem 2.
-To remove the $\log d$ factor, we refine the evaluation of the two key quantities $\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\|$ and $E\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\|$.
-Recently, Bandeira et al.
-[2023] introduce new matrix concentration inequalities for random matrices with Gaussian entries using novel mathematical tools.
-Later, Brailovskaya and van Handel [2024] extend the results for more general distributions.
-This type of concentration inequalities fully take the non-commutativity of random matrices into account and provide sharper bounds.
-For instance, their results on the expectation of the spectral norm of a random matrix $X\in\mathbb{R}^{m_{1}\times m_{2}}$ typically take the form:
-$E\|X\|\apprle\max(\|EX^{T}X\|,\|EXX^{T}\|)^{1/2}+v(X)\log^{\beta}d,$
-where $\max(\|EX^{T}X\|,\|EXX^{T}\|)^{1/2}$ is the desirable part and $\beta$ is a positive constant.
-The term $v(X)$ is a statistic that captures the non-commutativity of the random matrix $X$ and is often smaller than $\max(\|EX^{T}X\|,\|EXX^{T}\|)^{1/2}$ in many situations.
-If we are given that
-$v(X)\log^{\beta}d\leq\max(\|EX^{T}X\|,\|EXX^{T}\|)^{1/2},$
-the bound is free of the dimensional factor $\log^{\beta}d$.
-Our computations confirm that this holds for $\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\|$ and $E\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\|$.
-Assuming some mild conditions, we can obtain
-$P\left(\left\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\geq C_{1}\sigma\sqrt{\dfrac{1}{nm}}\right)\leq\dfrac{C_{2}}{d},$
-and
-$E\left\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|\leq C\sqrt{\dfrac{1}{nm}}.$
-As a result, the $\log d$ factor is successfully removed. See Lemmas 12 and 13 for more details.
-A similar issue appears in the existing results for the other four estimators.
-The advanced inequalities in Brailovskaya and van Handel [2024] can also be adapted to resolve the issue.
-We present sharper bounds for the remaining estimators in Sections 2.2-2.4.
-2.2 Matrix completion with unknown variance of the noise
-The choice of $\lambda$ in Theorem 2 is dependent on the variance of the noise.
-In practice, $\sigma$ is often unknown.
-To deal with this issue, Klopp [2014] proposes the following square-root lasso type estimator and studies its convergence rate.
-$\hat{A}_{S}=\operatorname*{arg\,min}_{\|A\|_{\infty}\leq a}\left\{\sqrt{\dfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}}+\lambda\|A\|_{1}\right\}.$ (8)
-We obtain a new result for (8).
-Theorem 3.
-Let Assumptions 1 and 2 hold. Let $n\geq C_{1}m\log^{6}d\log^{2}n$. If we take
-$\lambda=C_{2}\sqrt{\frac{1}{nm}},$
-with probability at least $1-\frac{2}{d}-2\exp(-C_{3}n)$, the following bound holds
-$\dfrac{\|\hat{A}_{S}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq\dfrac{C_{4}\mu^{2}\max(\sigma^{2},a^{2})rM}{n},$ (9)
-provided that $n\geq\frac{C_{2}}{10}\mu rM$.
-Here, $C_{1}$ is a universal constant, $C_{3}$ depends on $L_{1}$, and $C_{2},C_{4}$ depend on $L_{1},L_{2},L_{3}$, where $L_{i}$’s are constants from Assumptions 1-2.
-We can compare the above bound with the one obtained in Klopp [2014].
-Theorem 4 (Theorem 10 in Klopp [2014]).
-Let Assumption 1 and Conditions (1)-(3) in Assumption 2 hold. If we take
-$\lambda=C_{1}\sqrt{\frac{\log d}{nm}},$
-the following upper bound holds with probability $1-\frac{3}{d}-2\exp(-C_{2}n)$,
-$\dfrac{\|\hat{A}_{S}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{3}\left\{\dfrac{\mu^{2}\max(a^{2},\sigma^{2})rM\log d}{n},a^{2}\mu\sqrt{\frac{\log d}{n}}\right\},$
-provided that $n\geq 8C_{2}\mu rM\log d$. Here $C_{1},C_{3}$ depend on $L_{1},L_{2}$ and $C_{2}$ depends on $L_{1}$, where $L_{1},L_{2}$ are constants from Assumptions 1-2.
-Comparing the convergence rate in Theorem 4 and the lower bound in (4), they are not matched due to the existence of the $\log d$ factor.
-Our new bound in (9) resolves the issue and shows that the estimator (8) is indeed minimax rate optimal.
-Regarding more comparisons between Theorems 3 and 4, we refer to remarks after Theorem 2 for similar comments.
+In following subsections, we introduce our $\log d$-free error bounds.
 2.1 Matrix completion with heavy tailed noise
+We start with our improvement on a recent advance in matrix completion [Yu et al., 2024].
 Heavy-tailed data is ubiquitous in many fields such as finance, microeconomics and biology.
@@
 $l_{\tau}(x)=\begin{cases}\frac{x^{2}}{2},\leavevmode\nobreak\ \leavevmode\nobreak\ |x|\leq\tau,\\ \tau|x|-\frac{\tau^{2}}{2},\leavevmode\nobreak\ \leavevmode\nobreak\ |x|>\tau,\\ \end{cases}$
-and Huber loss is defined as $\frac{1}{n}\sum_{i=1}^{n}l_{\tau}(Y_{i}-\langle X_{i},A\rangle)$. The derivative of $l_{\tau}(\cdot)$ is denoted by
-$\psi_{\tau}(x)=\begin{cases}x,|x|\leq\tau,\\ \operatorname{sign}(x)\tau,|x|>\tau.\end{cases}$
-The estimator to be studied is
-$\hat{A}_{H}=\operatorname*{arg\,min}_{\|A\|_{\infty}\leq a}\left\{\dfrac{1}{n}\sum_{i=1}^{n}l_{\tau}(Y_{i}-\langle X_{i},A\rangle)+\lambda\|A\|_{1}\right\}.$ (10)
-We formally propose the following assumption.
-Assumption 3.
-$\xi_{i}$ have finite second moments, namely
-$E[\xi_{i}^{2}]=\sigma^{2}.$
+and Huber loss is defined as $\frac{1}{n}\sum_{i=1}^{n}l_{\tau}(Y_{i}-\langle X_{i},A\rangle)$. The estimator to be studied is
+$\hat{A}_{H}=\operatorname*{argmin}_{\|A\|_{\infty}\leq a}\left\{\dfrac{1}{n}\sum_{i=1}^{n}l_{\tau}(Y_{i}-\langle X_{i},A\rangle)+\lambda\|A\|_{*}\right\}.$ (4)
 We obtain the following new convergence rate for $\hat{A}_{H}$.
 Theorem 1.
-Let Assumptions 2 and 3 hold. We additionally assume $n\geq 160m\log^{4}d$ and $\xi_{i}$ are symmetric. If we take
+Let Assumptions 1 and 3 hold. We additionally assume $n\geq 160m\log^{4}d$ and $\xi_{i}|X_{i}$ are symmetric. If we take
 $\lambda=C_{1}\max(\sigma,a)\sqrt{\frac{1}{nm}},$
@@
 with probability at least $1-\frac{2}{d}$, the following bound holds
-$\dfrac{\|\hat{A}_{H}-A_{0}\|^{2}_{F}}{m_{1}m_{2}}\leq C_{2}\frac{\max(\sigma^{2},a^{2})rM}{n}.$ (11)
+$\dfrac{\|\hat{A}_{H}-A_{0}\|^{2}_{F}}{m_{1}m_{2}}\leq C_{2}\frac{\max(a^{2},\sigma^{2})rM}{n}.$ (5)
 Here, $C_{1}$ and $C_{2}$ depend on the constants $L_{2},L_{3}$ from Assumption 3.
@@
 then with probability at least $1-\frac{3}{d}$, we have
-$\dfrac{\|\hat{A}_{H}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{3}\left\{\max(a^{2},\sigma^{2})\dfrac{rM\log d}{n},a^{2}\sqrt{\frac{\log d}{n}}\right\}$
-provided that $n\geq C_{4}m\log d$. Here, $C_{1},\ldots,C_{3}$ are universal constants.
-We have the following remarks for the comparison of Theorems 5 and 6.
+$\dfrac{\|\hat{A}_{H}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{3}\max\left\{\max(a^{2},\sigma^{2})\dfrac{rM\log d}{n},a^{2}\sqrt{\frac{\log d}{n}}\right\}$
+provided that $n\geq C_{4}m\log d$. Here, $C_{1}\sim C_{3}$ are universal constants.
+Theorem 1 improves over Theorem 2 in several aspects.
   1. 1.
-    Theorem 5 improves over Theorem 6 in several aspects: (a) Theorem 5 allows for a more general sampling distribution; (b) It drops the nuisance term $O(\sqrt{(\log d)/n})$ from the convergence rate, via the peeling by infinity and nuclear norm to the Huber loss function; (c) Like in Theorems 1 and 3, the $\log d$ factor is removed by utilizing the advanced matrix concentration inequalities in Brailovskaya and van Handel [2024].
+    Our new result eliminates the $\log d$ factor in the convergence rate.
+It is achieved by utilizing new advanced concentration inequalities in Brailovskaya and van Handel [2024].
+See our Lemma 14 for a detailed analysis.
+Koltchinskii et al.
+[2011] provides the following minimax lower bound in the case of gaussian noise and uniform sampling distribution:
+    $\inf_{\hat{A}}\sup_{A_{0}}\mathbb{P}\left(\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\geq C_{1}\dfrac{\min(\sigma^{2},a^{2})rM}{n}\right)\geq C_{2},$ (6)
+    where $C_{1},C_{2}$ are universal constants, and the infimum and supremum are taken over all possible estimators and rank-$r$ matrices respectively.
+Combining the new bound (5) with the minimax lower bound (6) shows that the removal of the dimension factor $\log d$ establishes the minimax rate optimality of the estimator (4).
   2. 2.
-    The lower bound in (4) is derived assuming that the noises are Gaussian.
-Since the Gaussian distribution is of finite second moments, the lower bound also applies to the heavy tailed case.
-Thus, our new bound (11) together with (4) proves the minimax optimality for the estimator (10).
+    Theorem 1 allows for a more general sampling distribution while Theorem 2 requires uniform distribution.
+We also eliminate the nuisance term $O(\sqrt{(\log d)/n})$ from the convergence rate, via a different peeling argument [Wainwright, 2019] in the analysis of certain empirical processes.
   3. 3.
+    Our analysis reveals a more precise choice for the tuning parameter $\lambda$: the optimal choice for $\lambda$ should be of the order $O(\sqrt{1/(nm)})$ instead of $O(\sqrt{(\log d)/(nm)})$, which sheds light on the practice of parameter tuning.
+  4. 4.
     Theorem 1 requires symmetry of the noise, while Theorem 2 does not.
@@
 While this condition is very mild, this seems to be a pure technical condition and is expected to be unnecessary under future investigation.
-See our analysis in the proof of (11).
-2.4 Matrix completion under corruption
-In real applications, the true matrix $A_{0}$ may be corrupted.
-We consider the situation where the observed noisy entry is from a corrupted matrix $A_{0}=L_{0}+S_{0}$ where $L_{0}$ is of rank $r$ and $S_{0}$ is a sparse corruption matrix.
-Klopp et al.
-[2017] considers two corruption models: entrywise corruption and columnwise corruption.
-We adopt the notation of Klopp et al.
+Moreover, assuming that the $\xi_{i}$ have a finite moment of order $2+\kappa$ for some $\kappa>0$, a $\log d$-free bound can be established under a slightly weaker condition on sample size:
+    $n\geq\max\left(m\log^{4+\frac{1}{\kappa+1}}d,\log^{\frac{4}{\kappa}}d\right).$
+    See our analysis in the proof of inequality (5).
+2.2 Matrix completion with known variance
+The setting that the noise have distributions with exponential decay is most studied, see, e.g., Koltchinskii et al.
+[2011], Negahban and Wainwright [2012], Klopp [2014].
+Under the assumption of light tailed noise, an estimator proposed in Klopp [2014] is:
+$\hat{A}=\operatorname*{argmin}_{\|A\|_{\infty}\leq a}\left\{\dfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}+\lambda\|A\|_{*}\right\}.$ (7)
+We now present our new result:
+Theorem 3.
+Let Assumptions 2 and 3 hold. Assume that $n\geq C_{1}m\log^{4}d\leavevmode\nobreak\ (\log d+\log n)$. If we take
+$\lambda=C_{2}\sigma\sqrt{\frac{1}{nm}},$
+the following bound holds with probability at least $1-\frac{2}{d}$
+$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq\dfrac{C_{3}\mu^{2}\max(a^{2},\sigma^{2})rM}{n}.$ (8)
+Here, $C_{1}$ is a universal constant, and $C_{2},C_{3}$ depend on the constants $L_{1},L_{2},L_{3}$ from Assumptions 2-3.
+To compare with existing bounds, we mention the following theorem:
+Theorem 4 (Theorem 7 in Klopp [2014]).
+Let Assumption 2, Conditions (1)-(3) in Assumption 3 hold. Assume that $n\geq C_{1}m\log^{2}d$. If we take
+$\lambda=C_{2}\sigma\sqrt{\dfrac{\log d}{nm}},$
+the following bound holds with probability $1-\frac{3}{d}$:
+$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{3}\max\left\{\dfrac{\mu^{2}\max(a^{2},\sigma^{2})rM\log d}{n},a^{2}\mu\sqrt{\frac{\log d}{n}}\right\}.$
+Here, $C_{1}$ is a universal constant, and $C_{2},C_{3}$ depend on the constants $L_{1},L_{2}$ from Assumptions 2-3.
+Comparing Theorems 3 and 4, we make several remarks:
+  1. 1.
+    Theorem 3 eliminates a $\log d$ factor in the convergence rate. Comparing (8) and (6), the removal of the $\log d$ factor establishes the minimax optimality of estimator (7).
+  2. 2.
+    Theorem 3 also provides the correct order of tuning parameter $\lambda$ for (7).
+The old one is $O(\sqrt{\log d/(nm)})$ while the correct one in Theorem 3 is $O(\sqrt{1/(nm)})$.
+The $\log d$ difference can be significant in high dimensional settings.
+  3. 3.
+    Theorem 4 involves an additional nuisance term $O(\sqrt{\log d/n})$, similarly to Theorem 2.
+In the high- dimensional regime where $n\leq m_{1}m_{2}$, it is straightforward to verify that the first term–the desirable term–is the larger one.
+However, in less challenging scenarios in which $n$ is sufficiently large, the nuisance term dominates.
+This limitation arises from a Frobenius norm based peeling argument that Klopp [2014] used to prove restricted strong convexity.
+In contrast, we have adopted a different peeling scheme introduced in Wainwright [2019], which peels certain matrix space by the infinity norm and the nuclear norm.
+As a result, our new bound does not include the nuisance term $O(\sqrt{\log d/n})$.
+  4. 4.
+    Our new bound requires a slightly stronger condition on $n$.
+The $\log n$ term arises from the truncation of the sub-Gaussian noise, in order to apply the matrix concentration inequalities from Brailovskaya and van Handel [2024].
+In the most interesting regime where $n\leq m_{1}m_{2}$, the sample size requirement in Theorem 3 becomes $n\geq Cm\log^{5}d$ which is in the typical form of $n\geq m\text{Poly}(\log d)$ in the matrix completion literature.
+Under both with-replacement and without-replacement settings, the order of $\text{Poly}(\log d)$ often vary with problem assumptions and methods used; see, for example, Koltchinskii et al.
+[2011], Klopp [2014], Chen et al.
 [2020].
-The total number of observations is denoted by $N$.
-The set of uncorrupted observations is denoted by $\Omega$, namely $Y_{i}=\langle X_{i},L_{0}\rangle+\xi_{i}$ for $i\in\Omega$.
-The set of corrupted observations is denoted by $\tilde{\Omega}$, namely $Y_{i}=\langle X_{i},L_{0}+S_{0}\rangle+\xi_{i}$ for $i\in\tilde{\Omega}$.
-We denote $n=|\Omega|$, $N-n=|\tilde{\Omega}|$, and $\alpha=\frac{N}{n}$.
-2.4.1 Entrywise corruption
-For entrywise corruption, the set of corrupted entries is denoted by
-$\mathcal{I}_{1}=\{(j,k):S_{0}(j,k)\neq 0\},$
-where $S_{0}(j,k)$’s are entries of the corruption matrix $S_{0}$. We assume that $S_{0}$ has at most $s$ non-zero entries, i.e. $|\mathcal{I}_{1}|\leq s$. We also need the following assumption.
-Assumption 4.
-For $i\in\Omega$ and $i\in\tilde{\Omega}$, the sampling is uniform:
-$P(X_{i}\in e_{j}(m_{1})e_{k}^{T}(m_{2}))=\dfrac{1}{|\mathcal{I}_{1}|},\quad{\rm for\leavevmode\nobreak\ }(j,k)\in\mathcal{I}_{1},i\in\tilde{\Omega}$
-and
-$P(X_{i}\in e_{j}(m_{1})e_{k}^{T}(m_{2}))=\dfrac{1}{m_{1}m_{2}-|\mathcal{I}_{1}|},\quad{\rm for\leavevmode\nobreak\ }(j,k)\in\mathcal{I}_{1}^{c},i\in\Omega.$
-Klopp et al.
-[2017] considers general sampling distributions.
-To ease the notation and demonstrate more clearly the ideas of our method, we focus on the uniform sampling distribution in Assumption 4.
-Our results can be easily extended to the general distribution case.
-For this corruption model, Klopp et al.
-[2017] considers the following estimator:
-$(\hat{L}_{E},\hat{S}_{E})=\operatorname*{arg\,min}_{\begin{subarray}{c}\|L\|_{\infty}\leq a\\ \|S\|_{\infty}\leq a\end{subarray}}\left\{\dfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},L+S\rangle)^{2}+\lambda_{1}\|L\|_{1}+\lambda_{2}\|S\|_{1,*}\right\}.$ (12)
-We obtain the following theorem.
+In our results, the leverage of the concentration inequalities [Brailovskaya and van Handel, 2024] causes a small change of the order of $\text{Poly}(\log d)$.
+This fluctuation is acceptable in light of the substantial gain in sharpness of the resulting bounds.
+We emphasize that previous concentration inequalities cannot yield bounds as sharp as ours, even when the sample size is fixed at the same level as in our results.
+2.3 Matrix completion with unknown variance of the noise
+The choice of $\lambda$ in Theorem 4 is dependent on the variance of the noise.
+In practice, $\sigma$ is often unknown.
+To deal with this issue, Klopp [2014] proposes the following square-root lasso type estimator and studies its convergence rate.
+$\hat{A}_{S}=\operatorname*{argmin}_{\|A\|_{\infty}\leq a}\left\{\sqrt{\dfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}}+\lambda\|A\|_{*}\right\}.$ (9)
+We obtain a new result for (9).
 Theorem 5.
-Let Assumptions 1 and 4 hold. Let $n\geq C_{1}m\log^{6}d\log^{2}n$ and $N\leq m_{1}m_{2}$. Taking
-$\lambda_{1}=C_{2}\max(a,\sigma)\sqrt{\frac{1}{nm}},$
-$\lambda_{2}=C_{3}\max(a,\sigma)\frac{\log d}{N},$
-the following bound holds with probability at least $1-\frac{6}{d}$,
-$\dfrac{\|\hat{L}_{E}-L_{0}\|_{F}^{2}}{m_{1}m_{2}}+\dfrac{\|\hat{S}_{E}-S_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{4}\Bigg{(}\alpha\max(\sigma^{2},a^{2})\left(\dfrac{rM}{n}+\dfrac{|\tilde{\Omega}|\log d}{n}\right)+\dfrac{a^{2}s}{m_{1}m_{2}}\Bigg{)}.$ (13)
-Here, $C_{1}$ is a universal constant, and $C_{2},C_{3},C_{4}$ depend on the constant $L_{1}$ from Assumption 1.
-Klopp et al. [2017] gives the following result.
-Theorem 8 (Corollary 11 of Klopp et al. [2017]).
-Let Assumptions 1 and 4 hold. Assume that $n\geq C_{1}m\log d\log^{2}m$ and $N\leq m_{1}m_{2}$. Taking
-$\lambda_{1}=C_{2}\max(a,\sigma)\sqrt{\frac{\log d}{nm}},$
-$\lambda_{2}=C_{3}\max(a,\sigma)\frac{\log d}{N},$
-the following bound holds with probability at least $1-\frac{6}{d}$,
-$\dfrac{\|\hat{L}_{E}-L_{0}\|_{F}^{2}}{m_{1}m_{2}}+\dfrac{\|\hat{S}_{E}-S_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{4}\left(\alpha\max(\sigma^{2},a^{2})\left(\dfrac{rM\log d}{n}+\dfrac{|\tilde{\Omega}|\log d}{n}\right)+\dfrac{a^{2}s}{m_{1}m_{2}}\right).$ (14)
-Here, $C_{1}$ is a universal constant, and $C_{2},C_{3},C_{4}$ depend on the constant $L_{1}$ from Assumption 1.
-Klopp et al. [2017] also obtains the minimax lower bound under this setting assuming $\xi_{i}\sim\mathcal{N}(0,\sigma^{2})$:
-$\inf_{(\hat{L},\hat{S})}\sup_{\mathcal{A}}P\left(\dfrac{\|\hat{L}-L_{0}\|_{F}^{2}}{m_{1}m_{2}}+\dfrac{\|\hat{S}-S_{0}\|_{F}^{2}}{m_{1}m_{2}}\apprge\min(\sigma^{2},a^{2})\left(\dfrac{rM+|\tilde{\Omega}|}{n}+\dfrac{s}{m_{1}m_{2}}\right)\right)\geq C.$ (15)
-Here, $C\in(0,1)$, the infimum is taken over all possible estimators for $(L_{0},S_{0})$, and $\mathcal{A}$ is the parameter space such that $\text{rank}(L_{0})\leq r$ and $|\mathcal{I}_{1}|\leq s$.
-Our new result removes the $\log d$ factor in the term $\frac{rM\log d}{n}$ in (14).
-When $rM\geq|\tilde{\Omega}|\log d$, namely the number of corrupted observations is not too large, (13) and (15) together show that the estimator (12) is minimax rate optimal.
-2.4.2 Columnwise corruption
-For columnwise corruption model, we assume there are at most $s$ corrupted columns, namely $|\mathcal{I}_{2}|\leq s$, where $\mathcal{I}_{2}=\{1\leq i\leq m_{2};S_{0}^{i}\neq\textbf{0}\in\mathbb{R}^{m_{1}}\}$ includes the indices of non-zero columns of $S_{0}$.
-And we make the following assumption.
-Assumption 5.
-For $i\in\Omega$, the sampling distribution is uniform:
-$P(X_{i}=e_{j}(m_{1})e_{k}^{T}(m_{2}))=\dfrac{1}{m_{1}(m_{2}-|\mathcal{I}_{2}|)}.$
-For $i\in\tilde{\Omega}$, the sampling is uniform:
-$P(X_{i}=e_{j}(m_{1})e_{k}^{T}(m_{2}))=\dfrac{1}{m_{1}|\mathcal{I}_{2}|}.$
-Klopp et al. [2017] proposes the following estimator:
-$(\hat{L}_{C},\hat{S}_{C})=\operatorname*{arg\,min}_{\begin{subarray}{c}\|L\|_{\infty}\leq a\\ \|S\|_{\infty}\leq a\end{subarray}}\left\{\dfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},L+S\rangle)^{2}+\lambda_{1}\|L\|_{1}+\lambda_{2}\|S\|_{2,1}\right\}.$ (16)
-We obtain the following theorem.
-Theorem 9.
-Let Assumptions 1 and 5 hold. Assume $n\geq 2\log d\max(m_{2},m\log^{2}m)$, $N\leq m_{1}m_{2}$, and $n\geq C_{1}m\log^{6}d\log^{2}n$. Taking
-$\lambda_{1}=C_{2}\max(a,\sigma)\sqrt{\frac{1}{Nm}},$
-$\lambda_{2}=C_{3}\max(a,\sigma)\sqrt{\frac{\log d}{Nm_{2}}},$
-with probability at least $1-\frac{6}{d}$, we have
-$\dfrac{\|\hat{L}_{C}-L_{0}\|_{F}^{2}}{m_{1}m_{2}}+\dfrac{\|\hat{S}_{C}-S_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{4}\left(\alpha\max(\sigma^{2},a^{2})\left(\dfrac{rM}{n}+\dfrac{|\tilde{\Omega}|\log d}{n}\right)+\dfrac{a^{2}s}{m_{2}}\right).$ (17)
-Here, $C_{1}$ is a universal constant, and $C_{2},C_{3},C_{4}$ depend on the constant $L_{1}$ from Assumption 1.
-Klopp et al. [2017] obtains the following result.
-Theorem 10 (Corollary 7 in Klopp et al. [2017]).
-Let Assumptions 1 and 5 hold. Assume $N\leq m_{1}m_{2}$, and $n\geq 2\log d\max(m_{2},m\log^{2}m)$. Taking
-$\lambda_{1}=C_{2}\max(a,\sigma)\sqrt{\frac{\log d}{Nm}},$
-$\lambda_{2}=C_{3}\max(a,\sigma)\sqrt{\frac{\log d}{Nm_{2}}},$
-the following bound holds with probability at least $1-\frac{6}{d}$,
-$\dfrac{\|\hat{L}_{C}-L_{0}\|_{F}^{2}}{m_{1}m_{2}}+\dfrac{\|\hat{S}_{C}-S_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{4}\left(\alpha\max(\sigma^{2},a^{2})\left(\dfrac{rM\log d}{n}+\dfrac{|\tilde{\Omega}|\log d}{n}\right)+\dfrac{a^{2}s}{m_{2}}\right),$ (18)
-where $C_{1}$ is a universal constant, and $C_{2},C_{3},C_{4}$ depend on the constant $L_{1}$ from Assumption 1.
-According to Theorem 3 in Klopp et al. [2017], we have the following minimax lower bound for columnwise corruption assuming $\xi_{i}\sim\mathcal{N}(0,\sigma^{2})$
-$\inf_{(\hat{L},\hat{S})}\sup_{\mathcal{A}}P_{A_{0}}\left(\dfrac{\|\hat{L}-L_{0}\|_{F}^{2}}{m_{1}m_{2}}+\dfrac{\|\hat{S}-S_{0}\|_{F}^{2}}{m_{1}m_{2}}\apprge\min(\sigma^{2},a^{2})\left(\dfrac{rM+|\tilde{\Omega}|}{n}+\dfrac{s}{m_{2}}\right)\right)\geq C,$ (19)
-where $C\in(0,1)$ is a universal constant, the infimum is taken over all the estimators, and $\mathcal{A}$ is the parameter space such that $\text{rank}(L_{0})\leq r$ and $|\mathcal{I}_{2}|\leq s$.
-We remove the $\log d$ factor attached to $\frac{rM}{n}$ in (18).
-Comparing (19) and (17), we can conclude that the estimator (16) is minimax rate optimal provided that $rM\geq|\tilde{\Omega}|\log d$, i.e.
-the set of corrupted observations is not too large.
+Let Assumptions 2, 3 hold. Let $n\geq C_{1}m\log^{4}d\leavevmode\nobreak\ (\log d+\log n)$. We can take
+$\lambda=C_{2}\sqrt{\frac{1}{nm}},$
+and with probability at least $1-\frac{2}{d}-2\exp(-C_{3}n)$, the following bound holds
+$\dfrac{\|\hat{A}_{S}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq\dfrac{C_{4}\mu^{2}\max(a^{2},\sigma^{2})rM}{n},$ (10)
+provided that $n\geq 16\mu Mr$. Here, $C_{1}$ is a universal constant, $C_{3}$ depends on $L_{1}$, and $C_{2},C_{4}$ depend on $L_{1},L_{2},L_{3}$, where $L_{i}$’s are constants from Assumptions 2-3.
+We can compare the above bound with the one obtained in Klopp [2014].
+Theorem 6 (Theorem 10 in Klopp [2014]).
+Let Assumption 2 and and Conditions (1)-(3) in Assumption 3 hold. We can take
+$\lambda=C_{1}\sqrt{\frac{\log d}{nm}},$
+and the following upper bound holds with probability $1-\frac{3}{d}-2\exp(-C_{2}n)$,
+$\dfrac{\|\hat{A}_{S}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{3}\max\left\{\dfrac{\mu^{2}\max(a^{2},\sigma^{2})rM\log d}{n},a^{2}\mu\sqrt{\frac{\log d}{n}}\right\},$
+provided that $n\geq 8C_{1}\mu rM\log d$. Here $C_{1},C_{3}$ depend on $L_{1},L_{2}$ and $C_{2}$ depends on $L_{1}$, where $L_{1},L_{2}$ are constants from Assumptions 2-3.
+Comparing the convergence rate in Theorem 6 and the lower bound in (6), they are not matched due to the existence of the $\log d$ factor.
+Our new bound in (10) resolves the issue and shows that the estimator (9) is indeed minimax rate optimal.
+Regarding more comparisons between Theorems 5 and 6, we refer to remarks after Theorem 4 for similar comments.
 3 Conclusion
-In this paper, we considered three popular matrix completion settings: (1) the unknown matrix is of low rank and the noise is sub-Gaussian; (2) the unknown matrix is of low rank and the noise is heavy tailed; (3) the unknown matrix is a low rank matrix corrupted by a sparse matrix in the entrywise or columnwise sense, and the noise is sub-Gaussian.
-We revisited five existing estimators in these settings, and developed sharper upper bounds to establish their minimax rate optimality.
-The key technical contribution lies in employing powerful matrix concentration inequalities introduced in Brailovskaya and van Handel [2024] to successfully remove the dimension factor $\log d$ that has remained in the upper bounds prior to our work.
-The $\log d$ factor problem not only exists for the fives estimators studied in this paper, but is ubiquitous throughout the literature.
-We believe that many other existing results on matrix completion can be possibly improved by applying these new advanced inequalities.
+In this paper, we considered three common matrix completion settings: (1) the unknown matrix is of low rank and the noise is heavy tailed; (2) the unknown matrix is of low rank and the noise is sub-Gaussian with known variance; (3) the unknown matrix is of low rank and the noise is sub-Gaussian with unknown variance.
+We revisited three popular estimators in these settings and developed sharper upper bounds to establish their minimax rate optimality.
+A key technical contribution lies in employing powerful matrix concentration inequalities introduced in Brailovskaya and van Handel [2024] to successfully remove the dimension factor $\log d$ that has remained in the upper bounds prior to our work.
+As previously noted, the $\log d$ factor is ubiquitous in matrix completion problems under the sampling with replacement setting.
+Our sharp analyses (e.g., Lemma 14 establishes dimension-free results for a variety of random matrices) can be applied or adapted to improve other existing results on matrix completion [Koltchinskii et al., 2011, Negahban and Wainwright, 2012, Klopp et al., 2017, Chen et al., 2024].
 4 Proof
-4.1 Proof of sharp bounds for spectral norms
-In this section, we present advanced inequalities introduced in [Brailovskaya and van Handel, 2024] and how to use them.
-Let $Q=\sum_{i=1}^{n}Q_{i}\in\mathbb{R}^{m_{1}\times m_{2}}$ be a random matrix where $Q_{i}$ are independent random matrices with mean zero.
-We define
-$\gamma(Q)$ $=\left(\max(\|EQQ^{T}\|,\|EQ^{T}Q\|)\right)^{1/2},$
-$\gamma_{*}(Q)$ $=\sup_{\|y\|_{2}=\|z\|_{2}=1}\left(E(y^{T}Qz)^{2}\right)^{1/2},$
-$g(Q)$ $=\|\operatorname{Cov}(Q)\|^{1/2},$
-$R(Q)$ $=\max_{1\leq i\leq n}\|Q_{i}\|,$
-where $\operatorname{Cov}(Q)$ is the covariance matrix of all the entries of $Q$ and its dimension is $m_{1}m_{2}\times m_{1}m_{2}$.
-Theorem 11 (Corrollary 2.17 in [Brailovskaya and van Handel, 2024]).
-With the above notations, there exist a universal constant $C$ such that
-$P(\|Q\|\geq 2\gamma(Q)+C(g(Q)^{1/2}\gamma(Q)^{1/2}\log^{3/4}d+\gamma_{*}(Q)t^{1/2}+R(Q)^{1/3}\gamma(Q)^{2/3}t^{2/3}+R(Q)t))\leq de^{-t}$ (20)
-and a universal constant $C$ such that
-$E\|Q\|\leq 2\gamma(Q)+C(g(Q)^{1/2}\gamma(Q)^{1/2}\log^{3/4}d+R(Q)^{1/3}\gamma(Q)^{2/3}\log^{2/3}d+R(Q)\log d).$ (21)
-The matrix inequalities in [Brailovskaya and van Handel, 2024] are formulated for $Q_{i}\in\mathbb{R}^{d\times d}$ (not necessarily self-adjoint), and can be extended to rectangular matrices in the above form.
-See Remark 2.1 in [Brailovskaya and van Handel, 2024].
-Applying the above inequalities, we have the following results.
+Throughout this section, we denote
+$L(A)=\frac{1}{n}\sum_{i=1}^{n}l_{\tau}(Y_{i}-\langle X_{i},A\rangle).$
+The derivative of $l_{\tau}(\cdot)$ is denoted by $\phi_{\tau}(\cdot)$, where
+$\phi_{\tau}(x)=\begin{cases}x,|x|\leq\tau,\\ \operatorname{sign}(x)\tau,|x|>\tau.\end{cases}$
+In this section, $c_{0}\sim c_{4}$ are specific constants whose values do not change at each occurrence.
+4.1 Proof of Theorem 1
+We begin with the following general result.
 Lemma 7.
-Assume the sampling matrices $X_{i}$ satisfy Assumption 2 and the noise random variables $\xi_{i}$ satisfy Assumption 1. The following inequality holds with probability at least $1-\frac{1}{d}$,
-$\left\|\dfrac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\leq C\sigma\sqrt{\dfrac{1}{nm}},$ (22)
-given $n\geq m\log^{6}d\cdot\log^{2}n$.
-Here $C$ is a constant dependent on $L_{1}\sim L_{3}$.
-Assume $X_{i}$ satisfy Assumption 2 , $\xi_{i}$ satisfy Assumption 3 and are symmetric.
-Taking $\tau=\frac{\max(\sigma,a)}{\log^{2}d}\sqrt{\frac{n}{m}}$, the following inequality holds with probability at least $1-\frac{1}{d}$,
-$\left\|\dfrac{1}{n}\sum_{i=1}^{n}\psi_{\tau}(\xi_{i})X_{i}\right\|\leq C\max(\sigma,a)\sqrt{\dfrac{1}{nm}}.$ (23)
-where $C$ is dependent on $L_{2}$ and $L_{3}$.
+Let Assumptions 1 and 3 hold. Suppose $n\geq 160m\log^{4}d$ and $\tau=\frac{\max(\sigma,a)}{\log^{2}d}\sqrt{\frac{n}{m}}$. If we take
+$\lambda\geq 3\left\|\frac{1}{n}\sum_{i=1}^{n}\phi_{\tau}(\xi_{i})X_{i}\right\|,$
+then the following bound holds with probability at least $1-\frac{1}{d}$,
+$\frac{\|\hat{A}_{H}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\lesssim\mu^{2}rm_{1}m_{2}(\lambda^{2}+a^{2}\mathbb{E}^{2}\|\mathcal{R}\|)+\frac{\mu^{2}\log d}{n}.$
+Here $\mathcal{R}$ is defined as $\mathcal{R}=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}},$ where $\chi_{i}=\{|\xi_{i}|\leq\tau/2\}$.
+Proof of Theorem 1.
+Recall $\tau=\frac{\max(a,\sigma)}{\log^{2}d}\sqrt{\frac{n}{m}}$. By (30), we obtain with probability at least $1-\frac{1}{d}$,
+$\left\|\frac{1}{n}\sum_{i=1}^{n}\phi_{\tau}(\xi_{i})X_{i}\right\|\leq C_{1}\max(\sigma,a)\sqrt{\dfrac{1}{nm}},$
+where $C_{1}$ depends on $L_{2}\sim L_{3}$. With this specific $C_{1}$, one can take $\lambda=3C_{1}\max(\sigma,a)\sqrt{\frac{1}{nm}}$ to apply Lemma 7. By (32), we obtain
+$\mathbb{E}\left\|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}\right\|\leq C_{2}\sqrt{\frac{1}{nm}},$ (11)
+where $C_{2}$ depends on $L_{2}\sim L_{3}$. Substituting this bound and the above choice of $\lambda$ into the inequality from Lemma 7, we complete the proof of Theorem 2. ∎
+Remark.
+Without the condition that $\xi_{i}$ are symmetric, $\nabla L(A_{0})$ may not have mean 0. In this case, it is necessary to control the spectral norm of its expectation. Specifically, we write:
+$\frac{1}{n}\sum_{i=1}^{n}\phi_{\tau}(\xi_{i})X_{i}=\frac{1}{n}\sum_{i=1}^{n}(\phi_{\tau}(\xi_{i})-\mathbb{E}[\phi_{\tau}(\xi_{i})|X_{i}])X_{i}+\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\phi_{\tau}(\xi_{i})|X_{i}]X_{i}.$
+Denote $E[\phi_{\tau}(X_{i})|X_{i}]$ by $u_{i}$ (we omit the dependence on $\tau$ to ease the notation).
+While (30) still applies to $\frac{1}{n}\sum_{i=1}^{n}(\phi_{\tau}(\xi_{i})-u_{i})X_{i}$ given $\tau=\max(\sigma,a)\sqrt{\frac{n}{m\log^{4}d}}$, we need to estimate
+$\left\|\frac{1}{n}\sum_{i=1}^{n}u_{i}X_{i}\right\|$ $=\left\|\dfrac{1}{n}\sum_{i=1}^{n}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])+\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[u_{i}X_{i}]\right\|$
+$\leq\left\|\dfrac{1}{n}\sum_{i=1}^{n}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])\right\|+\left\|\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[u_{i}X_{i}]\right\|.$
+Assume a random variable $Y$ has 0 mean and finite moment of order $2+\kappa$ with $\mathbb{E}^{\frac{1}{2+\kappa}}[|Y|^{2+\kappa}]\leq\sigma$ where $\kappa\geq 0$.
+By Cauchy-Schwarz inequality, for a positive number $\tau$,
+$|\mathbb{E}\phi_{\tau}(Y)|$ $=|\mathbb{E}[Y-\phi_{\tau}(Y)]|$
+$=|\mathbb{E}[(Y-\phi_{\tau}(Y)\textbf{1}(|Y|>\tau)]$
+$\leq\mathbb{E}^{1/p}[|Y|^{p}]\mathbb{P}^{1/q}(Y>\tau)$
+where $11/p+1/q=$ 1. Note $\mathbb{P}(Y>\tau)\leq\mathbb{E}[|Y|^{2+\kappa}]/\tau^{2+\kappa}$. Take $1/q=1-1/(2+\kappa)$. Then, we obtain
+$|\mathbb{E}[(Y-\phi_{\tau}(Y))\textbf{1}(|Y|>\tau)]|\leq\mathbb{E}^{\frac{1}{2+\kappa}}[|Y|^{2+\kappa}](\mathbb{E}^{\frac{1+\kappa}{2+\kappa}}[|Y|^{2+\kappa}]\tau^{-(1+\kappa)}).$
+Let $\tau=\max(\sigma,a)\sqrt{\frac{n}{m\log^{4}d}}$.
+By the above inequality, we obtain $|u_{i}|\leq\max(\sigma,a)(\frac{m\log^{4}d}{n})^{(1+\kappa)/2}$.
+Assuming $P_{jk}=\frac{1}{m_{1}m_{2}}$, we apply Theorem 7.3.1 in Tropp et al.
+[2015] to $\frac{1}{n}\sum_{i=1}^{n}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])$.
+Since $\frac{1}{n}\|u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}]\|$ is bounded by $\max(\sigma,a)\frac{2}{n}(\frac{m\log^{4}d}{n})^{(1+\kappa)/2}$, the following holds with probability at least $1-\frac{1}{d}$,
+$\left\|\frac{1}{n}\sum_{i=1}^{n}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])\right\|\lesssim\max(\sigma,a)\left(\frac{m\log^{4}d}{n}\right)^{(1+\kappa)/2}\sqrt{\frac{\log d}{nm}}+\max(\sigma,a)\frac{\log d}{n}\left(\frac{m\log^{4}d}{n}\right)^{(1+\kappa)/2},$
+where we use easy estimates $\|\mathbb{E}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])^{T}\|\lesssim\max(\sigma,a)\frac{L_{2}}{m}(\frac{m\log^{4}d}{n})^{(1+\kappa)/2}$ and $\|\mathbb{E}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])^{T}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])\|\lesssim\max(\sigma,a)\frac{L_{2}}{m}(\frac{m\log^{4}d}{n})^{(1+\kappa)/2}$.
+Given $n\geq m\log^{4+1/(1+\kappa)}d$, we have
+$\left\|\dfrac{1}{n}\sum_{i=1}^{n}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])\right\|\lesssim\sigma\sqrt{\frac{1}{nm}}.$
+Denote $u_{i}(j,k)=\mathbb{E}[\phi_{\tau}(\xi_{i})|X_{i}=e_{j}e^{T}_{k}]$. Note
+$\|\mathbb{E}[u_{i}X_{i}]\|$ $=\sup_{\|z\|_{2}=1,\|y\|_{2}=1}z^{T}\mathbb{E}[u_{i}X_{i}]y$
+$\leq\left(\sum_{j,k}P^{2}_{jk}\right)^{1/2}\left(\sum_{j,k}z_{j}^{2}y_{k}^{2}u^{2}_{i}(j,k)\right)^{1/2}$
+$\leq\max(\sigma,a)\sqrt{\frac{1}{m_{1}m_{2}}}\left(\frac{m\log^{4}d}{n}\right)^{(1+\kappa)/2}.$
+Given $Mn^{\kappa}\geq m\log^{4}d$, the RHS is less than $\max(\sigma,a)\sqrt{\frac{1}{nm}}$.
+An alternative condition is $n\geq\log^{4/\kappa}d$.
+When the noise only has second moment, namely $\kappa=0$, we can derive a $\log d$-free bound given $M\geq m\log^{4}d$ and $n\geq m\log^{5}d$.
+For non-symmetric noise, a $\log d$-free bound may be obtained under a better condition (i.e.
+$M\geq m\log^{\beta}d$ where $54\leq\beta<$ 5) by applying Theorem 13 to $\frac{1}{n}\sum_{i=1}^{n}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])$.
+We do not make further efforts here since it cannot remove the condition entirely.
+This technical requirement will possibly be not necessary with future refinement of the sharp concentration inequalities in Brailovskaya and van Handel [2024].
+Before presenting the poof of Lemma 7, we state a technical lemma.
 Lemma 8.
-If $n\geq m\log^{4}d$, we have the following result:
-$E\left\|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|\leq C\sqrt{\frac{1}{nm}},$ (24)
-where $C$ is dependent on $L_{2}$ and $L_{3}$.
-Proof of Lemma 12.
-Under the notation of Theorem 11, $Q_{i}=\frac{\xi_{i}X_{i}}{n}$.
-Note $\|Q_{i}\|=\frac{1}{n}|\xi_{i}|$.
-Since $R(Q)$ requires $\|Q_{i}\|$ to be bounded, we need to truncate $\xi_{i}$ to $\xi_{i}\textbf{1}(|\xi_{i}|\leq\tau)$ for some $\tau\geq 0$.
-Define $\bar{\xi}_{i}=\xi_{i}\textbf{1}(|\xi_{i}|\leq\tau)-E[\xi_{i}\textbf{1}(|\xi_{i}|\leq\tau)]$.
-Note $E[\bar{\xi}_{i}]=0$ and $E[\bar{\xi}_{i}^{2}]\leq\sigma^{2}$.
-We now compute the four parameters for $Q=\frac{1}{n}\sum_{i=1}^{n}\bar{\xi}_{i}X_{i}$.
-1.
-$\gamma(Q)\leq\sqrt{\frac{L_{2}\sigma^{2}}{nm}}$.
-Note $E[X_{i}X_{i}^{T}]$ is a diagonal matrix whose elements on the diagonal line are $\sum_{k=1}^{m_{2}}P_{jk},j=1,2,\ldots,m_{1}$, which are less than $\frac{L_{2}}{m}$.
-The same bound holds for $E[X_{i}^{T}X_{i}]$.
-Since
-$E[QQ^{T}]=\frac{1}{n^{2}}\sum_{i=1}^{n}E[\bar{\xi}_{i}^{2}X_{i}X_{i}^{T}]=\frac{1}{n}E[\bar{\xi}_{1}^{2}X_{1}X_{1}^{T}],$
-we have $\|E[QQ^{T}]\|\leq\frac{L_{2}\sigma^{2}}{nm}$. Similarly, $\|E[Q^{T}Q]\|\leq\frac{L_{2}\sigma^{2}}{nm}$. Thus $\gamma(Q)\leq\sqrt{\frac{L_{2}\sigma^{2}}{nm}}$.
-2. $\gamma_{*}(Q)\leq\sqrt{\frac{L_{3}\sigma^{2}}{nm\log^{3}d}}$. For $y\in\mathbb{R}^{m_{1}},z\in\mathbb{R}^{m_{2}}$ with $\|y\|_{2}=\|z\|_{2}=1$, we have the following result:
-$E(y^{T}\bar{\xi}_{i}X_{i}z)^{2}$ $=E(y^{T}X_{i}z)^{2}E[\bar{\xi}_{i}^{2}]$
-$=E[\bar{\xi}_{i}^{2}]\sum_{j,k}P_{jk}y^{2}(j)z^{2}(k)$
-$\leq\sigma^{2}\sum_{j,k}P_{jk}y^{2}(j)z^{2}(k)$
-$\leq\dfrac{L_{3}\sigma^{2}}{m\log^{3}d}.$
-where we use $\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}y^{2}(j)z^{2}(k)=\|y\|_{2}^{2}\cdot\|z\|_{2}^{2}=1$ and $\max P_{jk}\leq\frac{L_{3}}{m\log^{3}d}$.
-Since $\bar{\xi}_{i}X_{i}$ are independent and mean-zero, $E(y^{T}Qz)^{2}=\dfrac{1}{n^{2}}\sum_{i=1}^{n}E\left(y^{T}\bar{\xi}_{i}X_{i}z\right)^{2}\leq\frac{L_{3}\sigma^{2}}{nm\log^{3}d}$.
-Thus $\gamma_{*}(Q)\leq\sqrt{\frac{L_{3}\sigma^{2}}{nm\log^{3}d}}$.
-(3) $g(Q)\leq\sqrt{\frac{L_{3}\sigma^{2}}{nm\log^{3}d}}$.
-Since $\bar{\xi}_{i}X_{i}$’s are i.i.d., $\operatorname{Cov}(Q)=\dfrac{1}{n^{2}}\sum_{i=1}^{n}\operatorname{Cov}(\bar{\xi}_{i}X_{i})=\frac{1}{n}\operatorname{Cov}(\bar{\xi}_{1}X_{1})$.
-Define $\textbf{1}_{(j,k)}=\textbf{1}(X_{1}=e_{j}(m_{1})e_{k}(m_{2})^{T})$ and denote $I(j,k)=\bar{\xi}_{i}\textbf{1}_{(j,k)}$.
-Note the following facts
-$E[I(j,k)]=0,$
-$I(j,k)I(l,h)=0\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ (j,k)\neq(l,h).$
-Thus $\text{Cov}\big{(}I(j,k),I(l,h)\big{)}=E[I(j,k)I(l,h)]-E[I(j,k)]E[I(l,h)]=0$ for $(j,k)\neq(l,h)$. Then we have
-$\text{Cov}(\bar{\xi}_{i}X_{i})\Big{(}(j,k),(l,h)\Big{)}=\begin{cases}E[\bar{\xi}^{2}_{i}]P_{jk}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ \leavevmode\nobreak\ (j,k)=(l,h),\\ 0\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ (j,k)\neq(l,h).\end{cases}$
-Since $P_{jk}\leq\frac{L_{3}}{m\log^{3}d}$, $\|\operatorname{Cov}(\bar{\xi}_{i}X_{1})\|\leq\frac{L_{3}\sigma^{2}}{m\log^{3}d}$. Thus, $g(Q)\leq\sqrt{\frac{L_{3}\sigma^{2}}{nm\log^{3}d}}$ holds.
-4.$R(Q)\leq\frac{2\tau}{n}$. Obviously, $\|\bar{\xi}_{i}X_{i}\|=|\bar{\xi}_{i}|$, and $\max_{1\leq i\leq n}|\bar{\xi}_{i}|\leq 2\tau$.
-Now, we can plug the four estimates in (20). Taking $t=3\log d$, we get that with probability at least $1-\frac{1}{d^{2}}$, the following holds
-$\|Q\|$ $\leq 2\sqrt{\frac{\sigma^{2}}{nm}}$
-$+C\Bigg{(}\left(\frac{\sigma^{2}}{nm}\right)^{1/2}\cdot\left(\dfrac{\log^{3/4}d}{\log^{3/4}d}\right)+\left(\frac{\sigma^{2}}{nm}\right)^{1/2}\cdot\left(\dfrac{\log^{1/2}d}{\log^{3/2}d}\right)$
-$+\left(\frac{\log^{2}d\cdot\tau}{n}\right)^{1/3}\cdot\left(\frac{\sigma^{2}}{nm}\right)^{1/3}+\dfrac{\log d\cdot\tau}{n}\Bigg{)},$ (25)
-where $C$ is dependent on $L_{2}\sim L_{3}$. We can see the quantities in the first line and the second line of RHS are less than $C\sqrt{\frac{\sigma^{2}}{nm}}$.
-To prove the dimension free bound for $\frac{1}{n}\sum_{i=1}^{n}\psi_{\tau}(\xi_{i})X_{i}$, we consider $Q_{i}=\frac{\psi_{\tau}(\xi_{i})}{n}X_{i}$.
-Since $E[\psi_{\tau}(\xi_{i})]=0$, $E[\psi(\xi_{i})^{2}]\leq\sigma^{2}$, $|\psi_{\tau}(\xi_{i})|\leq\tau$, the estimates for the four parameters still apply.
-Then, with $\tau=\frac{\max(\sigma,a)}{\log^{2}d}\sqrt{\frac{n}{m}}$, we obtain
-$\left(\frac{\log^{2}d}{n}\cdot\frac{\max(\sigma,a)}{\log^{2}d}\sqrt{\frac{n}{m}}\right)^{1/3}\cdot\left(\frac{\sigma^{2}}{nm}\right)^{1/3}\leq\max(\sigma,a)\sqrt{\dfrac{1}{nm}},$
-$\dfrac{\log d}{n}\frac{\max(\sigma,a)}{\log^{2}d}\sqrt{\frac{n}{m}}\leq\max(\sigma,a)\sqrt{\dfrac{1}{nm}}.$
-Then the upper bound of $\|\frac{1}{n}\sum_{i=1}^{n}\psi_{\tau}(\xi_{i})X_{i}\|$ is free of the $\log d$ factor.
-For the unbounded noise case, we take $\tau=C_{1}\sigma\log n\cdot\log d$.
-To make the quantities in the last line be less than $C\sqrt{\frac{\sigma^{2}}{nm}}$, we need
-$\left(\frac{\sigma\log^{3}d\cdot\log n}{n}\right)^{1/3}\cdot\left(\frac{\sigma^{2}}{nm}\right)^{1/3}\leq\sqrt{\dfrac{\sigma^{2}}{nm}},$
-$\dfrac{\sigma\log^{2}d\cdot\log n}{n}\leq\sqrt{\dfrac{\sigma^{2}}{nm}}.$
-Those imply
-$n\geq m\log^{6}d\cdot\log^{2}n,$
-$n\geq m\log^{4}d\log^{2}n.$
-On $\{\max_{1\leq i\leq n}|\xi_{i}|\leq\tau\}$,
-$\left\|\dfrac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|$ $\leq\|Q\|+\left\|\dfrac{1}{n}\sum_{i=1}^{n}E[\xi_{i}\textbf{1}(|\xi_{i}|\leq\tau)]X_{i}\right\|$
-$\leq\|Q\|+|E[\xi_{1}\textbf{1}(|\xi_{1}|\leq\tau)]|.$
-Since $E[\xi_{i}]=0$, by Cauchy-Schwarz inequality, for $d$ sufficiently large,
-$|E\xi_{i}\textbf{1}(|\xi_{i}|\leq\tau)|$ $=|E\xi_{i}\textbf{1}(|\xi_{i}|>\tau)|$
-$\leq E|\xi_{i}|\textbf{1}(|\xi_{i}|>\tau)$
-$=\sigma E[|\xi_{i}|/\sigma]\textbf{1}(|\xi_{i}|/\sigma>C_{1}\log d\cdot\log n)$
-$\leq\sigma E^{1/2}[|\xi_{i}|^{2}/\sigma^{2}]\cdot P^{1/2}(|\xi_{i}|/\sigma>C_{1}\log d\cdot\log n)$
-$\lesssim\dfrac{L_{1}\sigma}{n^{C\log d}}$
-$\lesssim\frac{L_{1}\sigma}{n},$
-for a suitably chosen constant $C_{1}$ and sufficiently large $d$.
-$\frac{\sigma}{n}\leq\sqrt{\frac{\sigma^{2}}{nm}}$ holds under $n\geq m\log^{6}d\cdot\log^{2}n$.
-By the concentration inequality for the maximum of sub-Gaussian random variables (see Theorem 2.2.27 in [Talagrand, 2014]), $P(\max_{1\leq i\leq n}|\xi_{i}|\leq\tau)\geq 1-\frac{1}{d^{2}}$ for some suitably chosen $C_{1}$.
-Thus, combining the pieces, $\left\|\dfrac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\leq C\sigma\sqrt{\frac{1}{nm}}$ holds with probability at least $1-\frac{1}{d}$, where $C$ is dependent on $L_{1}\sim L_{3}$.
-∎
-Proof of Lemma 13.
-Let $Q=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}$. Since $\epsilon_{i}\leq 1$, we have $R(Q)=\frac{1}{n}$. And we have estimated the four parameters in the above. In the last line of (25), we need
-$\left(\frac{\log^{2}d}{n}\right)^{1/3}\cdot\left(\frac{1}{nm}\right)^{1/3}\leq C\sqrt{\dfrac{1}{nm}},$
-$\dfrac{\log d}{n}\leq C\sqrt{\dfrac{1}{nm}},$
-which requires $n\geq m\log^{4}d.$
-4.2 Proof of Theorem 1
-We first derive a general bound in Lemma 14, which reveals the role of the two key quantities $\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\|$ and $E\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\|$ in the error rate.
-Then combining it with the sharp spectral norm results for $\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\|$ and $E\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\|$ from Section 4.1 will complete the proof.
-Lemma 14.
-Let Assumption 1 and 2 hold. Denote $R=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}$. We take
-$\lambda\geq 3\left\|\frac{2}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|.$
-Then, the following bound holds with probability at least $1-\frac{1}{d}$,
-$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C\left(\mu^{2}(\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu^{2}a^{2}\log d}{n}\right),$ (26)
-where $C$ is a universal constant.
-Now we give the proof of Theorem 1.
-Proof of Theorem 1.
-By (22), we know that with probability at least $1-\frac{1}{d}$,
-$\left\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\leq C_{1}\sigma\sqrt{\dfrac{1}{nm}}.$
-where $C_{1}$ is dependent on $L_{1}\sim L_{3}$. Thus we can take $\lambda=6C_{1}\sigma\sqrt{\frac{1}{nm}}$. And by (24),
-$\left\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|\leq C_{2}\sqrt{\dfrac{1}{nm}},$
-where $C_{2}$ depends on $L_{2}\sim L_{3}$. Plugging the two above estimates in (26), we obtain, with probability at least $1-\frac{2}{d}$,
-$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\lesssim\dfrac{\mu^{2}\max(a^{2},\sigma^{2})rM}{n}+\frac{\mu^{2}a^{2}\log d}{n},$
-where the second term on RHS can be covered by the first term. ∎
-Proof of Lemma 14.
-Since $\hat{A}$ is the solution of (2), we have
-$\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},\hat{A}\rangle)^{2}+\|\hat{A}\|_{1}\leq\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A_{0}\rangle)^{2}+\|A_{0}\|_{1},$
-which implies
-$\frac{1}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}-A_{0}\rangle^{2}$ $\leq 2\left\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\|\hat{A}-A_{0}\|_{1}+\lambda(\|A_{0}\|_{1}-\|\hat{A}\|_{1})$
-$\leq\frac{4}{3}\lambda\|\hat{A}-A_{0}\|_{1}.$
-By Lemma 22, we know that
-$\|\hat{A}-A_{0}\|_{1}\leq 3\|P_{A_{0}}(\hat{A}-A_{0})\|_{1}\leq 3\sqrt{r}\|\hat{A}-A_{0}\|_{F},$ (27)
-which implies
-$\frac{1}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}-A_{0}\rangle^{2}\leq 5\lambda\sqrt{r}\|\hat{A}-A_{0}\|_{F}.$
-By Lemma 18 and $\|\hat{A}-A_{0}\|^{2}_{w(F)}\geq\frac{\|\hat{A}-A_{0}\|^{2}_{F}}{\mu m_{1}m_{2}}$, we obtain that with probability at least $1-\frac{1}{d}$,
-$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{\mu m_{1}m_{2}}\leq C\left(\dfrac{\mu\|\hat{A}-A_{0}\|_{\infty}^{2}\log d}{n}+\|\hat{A}-A_{0}\|_{\infty}\|\hat{A}-A_{0}\|_{1}E\|R\|+\lambda\sqrt{r}\|\hat{A}-A_{0}\|_{F}\right).$
-By $\|\hat{A}-A_{0}\|_{\infty}\leq 2a$, we have
-$\dfrac{\mu\|\hat{A}-A_{0}\|_{\infty}^{2}\log d}{n}\leq\frac{4\mu a^{2}\log d}{n}.$
-By $\|\hat{A}-A_{0}\|_{\infty}\leq 2a$ and (27), we obtain
-$\quad\|\hat{A}-A_{0}\|_{\infty}\|\hat{A}-A_{0}\|_{1}E\|R\|+\lambda\sqrt{r}\|\hat{A}-A_{0}\|_{F}$
-$\leq 2a\sqrt{r}\|\hat{A}-A_{0}\|_{F}E\|R\|+\frac{1}{10C}\frac{\|\hat{A}-A_{0}\|_{F}^{2}}{\mu m_{1}m_{2}}+10C\mu\lambda^{2}rm_{1}m_{2}$
-$\leq\frac{1}{5C}\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{\mu m_{1}m_{2}}+40C\mu a^{2}E^{2}\|R\|rm_{1}m_{2}+10C\mu\lambda^{2}rm_{1}m_{2}.$
-Combining the pieces above, we obtain the following bound holds with probability at least $1-\frac{1}{d}$
-$\frac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\lesssim\mu^{2}(\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\dfrac{a^{2}\mu^{2}\log d}{n}.$
-4.3 Proof of Theorem 3
-In this section, denote $R=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}$ and $L(A)=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}}$.
-We use the following general result to prove Theorem 3.
-Lemma 15.
-Let Assumptions 1 and 2 hold. Assume $\mu\lambda^{2}rm_{1}m_{2}\leq\frac{1}{10}$. We take
-$\lambda\geq\frac{3\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\|}{L(A_{0})}.$
-Then the following bound holds with probability at least $1-\frac{1}{d}$,
-$\frac{\|\hat{A}_{S}-A_{0}\|^{2}_{F}}{m_{1}m_{2}}\lesssim\mu^{2}(L^{2}(A_{0})\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu^{2}a^{2}\log d}{n}.$ (28)
-Proof of Theorem 3.
-Note $L(A_{0})=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\xi_{i}^{2}}$. By Theorem 3.1.1. in [Vershynin, 2018], there exists a universal constant $C$ such that
-$\left|\sqrt{\frac{1}{n}\sum_{i=1}^{n}\xi^{2}_{i}}-\sigma\right|\leq C\sigma\sqrt{\frac{L_{1}t}{n}}$
-holds with probability at least $1-2\exp(-t)$. Then, there exist a constant $C_{1}$ dependent on $L_{1}$ such that
-$0.9\sigma\leq\sqrt{\frac{1}{n}\sum_{i=1}^{n}\xi_{i}}\leq 1.1\sigma$
-holds with probability at least $1-2\exp(-C_{1}n)$. By (22), with probability at least $1-\frac{1}{d}$,
-$\left\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\leq C_{2}\sigma\sqrt{\dfrac{1}{nm}}.$
-Thus, we can take $\lambda=4C_{2}\sqrt{\frac{1}{nm}}$. And by (24),
-$E\left\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|\leq C_{3}\sqrt{\dfrac{1}{nm}}.$
-Here $C_{2}$ depends on $L_{1}\sim L_{3}$ and $C_{3}$ depends on $L_{2}\sim L_{3}$. Plugging in the above estimates, we obtain that
-$\dfrac{\|\hat{A}_{S}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\lesssim\dfrac{\mu^{2}\max(a^{2},\sigma^{2})rM}{n}+\frac{\mu^{2}a^{2}\log d}{n},$
-holds with probability at least $1-\frac{2}{d}-\exp(-C_{1}n)$. And the second term on RHS can be covered by the first term. ∎
-Proof of Lemma 15.
-The basic steps follow the proof in [Klopp, 2014]. Note
-$\frac{1}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}_{S}-A_{0}\rangle^{2}=L^{2}(\hat{A}_{S})-L^{2}(A_{0})+\frac{2}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}_{S}-A_{0}\rangle.$
-Since $L(\hat{A}_{S})$ is the solution of (8),
-$L(\hat{A}_{S})-L(A_{0})\leq\lambda(\|A_{0}\|_{1}-\|\hat{A}_{S}\|_{1}).$
-Immediately, we obtain
-$L^{2}(\hat{A}_{S})-L^{2}(A_{0})\leq(L(\hat{A}_{S})+L(A_{0}))(\lambda\|\hat{A}_{s}-A_{0}\|_{1}).$
-Again, by the definition of $\hat{A}_{S}$, we obtain
-$L(\hat{A})+L(A_{0})$ $\leq 2L(A_{0})+\lambda(\|A_{0}\|_{1}-\|\hat{A}_{S}\|_{1})\leq 2L(A_{0})+\lambda\|\hat{A}-A_{0}\|_{1}.$
-Gathering the pieces, we get
-$\frac{1}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}_{S}-A_{0}\rangle^{2}$ $\leq\frac{2}{n}\sum_{i=1}^{n}\langle\xi_{i}X_{i},\hat{A}_{S}-A_{0}\rangle+(2L(A_{0})+\lambda\|\hat{A}_{S}-A_{0}\|_{1})(\lambda\|\hat{A}-A_{0}\|_{1})$
-$\leq\frac{8L(A_{0})}{3}\lambda\|\hat{A}_{S}-A_{0}\|_{1}+\lambda^{2}\|\hat{A}_{S}-A_{0}\|_{1}^{2}.$
-By Lemma 18, we get with probability at least $1-\frac{1}{d}$,
-$\|\hat{A}_{S}-A_{0}\|^{2}_{w(F)}$ $\leq C\left(\|\hat{A}-A_{0}\|^{2}_{\infty}\frac{\mu\log d}{n}+\|\hat{A}_{S}-A_{0}\|_{\infty}\|\hat{A}_{S}-A_{0}\|_{1}E\|R\|\right)$
-$+\frac{8L(A_{0})}{3}\lambda\|\hat{A}_{S}-A_{0}\|_{1}+\lambda^{2}\|\hat{A}_{S}-A_{0}\|_{1}^{2}.$
-Since $\|\hat{A}_{S}-A_{0}\|_{\infty}\leq 2a$,
-$\|\hat{A}_{S}-A_{0}\|^{2}_{\infty}\frac{\mu\log d}{n}\leq\frac{4\mu a^{2}\log d}{n}.$
-By lemma (22), $\|\hat{A}_{S}-A_{0}\|_{1}\leq 3\|P_{A_{0}}(\hat{A}_{S}-A_{0})\|_{1}\leq 3\sqrt{r}\|\hat{A}_{S}-A_{0}\|_{F}$. Then we have
-$\|\hat{A}_{S}-A_{0}\|_{\infty}\|\hat{A}_{S}-A_{0}\|_{1}E\|R\|$ $\leq 6a\sqrt{r}E\|R\|\|\hat{A}_{S}-A_{0}\|_{F}$
-$\leq 360C\mu a^{2}E^{2}\|R\|rm_{1}m_{2}+\frac{1}{10C}\frac{\|\hat{A}_{S}-A_{0}\|_{F}^{2}}{\mu m_{1}m_{2}}.$
-Similarly,
-$\quad\left\|\frac{2}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\|\hat{A}_{S}-A_{0}\|_{1}+L(A_{0})\lambda\|\hat{A}_{S}-A_{0}\|_{1}$ $\leq\frac{5}{3}L(A_{0})\lambda\sqrt{r}\|\hat{A}_{S}-A_{0}\|_{F}$
-$\leq 40\mu L^{2}(A_{0})\lambda^{2}rm_{1}m_{2}+\frac{1}{10}\frac{\|\hat{A}_{S}-A_{0}\|_{F}^{2}}{\mu m_{1}m_{2}}.$
-Given $\mu\lambda^{2}rm_{1}m_{2}\leq\frac{1}{10}$,
-$\lambda^{2}\|\hat{A}_{S}-A_{0}\|_{1}^{2}\leq\frac{1}{10}\frac{\|\hat{A}_{S}-A_{0}\|_{F}^{2}}{\mu m_{1}m_{2}}.$
-Combining the pieces, we get
-$\|\hat{A}_{S}-A_{0}\|^{2}_{w(F)}\leq C\left(\mu^{2}(L^{2}(A_{0})\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu^{2}a^{2}\log d}{n}\right)+\frac{3}{10}\frac{\|\hat{A}_{S}-A_{0}\|_{F}^{2}}{\mu m_{1}m_{2}}.$
-Since $\|\hat{A}_{S}-A_{0}\|^{2}_{w(F)}\geq\frac{\|\hat{A}_{S}-A_{0}\|^{2}_{F}}{\mu m_{1}m_{2}}$, we obtain the following bound holds with probability at least $1-\frac{1}{d}$,
-$\frac{\|\hat{A}_{S}-A_{0}\|^{2}_{F}}{m_{1}m_{2}}\lesssim\mu^{2}(L^{2}(A_{0})\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu^{2}a^{2}\log d}{n}.$
-4.4 Proof of Theorem 5
-The proof is inspired by [Yu et al., 2024].
-Throughout this section, we set $L(A)=\frac{1}{n}\sum_{i=1}^{n}l_{\tau}(Y_{i}-\langle X_{i},A\rangle)$.
-In this section, $c_{3}$ and $c_{4}$ are specific constants those values do not change at each occurrence.
-We have the following general result.
-Lemma 16.
-Let Assumptions 2 and 3 hold. Let $n\geq 160m\log^{4}d$ and $\tau=\frac{\max(\sigma,a)}{\log^{2}d}\sqrt{\frac{n}{m}}$. If we take
-$\lambda\geq 3\left\|\frac{1}{n}\sum_{i=1}^{n}\psi_{\tau}(\xi_{i})X_{i}\right\|,$
-then the following bound holds with probability at least $1-\frac{1}{d}$,
-$\frac{\|\hat{A}_{H}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\lesssim\mu^{2}(\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu^{2}\log d}{n}.$
-Here $R$ is defined as $R=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}},$ where $\chi_{i}=\{|\xi_{i}|\leq\tau/2\}$.
-Proof of Theorem 5.
-We let $\tau=\frac{\max(a,\sigma)}{\log^{2}d}\sqrt{\frac{n}{m}}$. By (23), with probability at least $1-\frac{1}{d}$,
-$\left\|\frac{1}{n}\sum_{i=1}^{n}\psi_{\tau}(\xi_{i})X_{i}\right\|\leq C_{1}\max(\sigma,a)\sqrt{\dfrac{1}{nm}},$ (29)
-where $C_{1}$ depends on $L_{2}\sim L_{3}$. Since $E[\epsilon_{i}\textbf{1}_{\chi_{i}}]=0$ and $|\epsilon_{i}\textbf{1}_{\chi_{i}}|\leq 1$,by (24),
-$E\left\|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}\right\|\leq C_{2}\sqrt{\frac{1}{nm}},$ (30)
-where $C_{2}$ depends on $L_{2}\sim L_{3}$. Then all we need to do is to plug in the above two estimates in the bound in Lemma 16.
-In the above, if we do not assume $\xi_{i}$ are symmetric, we need to estimate
-$\left\|\frac{1}{n}\sum_{i=1}^{n}(E[\psi_{\tau}(\xi_{i})])X_{i}\right\|$ $=|E[\psi_{\tau}(\xi_{1})]|\left\|\dfrac{1}{n}\sum_{i=1}^{n}X_{i}\right\|$
-$\leq|E[\psi_{\tau}(\xi_{1})]|\left(\left\|\dfrac{1}{n}\sum_{i=1}^{n}(X_{i}-E[X_{i}])\right\|+\|E[X_{1}]\|\right).$
-By Cauchy-Schwarz inequality, $|E[\psi_{\tau}(\xi_{i})]|\leq\sqrt{E[4\xi_{i}^{2}]}\sqrt{P(|\xi_{i}|\geq\tau)}\leq 2\sigma\sqrt{\frac{m\log^{4}d}{n}}$.
-Assuming $P_{jk}=\frac{1}{m_{1}m_{2}}$, we apply Theorem 7.3.1 in[Tropp et al., 2015] to $\frac{1}{n}\sum_{i=1}^{n}(X_{i}-E[X_{i}])$.
-Since $\frac{1}{n}\|(X_{i}-E[X_{i}])\|$ is bounded by $\frac{2}{n}$, we obtain the following holds with probability at least $1-\frac{1}{d}$,
-$\left\|\frac{1}{n}\sum_{i=1}^{n}(X_{i}-E[X_{i}])\right\|\lesssim\sqrt{\frac{\log d}{nm}}+\frac{\log d}{n},$
-where we use easy estimates $\|E(X_{i}-E[X_{i}])(X_{i}-E[X_{i}])^{T}\|\lesssim\frac{L_{2}}{m}$ and $\|E(X_{i}-E[X_{i}])^{T}(X_{i}-E[X_{i}])\|\lesssim\frac{L_{2}}{m}$. Given $n\geq m\log^{5}d$, we have
-$|E[\psi_{\tau}(\xi_{1})]|\left\|\dfrac{1}{n}\sum_{i=1}^{n}(X_{i}-E[X_{i}])\right\|\lesssim\sigma\sqrt{\frac{1}{nm}}.$
-All the entries of $E[X_{i}]$ are $\frac{1}{m_{1}m_{2}}$, then we have $\|E[X_{i}]\|\leq\frac{1}{\sqrt{m_{1}m_{2}}}$.
-Note $\sqrt{\frac{m\log^{4}d}{n}}\cdot\frac{1}{\sqrt{m_{1}m_{2}}}\leq\sqrt{\frac{1}{nm}}$ holds only when $M\geq m\log^{4}d$.
-∎
+With $\lambda$ in (4) satisfying $\lambda\geq 3\|\frac{1}{n}\sum_{i=1}^{n}\phi_{\tau}(\xi_{i})X_{i}\|$ , we can obtain
+$\|P_{A_{0}}^{\perp}(\hat{A}_{H}-A_{0})\|_{*}\leq 2\|P_{A_{0}}(\hat{A}_{H}-A_{0})\|_{*}.$
+Proof.
+As a solution of a convex program, by Proposition 1.3 in Bubeck [2014] , $\hat{A}_{H}$ satisfies:
+$\langle\nabla L(\hat{A}_{H})+\lambda Z,\bar{A}-A_{0}\rangle\leq 0.$
+where $Z\in\partial\|\hat{A}_{H}\|_{*}$. Since $\|\cdot\|_{*}$ is convex, we have
+$\langle Z,A_{0}-\hat{A}_{H}\rangle\leq\|A_{0}\|_{*}-\|\hat{A}_{H}\|_{*}.$
+By the convexity of $L(\cdot)$,
+$\langle\nabla L(\hat{A}_{H})-\nabla L(A_{0}),\hat{A}_{H}-A_{0}\rangle\geq 0.$
+Combining the results above, we obtain:
+$0\leq\langle\nabla L(\hat{A}_{H})-\nabla L(A_{0}),\hat{A}_{H}-A_{0}\rangle\leq\lambda(\|A_{0}\|_{*}-\|\hat{A}_{H}\|_{*})+\|\nabla L(A_{0})\|\|\hat{A}_{H}-A_{0}\|_{*}.$ (12)
+Note
+$\|\hat{A}_{H}\|_{*}$ $=\|P_{A_{0}}(\hat{A}_{H}-A_{0})+P_{A_{0}}^{\perp}(\hat{A}_{H}-A_{0})+P_{A_{0}}(A_{0})\|_{*}$
+$\geq\|P_{A_{0}}^{\perp}(\hat{A}_{H}-A_{0})\|_{*}+\|P_{A_{0}}(A_{0})\|_{*}-\|P_{A_{0}}(\hat{A}_{H}-A_{0})\|_{*}.$
+Then, we get
+$0\leq\frac{4}{3}\|P_{A_{0}}(\hat{A}_{H}-A_{0})\|_{*}-\frac{2}{3}\|P_{A_{0}}^{\perp}(\hat{A}_{H}-A_{0})\|_{*},$
+which leads to the statement of the lemma. x∎
 Proof Lemma 7.
 As a solution of the convex program (4), by (12), $\hat{A}_{H}$ satisfies:
-$\langle\nabla L(\hat{A}_{H})-L(A_{0}),\hat{A}_{H}-A_{0}\rangle\leq\frac{4}{3}\|\hat{A}_{H}-A_{0}\|_{1}.$
+$\langle\nabla L(\hat{A}_{H})-L(A_{0}),\hat{A}_{H}-A_{0}\rangle\leq\frac{4}{3}\lambda\|\hat{A}_{H}-A_{0}\|_{*}.$
+From Wainwright [2019] Chapter 10, we know, for a matrix $B$,
+$\|P_{A_{0}}(B)\|_{*}\leq\sqrt{2\text{rank}(A_{0})}\|B\|_{F}.$
 By Lemma 8,
-$\|\hat{A}_{H}-A_{0}\|_{1}\leq 3\|P_{A_{0}}(\hat{A}_{H}-A_{0})\|_{1}\leq 3\sqrt{r}\|\hat{A}_{H}-A_{0}\|_{F}.$
+$\|\hat{A}_{H}-A_{0}\|_{*}\leq 3\|P_{A_{0}}(\hat{A}_{H}-A_{0})\|_{*}\leq 3\sqrt{2r}\|\hat{A}_{H}-A_{0}\|_{F}.$
 By Lemma 9, assuming $\|\hat{A}_{H}-A_{0}\|^{2}_{w(F)}\leq s^{2}$, there is a constant $c_{3}$ such that the following holds with probability at least $1-\frac{1}{d}$,
-$\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\left(\frac{9}{10}-\dfrac{4\sigma^{2}}{\tau^{2}}\right)\leq 4\sqrt{r}\lambda\|\hat{A}_{H}-A_{0}\|_{F}+\dfrac{64a^{2}s^{2}}{\tau^{2}}$
-$+c_{3}\Bigg{(}\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{1}E\|R\|+\frac{\mu a^{2}\log d}{n}\Bigg{)},$ (31)
-where we use $\|\hat{A}_{H}-A_{0}\|_{\infty}\leq 2a$. On the RHS of (31), we have
+$\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\left(\frac{9}{10}-\dfrac{4\sigma^{2}}{\tau^{2}}\right)\leq 4\sqrt{2r}\lambda\|\hat{A}_{H}-A_{0}\|_{F}+\dfrac{64a^{2}s^{2}}{\tau^{2}}$
+$+c_{3}\Bigg{(}\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{*}\mathbb{E}\|R\|+\frac{\mu a^{2}\log d}{n}\Bigg{)},$ (13)
+where we use $\|\hat{A}_{H}-A_{0}\|_{\infty}\leq 2a$. On the RHS of (13), it holds
 $\lambda\sqrt{r}\|\hat{A}_{H}-A_{0}\|_{F}\leq 40\mu\lambda^{2}rm_{1}m_{2}+\frac{1}{40}\frac{\|\hat{A}_{H}-A_{0}\|^{2}_{F}}{\mu m_{1}m_{2}}.$
-By $\|\hat{A}_{H}-A_{0}\|_{\infty}\leq 2a$, we get
-$\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{1}E\|R\|$ $\leq 2a\sqrt{r}\|\hat{A}_{H}-A_{0}\|_{F}E\|R\|$
-$\leq 40c_{3}\mu a^{2}E^{2}\|R\|rm_{1}m_{2}+\frac{1}{10c_{3}}\frac{\|\hat{A}_{H}-A_{0}\|^{2}_{F}}{\mu m_{1}m_{2}}.$
+By $\|\hat{A}_{H}-A_{0}\|_{\infty}\leq 2a$, we obtain
+$\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{*}\mathbb{E}\|R\|$ $\leq 2a\sqrt{r}\|\hat{A}_{H}-A_{0}\|_{F}\mathbb{E}\|R\|$
+$\leq 40c_{3}\mu a^{2}\mathbb{E}^{2}\|R\|rm_{1}m_{2}+\frac{1}{10c_{3}}\frac{\|\hat{A}_{H}-A_{0}\|^{2}_{F}}{\mu m_{1}m_{2}}.$
 Then we obtain
-$\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\left(\frac{9}{10}-\dfrac{4\sigma^{2}}{\tau^{2}}\right)\leq 160\mu\lambda^{2}rm_{1}m_{2}+\frac{1}{5}\frac{\|\hat{A}_{H}-A_{0}\|^{2}_{F}}{\mu m_{1}m_{2}}+\dfrac{64a^{2}s^{2}}{\tau^{2}}$
-$+c_{3}\Bigg{(}40c_{3}\mu a^{2}E^{2}\|R\|rm_{1}m_{2}+\frac{4\mu a^{2}\log d}{n}\Bigg{)},$
-If $n\geq 160m\log^{4}d$, we have $\frac{4\sigma^{2}}{\tau^{2}}\leq\frac{1}{10}$ and $\frac{64a^{2}}{\tau^{2}}\leq\frac{2}{5}$.
-Then $\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\left(\frac{9}{10}-\frac{4\sigma^{2}}{\tau^{2}}\right)\geq\frac{4}{5}\|\hat{A}_{H}-A_{0}\|^{2}_{w(F)}$.
+$\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\left(\frac{9}{10}-\dfrac{4\sigma^{2}}{\tau^{2}}\right)\leq 160\sqrt{2}\mu\lambda^{2}rm_{1}m_{2}+\frac{1}{5}\frac{\|\hat{A}_{H}-A_{0}\|^{2}_{F}}{\mu m_{1}m_{2}}+\dfrac{64a^{2}s^{2}}{\tau^{2}}$
+$+c_{3}\Bigg{(}40c_{3}\mu a^{2}\mathbb{E}^{2}\|R\|rm_{1}m_{2}+\frac{4\mu a^{2}\log d}{n}\Bigg{)},$
+If $n\geq 160m\log^{4}d$, we have $\frac{4\sigma^{2}}{\tau^{2}}\leq\frac{1}{10}$ and $\frac{64a^{2}}{\tau^{2}}\leq\frac{2}{5}$, which leads to $\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\left(\frac{9}{10}-\frac{4\sigma^{2}}{\tau^{2}}\right)\geq\frac{4}{5}\|\hat{A}_{H}-A_{0}\|^{2}_{w(F)}$.
 Then we obtain that there is a constant $c_{4}$ such that the following holds with probability at least $1-\frac{1}{d}$,
-$\frac{3}{5}\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\leq c_{4}\left(\mu(\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu\log d}{n}\right)+\frac{2}{5}s^{2}.$
-This implies
-$\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\leq\frac{5c_{4}}{3}\left(\mu(\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu\log d}{n}\right)+\frac{2}{3}s^{2}.$
-Denote $\delta=\frac{5c_{4}}{3}\left(\mu(\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu\log d}{n}\right)$.
+$\frac{3}{5}\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\leq c_{4}\left(\mu(\lambda^{2}+a^{2}\mathbb{E}^{2}\|R\|)rm_{1}m_{2}+\frac{\mu\log d}{n}\right)+\frac{2}{5}s^{2}.$
+The above implies
+$\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\leq\frac{5c_{4}}{3}\left(\mu(\lambda^{2}+a^{2}\mathbb{E}^{2}\|R\|)rm_{1}m_{2}+\frac{\mu\log d}{n}\right)+\frac{2}{3}s^{2}.$
+Denote $\delta=\frac{5c_{4}}{3}\left(\mu(\lambda^{2}+a^{2}\mathbb{E}^{2}\|R\|)rm_{1}m_{2}+\frac{\mu\log d}{n}\right)$.
 We take $s^{2}=4\delta$.
@@
 Note that $\tilde{A}_{H}$ also satisfies
-$\|\tilde{A}_{H}-A_{0}\|_{1}\leq 3\sqrt{r}\|\tilde{A}_{H}-A_{0}\|_{F}.$
-By Lemma F.2 in [Fan et al., 2018], we have
+$\|\tilde{A}_{H}-A_{0}\|_{*}\leq 3\sqrt{r}\|\tilde{A}_{H}-A_{0}\|_{F}.$
+By Lemma F.2 in Fan et al. [2018], we have
 $\langle\nabla L(\tilde{A}_{H})-\nabla L(A_{0}),\tilde{A}_{H}-A_{0}\rangle$ $\leq\eta\langle\nabla L(\hat{A}_{H})-\nabla L(A_{0}),\hat{A}_{H}-A_{0}\rangle$
-$\leq 4\lambda\|P_{A_{0}}(\tilde{A}-A_{0})\|_{1}$
+$\leq 4\lambda\|P_{A_{0}}(\tilde{A}-A_{0})\|_{*}$
 Then, we can repeat the previous steps and Lemma 9 also applies to $\tilde{A}_{H}-A_{0}$. Finally we get that
-$\|\tilde{A}_{H}-A_{0}\|_{w(F)}^{2}\leq\delta+\frac{2}{3}s^{2}\leq\frac{11}{12}s^{2}.$
-This contradicts with that $\|\tilde{A}_{H}-A_{0}\|_{w(F)}^{2}=s^{2}$. Thus, we get
-$\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\leq\frac{20c_{4}}{3}\left(\mu(\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu\log d}{n}\right)$
+$\|\tilde{A}_{H}-A_{0}\|_{w(F)}^{2}\leq\delta+\frac{2}{3}s^{2}\leq\frac{11}{12}s^{2},$
+which contradicts with that $\|\tilde{A}_{H}-A_{0}\|_{w(F)}^{2}=s^{2}$. Thus, we get
+$\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\leq\frac{20c_{4}}{3}\left(\mu(\lambda^{2}+a^{2}\mathbb{E}^{2}\|\mathcal{R}\|)rm_{1}m_{2}+\frac{\mu\log d}{n}\right)$
 and
-$\frac{\|\hat{A}_{H}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\lesssim\mu^{2}(\lambda^{2}+a^{2}E^{2}\|R\|)rm_{1}m_{2}+\frac{\mu^{2}\log d}{n}.$
-4.5 Proof of Theorem 7 and 9
-Based on [Klopp et al., 2017], (12) and (16) can be summarized in the same form
-$(\hat{L},\hat{S})\in\underset{\begin{subarray}{c}\|L\|_{\infty}\leq\mathrm{a}\\ \|S\|_{\infty}\leq\mathrm{a}\end{subarray}}{\arg\min}\left\{\frac{1}{N}\sum_{i=1}^{N}\left(Y_{i}-\left\langle X_{i},L+S\right\rangle\right)^{2}+\lambda_{1}\|L\|_{1}+\lambda_{2}\mathcal{R}(S)\right\}.$
-where $\mathcal{R}$ is certain matrix norm for $S$.
-Denote $\mathcal{R}^{*}(B)=\sup_{\mathcal{R}(A)\leq 1}\langle A,B\rangle$.
-[Klopp et al., 2017] offers the following general results for robust matrix completion.
-Theorem 17.
-Let Assumptions 1 and 4 be satisfied.
-Let $\lambda_{2}\geq 4\left(\mathcal{R}^{*}(\frac{1}{N}\sum_{i\in\Omega}\xi_{i}X_{i})+2a\mathcal{R}^{*}(\frac{1}{N}\sum_{i\in\Omega}X_{i})\right)$, and $\lambda_{1}>4\|\frac{1}{N}\sum_{i\in\Omega}\xi_{i}X_{i}\|$.
-Then, with probability at least $1-\frac{4.5}{d}$,
-$\frac{\left\|\hat{L}-L_{0}\right\|_{2}^{2}}{m_{1}m_{2}}+\frac{\left\|\hat{S}-S_{0}\right\|_{2}^{2}}{m_{1}m_{2}}\leq C\left\{\Psi_{1}+\Psi_{2}+\Psi_{3}\right\},$
-where $C$ is an absolute constant, and $\Psi_{1}$ is defined as
-$\Psi_{1}=\mu^{2}m_{1}m_{2}r\left(\alpha^{2}\lambda_{1}^{2}+\mathbf{a}^{2}\left(E\left\|\frac{1}{n}\sum_{i\in\Omega}\epsilon_{i}X_{i}\right\|\right)^{2}\right)+\mathbf{a}^{2}\mu\sqrt{\frac{\log(d)}{n}}.$
-Here, we do not need to care about the definitions of $\Psi_{2}$ and $\Psi_{3}$ since they are unrelated to $O(\frac{rM}{n})$ part in the final bounds.
-The term $\frac{rM\log d}{n}$ only results from $\Psi_{1}$.
-Note $\frac{1}{N}\sum_{i\in\Omega}\xi_{i}X_{i}=\frac{1}{\alpha n}\sum_{i\in\Omega}\xi_{i}X_{i}$.
-By (22) and (24), we can obtain
-$\left\|\dfrac{1}{N}\sum_{i\in\Omega}\xi_{i}X_{i}\right\|\leq C\frac{\sigma}{\alpha}\sqrt{\dfrac{1}{nm}}=C\sigma\sqrt{\frac{1}{\alpha Nm}}$
-with probability at least $1-\frac{1}{d}$ and
-$E\left\|\dfrac{1}{n}\sum_{i\in\Omega}\epsilon_{i}X_{i}\right\|\leq C\sqrt{\dfrac{1}{nm}}.$
-Plug the above two bounds in $\Psi_{1}$, then we remove the $\log d$ factor.
-4.6 Proof of restricted strong convexity for $L^{2}$ loss function
-In this section, $c_{0}\sim c_{4}$ are positive constants whose values do not change at each occurrence.
-Lemma 18.
-Recall $\|A\|^{2}_{w(F)}=\sum_{i,j}a_{ij}^{2}P_{ij}$.
-Assume $n\geq m\log^{6}d\log^{2}n$.
-For a non-zero matrix $A$, the following inequality holds, uniformly with probability at least $1-\dfrac{1}{d}$,
-$\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},A\rangle^{2}\geq\dfrac{3}{4}\|A\|_{w(F)}^{2}-c_{3}\left(\|A\|_{\infty}^{2}\dfrac{\mu\log d}{n}+\|A\|_{\infty}\|A\|_{1}E\|R\|\right),$ (32)
-where $c_{3}$ is a universal positive constant and $R$ is defined as
-$R=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}.$
-Proof of Lemma 18.
-At first, we introduce a concentration inequality for empirical processes.
-Lemma 19 (Theorem 3.27 in [Wainwright, 2019]).
-Let $Z_{i},1\leq i\leq n$, be independent but not necessarily identically distributed random variables taking values in measurable spaces $\mathcal{Z},1\leq i\leq n$.
-$\mathscr{F}$ is a function class.
-Consider the random variable
-$Z=\sup_{f\in\mathscr{F}}\frac{1}{n}\sum_{i=1}^{n}f(Z_{i}).$
-If $\|f\|_{\infty}\leq b$ for all $f\in\mathscr{F}$, we have the following inequality, for $t>0$,
-$P(Z\geq c_{0}E[Z]+c_{1}\tilde{\sigma}\sqrt{t}+c_{2}bt)\leq\exp\left(-nt\right)$ (33)
-where $\tilde{\sigma}^{2}=\sup_{f\in\mathscr{F}}\frac{1}{n}\sum_{i=1}^{n}Ef^{2}(Z_{i})$ and $c_{0},c_{1},c_{2}$ are some universal constants.
-Without loss of generality, we can assume $\|A\|_{w(F)}=1$. For a pair of fixed parameters $(\alpha,\rho)$, define:
-$S(\alpha,\rho)=\{A\in\mathbb{R}^{m_{1}\times m_{2}};\|A\|_{w(F)}=1,\|A\|_{\infty}\leq\alpha,\|A\|_{1}\leq\rho\}.$
-Introducing the convenient short hand notation $F_{A}(X)=\langle X,A\rangle^{2}$, we define
-$Z(\alpha,\rho)=\sup_{A\in S(\alpha,\rho)}\left|\dfrac{1}{n}\sum_{i=1}^{n}F_{A}(X_{i})-\|A\|_{w(F)}^{2}\right|.$
-For all $A\in S(\alpha,\rho)$, $|F_{A}(X_{i})|\leq\alpha^{2}$ for all $1\leq i\leq n$. Thus, we obtain
-$\operatorname{Var}\left(F_{A}(X_{i})\right)$ $=E[F_{A}(X_{i})^{2}]-\left(EF_{A}(X_{i})\right)^{2}$
-$\leq E[F_{A}(X_{i})^{2}]$
-$\leq\alpha^{2}\|A\|^{2}_{w(F)}.$
-Then we have $\tilde{\sigma}^{2}\leq\alpha^{2}$. Applying the concentration inequality (33) for the empirical processes with $t=\frac{2\mu\log d}{N}$, we get
-$P\left(Z(\alpha,\rho)\geq c_{0}E[Z(\alpha,\rho)]+\sqrt{2}c_{1}\alpha\sqrt{\frac{\mu\log d}{n}}+2c_{2}\frac{\alpha^{2}\mu\log d}{n}\right)$ $\leq\exp\left(-2\mu\log d\right).$
-Now we estimate $E[Z(\alpha,\rho)]$. By Rademacher symmetrization and Ledoux-Talagrand contraction (see [Wainwright, 2019] Proposition 4.11) , we get
-$E[Z(\alpha,\rho)]\leq 2E\sup_{A\in S(\alpha,\rho)}\Big{|}\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\langle X_{i},A\rangle^{2}\Big{|}\leq 4\alpha E\sup_{A\in S(\alpha,\rho)}\Big{|}\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\langle X_{i},A\rangle\Big{|}.$
-Since $\|A\|_{1}\leq\rho$,
-$E\sup_{A\in S(\alpha,\rho)}\Big{|}\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\langle X_{i},A\rangle\Big{|}\leq\rho E\left\|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|=\rho E\|R\|,$
-where we use $|\langle A,B\rangle|\leq\|A\|\|B\|_{1}$. Thus, we have
-$P\left(Z(\alpha,\rho)\geq 4c_{0}\alpha\rho E\|R\|+\sqrt{2}c_{1}\alpha\sqrt{\frac{\mu\log d}{n}}+2c_{2}\frac{\mu\alpha^{2}\log d}{n}\right)\leq\frac{1}{d^{2\mu}}.$ (34)
-Then we use the peeling argument to prove that the following:
-$\left|\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},A\rangle^{2}-\|A\|_{w(F)}^{2}\right|\leq 16c_{0}\|A\|_{\infty}\|A\|_{1}E\|R\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\dfrac{\mu\log d}{n}}+8c_{2}\frac{\mu\|A\|^{2}_{\infty}\log d}{n}$ (35)
-holds uniformly with probability at least $1-\frac{1}{d}$. Let $\mathcal{E}$ be the event that the bound (35) is violated for some $A$ with $\|A\|_{w(F)}=1$. Note that under Assumption 2,
-$\|A\|_{F}\leq\sqrt{\mu m_{1}m_{2}}\|A\|_{w(F)}=\sqrt{\mu m_{1}m_{2}}.$
-and
-$\|A\|_{F}\geq\|A\|_{w(F)}=1.$
-Then, we have
-$\|A\|_{1}\leq\sqrt{m}\|A\|_{F}\leq\sqrt{\mu d^{3}}$
-and
-$1\leq\|A\|_{F}\leq\|A\|_{1}.$
-Without loss of generality, we can assume $\|A\|\in[1,\sqrt{\mu d^{3}}]$. Similarly, we have
-$\|A\|_{\infty}\leq\|A\|_{F}\leq\sqrt{\mu m_{1}m_{2}}$
-and
-$\|A\|_{\infty}\geq\frac{\|A\|_{F}}{\sqrt{m_{1}m_{2}}}\geq\frac{1}{d}.$
-For an integer pair $(k,l)$, define
-$\tilde{S}(k,l)=\{A;\|A\|_{w(F)}=1,2^{k-1}\leq\|A\|_{\infty}\leq 2^{k},\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ 2^{l-1}\leq\|A\|_{1}\leq 2^{l}\}$
-Every matrix $A$ with $\|A\|_{w(F)}=1$ belongs to one $\tilde{S}(k,l)$ with $k=\lfloor\log_{2}\frac{1}{d}\rfloor,\ldots,\lceil\frac{1}{2}\log_{2}(\mu d)\rceil$ and $l=1,\ldots,\lceil\frac{1}{2}\log_{2}(\mu d^{3})\rceil$.
-Thus the total number of the pairs $(k,l)$ is less than $H^{2}=9\log_{2}^{2}(\mu d)$.
-Let $\mathcal{E}_{k,l}$ be the event that the bound (35) is violated for some $A\in\tilde{S}_{k,l}$.
-It can be seen that
-$\mathcal{E}\subseteq\bigcup_{1\leq k\leq H\atop 1\leq l\leq H}\mathcal{E}_{kl}.$
-Thus, if there exists a matrix $A$ with $\|A\|_{w(F)}=1$ that violates the bound (35), it must belong to some set $\tilde{S}_{k,l}$. Next, for $\rho=2^{l},\alpha=2^{k}$, define the event
-~$\widetilde{\mathcal{E}}_{k,l}:=\left\{Z(\alpha,\rho)\geq 4c_{0}\alpha\rho E\|R\|+\sqrt{2}c_{1}\alpha\sqrt{\dfrac{\mu\log d}{n}}+4c_{2}\frac{\mu\alpha^{2}\log d}{n}\right\}.$
-We claim that $\mathcal{E}_{k,l}\subseteq\widetilde{\mathcal{E}}_{k,l}$. Indeed, if event $\mathcal{E}_{k,l}$ occurs, then there must exist some $A\in\tilde{S}_{k,l}$ such that
-$\left|\frac{1}{n}\sum_{i=1}^{n}\langle X_{i},A\rangle^{2}-\|A\|_{w(F)}^{2}\right|$ $\geq 16c_{0}\|A\|_{\infty}\|A\|_{1}E\|R\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\dfrac{\mu\log d}{n}}+16c_{2}\dfrac{\|A\|^{2}_{\infty}\mu\log d}{n}$
-$\geq 16c_{0}2^{k-1}2^{l-1}E\|R\|+2\sqrt{2}c_{1}2^{k-1}\sqrt{\dfrac{\mu\log d}{n}}+16c_{2}2^{2(k-1)}\frac{\mu\log d}{n}$
-$\geq\tilde{4}c_{0}2^{k}2^{l}E\|R\|+\sqrt{2}c_{1}2^{k}\sqrt{\dfrac{\mu\log d}{n}}+4c_{2}2^{2k}\frac{\mu\log d}{n},$
-showing that ~$\widetilde{\mathcal{E}}_{k,l}$ occurs. Putting together the pieces, we have, for $d$ sufficiently large,
-$P(\mathcal{E})\stackrel{{\scriptstyle\text{ (i) }}}{{\leq}}\sum_{1\leq k\leq H\atop 1\leq l\leq H}P\left(\widetilde{\mathcal{E}}_{k,l}\right)\stackrel{{\scriptstyle\text{ (ii) }}}{{\leq}}\frac{H^{2}}{d^{2\mu}}\leq\frac{9\log^{2}\mu d}{d^{\mu}}\cdot\frac{1}{d^{\mu}}\leq\dfrac{1}{d},$
-where inequality (i) follows from the union bound applied to the inclusion $\mathcal{E}\subseteq\bigcup_{k,\ell=1}^{H}\widetilde{\mathcal{E}}_{k,l}$; inequality (ii) is a consequence of the tail bound (34).
-So far, we have proved (35).
-In the above argument, we normalize the matrix by its $\|\cdot\|_{w(F)}$ norm.
-Namely we consider $\frac{A}{\|A\|_{w(F)}}$, whose infinity norm and nuclear norm in (35) are $\dfrac{\|A\|_{\infty}}{\|A\|_{w(F)}}$ and $\dfrac{\|A\|_{1}}{\|A\|_{w(F)}}$.
-Multiplying $\|A\|_{w(F)}$ on both sides of formula (35), we obtain that with probability at least $1-\frac{1}{d}$,
-$\left|\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},A\rangle^{2}-\|A\|_{w(F)}^{2}\right|$ $\leq 16c_{0}\|A\|_{\infty}\|A\|_{1}\sqrt{\dfrac{1}{nm}}+2\sqrt{2}c_{1}\|A\|_{\infty}\|A\|_{w(F)}\sqrt{\dfrac{\mu\log d}{n}}$
-$+16c_{2}\|A\|^{2}_{\infty}\frac{\mu\log d}{n}.$
-which implies
-$\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},A\rangle^{2}$ $\geq\|A\|_{w(F)}^{2}-16c_{0}\|A\|_{\infty}\|A\|_{1}E\|R\|-\frac{\|A\|_{w(F)}^{2}}{4}-(8c_{1}^{2}+4c_{2})\|A\|^{2}_{\infty}\frac{\mu\log d}{n}$
-$=\frac{3}{4}\|A\|_{w(F)}^{2}-16c_{0}\|A\|_{\infty}\|A\|_{1}E\|R\|-(8c_{1}^{2}+4c_{2})\|A\|^{2}_{\infty}\frac{\mu\log d}{n}.$
+$\frac{\|\hat{A}_{H}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\lesssim\mu^{2}(\lambda^{2}+a^{2}\mathbb{E}^{2}\|\mathcal{R}\|)rm_{1}m_{2}+\frac{\mu^{2}\log d}{n}.$
 4.2 Proof of restricted strong convexity for Huber loss function
-In this section, $c_{0}\sim c_{3}$ are positive constants whose values do not change at each occurrence, $L(A)=\frac{1}{n}\sum_{i=1}^{n}l_{\tau}(Y_{i}-\langle X_{i},A\rangle)$.
 Lemma 9.
@@
 $\langle\nabla L(\hat{A}_{H})-\nabla L(A_{0}),\hat{A}_{H}-A_{0}\rangle$ $\geq\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\left(\frac{9}{10}-\dfrac{4\sigma^{2}}{\tau^{2}}\right)-\dfrac{16\|\hat{A}_{H}-A_{0}\|^{2}_{\infty}s^{2}}{\tau^{2}}$
-$-c_{3}\Bigg{(}\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{1}E\|R\|+\frac{\mu\|\hat{A}_{H}-A_{0}\|_{\infty}^{2}\log d}{n}\Bigg{)},$ (36)
+$-c_{3}\Bigg{(}\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{*}\mathbb{E}\|R\|+\frac{\mu\|\hat{A}_{H}-A_{0}\|_{\infty}^{2}\log d}{n}\Bigg{)},$ (14)
 where $c_{3}$ is a universal constant.
@@
 $\quad\langle\nabla L(\hat{A}_{H})-\nabla L(A_{0}),\hat{A}_{H}-A_{0}\rangle$
-$=\frac{1}{n}\sum_{i=1}^{n}(\psi_{\tau}(Y_{i}-\langle X_{i},A_{0}\rangle)-\psi_{\tau}(Y_{i}-\langle X_{i},\hat{A}_{H}\rangle))\langle X_{i},\hat{A}_{H}-A_{0}\rangle$
-$\geq\frac{1}{n}\sum_{i=1}^{n}(\psi_{\tau}(Y_{i}-\langle X_{i},A_{0}\rangle)-\psi_{\tau}(Y_{i}-\langle X_{i},\hat{A}_{H}\rangle))\langle X_{i},\hat{A}_{H}-A_{0}\rangle\textbf{1}_{H_{i}}$
+$=\frac{1}{n}\sum_{i=1}^{n}(\phi_{\tau}(Y_{i}-\langle X_{i},A_{0}\rangle)-\phi_{\tau}(Y_{i}-\langle X_{i},\hat{A}_{H}\rangle))\langle X_{i},\hat{A}_{H}-A_{0}\rangle$
+$\geq\frac{1}{n}\sum_{i=1}^{n}(\phi_{\tau}(Y_{i}-\langle X_{i},A_{0}\rangle)-\phi_{\tau}(Y_{i}-\langle X_{i},\hat{A}_{H}\rangle))\langle X_{i},\hat{A}_{H}-A_{0}\rangle\textbf{1}_{H_{i}}$
 $=\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}_{H}-A_{0}\rangle^{2}\textbf{1}_{H_{i}}.$
@@
 $\geq\left\langle X_{i},\dfrac{\hat{A}_{H}-A_{0}}{\|\hat{A}_{H}-A_{0}\|_{w(F)}}\right\rangle^{2}\textbf{1}\left(\dfrac{\langle X_{i},\hat{A}_{H}-A_{0}\rangle}{\|\hat{A}_{H}-A_{0}\|_{w(F)}}\leq\dfrac{\tau}{2s}\right)\textbf{1}_{\chi_{i}}$
-$\geq G_{\frac{\tau}{4s}}\left(\left\langle X_{i},\frac{\hat{A}_{H}-A_{0}}{\|\hat{A}_{H}-A_{0}\|_{w(F)}}\right\rangle\right)\textbf{1}_{\chi_{i}}.$
+$\geq G_{\frac{\tau}{4s}}\left(\left\langle X_{i},\frac{\hat{A}_{H}-A_{0}}{\|\hat{A}_{H}-A_{0}\|_{w(F)}}\right\rangle\right)\textbf{1}_{\chi_{i}},$
 where $\chi_{i}=\{|\xi_{i}|\leq\tau/2\}$. Hence, we have
@@
 For a matrix $A\in\mathbb{R}^{m_{1}\times m_{2}}$, we define
-$f(A):=E_{X_{i},\xi_{i}}[G_{\frac{\tau}{4s}}(X_{i},A)\textbf{1}_{\chi_{i}}].$
+$f(A):=\mathbb{E}_{X_{i},\xi_{i}}[G_{\frac{\tau}{4s}}(X_{i},A)\textbf{1}_{\chi_{i}}].$
 We have,
-$f(A)$ $\geq E[\langle X_{i},A\rangle^{2}]\textbf{1}\{|\langle X_{i},A\rangle|\leq\tau/4s\}\textbf{1}_{\chi_{i}}$
-$\geq E\langle X_{i},A\rangle^{2}-E\langle X_{i},A\rangle^{2}\textbf{1}\{|\langle X_{i},A\rangle|>\tau/4s\}-E\langle X_{i},A\rangle^{2}\textbf{1}\{|\xi_{i}|>\tau/2\}$
+$f(A)$ $\geq\mathbb{E}[\langle X_{i},A\rangle^{2}]\textbf{1}\{|\langle X_{i},A\rangle|\leq\tau/4s\}\textbf{1}_{\chi_{i}}$
+$\geq\mathbb{E}\langle X_{i},A\rangle^{2}-\mathbb{E}\langle X_{i},A\rangle^{2}\textbf{1}\{|\langle X_{i},A\rangle|>\tau/4s\}-\mathbb{E}\langle X_{i},A\rangle^{2}\textbf{1}\{|\xi_{i}|>\tau/2\}$
 $\geq\|A\|_{w(F)}^{2}-\dfrac{16s^{2}\|A\|^{2}_{\infty}}{\tau^{2}}\|A\|_{w(F)}^{2}-\dfrac{4\sigma^{2}}{\tau^{2}}\|A\|_{w(F)}^{2}$
@@
 Assume $\|A\|_{w(F)}^{2}=1$. Then there are universal constants $c_{0},c_{1},c_{2}$ such that
-$\left|\dfrac{1}{n}\sum_{i=1}^{n}G_{\frac{\tau}{4s}}(\langle X_{i},A\rangle)-f(A)\right|\leq 8c_{0}\|A\|_{\infty}\|A\|_{1}E\|R\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\frac{\mu\log d}{n}}+8c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}$ (38)
-for all such $A$, uniformly with probability at least $1-\frac{1}{d}$. Here $R=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}$ and $\chi_{i}=\{|\xi_{i}|\leq\tau/2\}$.
+$\left|\dfrac{1}{n}\sum_{i=1}^{n}G_{\frac{\tau}{4s}}(\langle X_{i},A\rangle)\textbf{1}_{\chi_{i}}-f(A)\right|\leq 8c_{0}\|A\|_{\infty}\|A\|_{*}\mathbb{E}\|\mathcal{R}\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\frac{\mu\log d}{n}}+8c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}$ (16)
+for all such $A$, uniformly with probability at least $1-\frac{1}{d}$. Here $\mathcal{R}=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}$ and $\chi_{i}=\{|\xi_{i}|\leq\tau/2\}$.
 By (15) and (16), we get
 $\quad\langle\nabla L(\hat{A}_{H})-\nabla L(A_{0}),\hat{A}_{H}-A_{0}\rangle\geq\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}(1-\dfrac{4\sigma^{2}}{\tau^{2}})-\dfrac{16a^{2}s^{2}}{\tau^{2}}$
-$-\Bigg{(}8c_{0}\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{1}E\|R\|+2\sqrt{2}c_{1}\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{w(F)}\sqrt{\frac{\mu\log d}{n}}$
+$-\Bigg{(}8c_{0}\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{*}\mathbb{E}\|\mathcal{R}\|+2\sqrt{2}c_{1}\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{w(F)}\sqrt{\frac{\mu\log d}{n}}$
 $+8c_{2}\frac{\mu\|\hat{A}_{H}-A_{0}\|_{\infty}^{2}\log d}{n}\Bigg{)}$
 $\geq\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}(1-\dfrac{4\sigma^{2}}{\tau^{2}})-\dfrac{16\|\hat{A}_{H}-A_{0}\|^{2}_{\infty}s^{2}}{\tau^{2}}$
-$-c_{3}\Bigg{(}\|\hat{A}_{H}-A_{0}\|_{1}\|\hat{A}_{H}-A_{0}\|_{\infty}+\frac{\mu\|\hat{A}_{H}-A_{0}\|^{2}_{\infty}\log d}{n}\Bigg{)}-\frac{\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}}{10}$ (39)
+$-c_{3}\Bigg{(}\|\hat{A}_{H}-A_{0}\|_{*}\|\hat{A}_{H}-A_{0}\|_{\infty}+\frac{\mu\|\hat{A}_{H}-A_{0}\|^{2}_{\infty}\log d}{n}\Bigg{)}-\frac{\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}}{10}$ (17)
 where in the last step, we apply the basic inequality ($2ab\leq\frac{a^{2}}{c^{2}}+c^{2}b^{2}$) and choose a suitable constant $c_{3}$. By (17), we get (14).
 Proof of Lemma 10.
-We follow a similar procedure as in the proof of Lemma 18. For fixed $(\alpha,\rho)\in\mathbb{R}^{2}$, define the space:
-$S(\alpha,\rho)=\{A:\|A\|_{w(F)}^{2}=1,\|A\|_{1}\leq\rho,\|A\|_{\infty}\leq\alpha\}.$
+At first, we introduce a classical concentration inequality for empirical processes, which can be found in many textbooks such as Wainwright [2019].
+Lemma 11 (Theorem 3.27 in Wainwright [2019]).
+Let $Z_{i},1\leq i\leq n$, be independent but not necessarily identically distributed random variables taking values in measurable spaces $\mathcal{Z},1\leq i\leq n$.
+$\mathscr{F}$ is a function class.
+Consider the random variable
+$Z=\sup_{f\in\mathscr{F}}\frac{1}{n}\sum_{i=1}^{n}f(Z_{i}).$
+If $\|f\|_{\infty}\leq b$ for all $f\in\mathscr{F}$, we have the following inequality, for $t>0$,
+$P(Z\geq c_{0}E[Z]+c_{1}\tilde{\sigma}\sqrt{t}+c_{2}bt)\leq\exp\left(-nt\right)$ (18)
+where $\tilde{\sigma}^{2}=\sup_{f\in\mathscr{F}}\frac{1}{n}\sum_{i=1}^{n}Ef^{2}(Z_{i})$ and $c_{0},c_{1},c_{2}$ are some universal constants.
+For fixed $(\alpha,\rho)\in\mathbb{R}^{2}$, we define the space:
+$S(\alpha,\rho)=\{A:\|A\|_{w(F)}^{2}=1,\|A\|_{*}\leq\rho,\|A\|_{\infty}\leq\alpha\}.$
 Denote
 $F_{A}(X_{i},\xi_{i})=G_{\frac{\tau}{4s}}(\langle X_{i},A\rangle)\textbf{1}_{\chi_{i}}.$
-for $A\in S(\alpha,\rho)$. Recall $f(A)=E_{X_{i},\xi_{i}}[F_{A}(X_{i},\xi_{i})]$. Since $x^{2}\geq x^{2}\textbf{1}_{|x|\leq 2R}\geq G_{R}(x)$, for $A\in S(\alpha,\rho)$,
-$\quad\leavevmode\nobreak\ E[F^{2}_{A}(X_{i},\xi_{i})]-f^{2}(A)$
-$\leq E[F^{2}_{A}(X_{i},\xi_{i})]$
-$\leq E[G^{2}_{\tau/4s}(\langle X_{i},A\rangle)]$
+for $A\in S(\alpha,\rho)$. Recall $f(A)=\mathbb{E}_{X_{i},\xi_{i}}[F_{A}(X_{i},\xi_{i})]$. Since $x^{2}\geq x^{2}\textbf{1}_{|x|\leq 2R}\geq G_{R}(x)$, for $A\in S(\alpha,\rho)$,
+$\quad\leavevmode\nobreak\ \mathbb{E}[F^{2}_{A}(X_{i},\xi_{i})]-f^{2}(A)$ $\leq\mathbb{E}[F^{2}_{A}(X_{i},\xi_{i})]$
+$\leq\mathbb{E}[G^{2}_{\tau/4s}(\langle X_{i},A\rangle)]$
 $\leq\alpha^{2}\|A\|_{w(F)}^{2}$
@@
 We have
-$E[Z(\alpha,\rho)]$ $\stackrel{{\scriptstyle\text{i}}}{{\leq}}E\sup_{A\in\Omega(\alpha,\rho)}\left|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}F_{A}(X_{i},\xi_{i})\right|$
-$\stackrel{{\scriptstyle\text{ii}}}{{\leq}}2\alpha E\sup_{A\in\Omega(\alpha,\rho)}\left|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\langle X_{i}\textbf{1}_{\chi_{i}},A\rangle\right|$
-$\stackrel{{\scriptstyle\text{iii}}}{{\leq}}4\alpha\rho E\left\|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}\right\|.$
+$\mathbb{E}[Z(\alpha,\rho)]$ $\stackrel{{\scriptstyle\text{i}}}{{\leq}}\mathbb{E}\sup_{A\in\Omega(\alpha,\rho)}\left|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}F_{A}(X_{i},\xi_{i})\right|$
+$\stackrel{{\scriptstyle\text{ii}}}{{\leq}}2\alpha\mathbb{E}\sup_{A\in\Omega(\alpha,\rho)}\left|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\langle X_{i}\textbf{1}_{\chi_{i}},A\rangle\right|$
+$\stackrel{{\scriptstyle\text{iii}}}{{\leq}}4\alpha\rho\mathbb{E}\left\|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}\right\|.$
 In step (i) we use symmetrization technique.
@@
 Then, in step (ii) we apply Talagrand-Ledoux contraction inequality.
-Step (iii) is due to $|\langle A,B\rangle|\leq\|A\|\|B\|_{1}$.
+Step (iii) is due to $|\langle A,B\rangle|\leq\|A\|\|B\|_{*}$.
 We take $t=\frac{2\mu\log d}{n}$ in (18). Since $\tilde{\sigma}\leq\alpha$ and $\|F_{A}(X_{i},\xi_{i})\|_{\infty}\leq\alpha^{2}$, with probability at least $1-\exp(-2\mu\log d)$,
-$Z(\alpha,\rho)\leq 4c_{0}\alpha\rho E\|R\|+\sqrt{2}c_{1}\alpha\sqrt{\frac{\mu\log d}{n}}+4c_{2}\frac{\mu\alpha^{2}\log d}{n}.$ (40)
+$Z(\alpha,\rho)\leq 4c_{0}\alpha\rho\mathbb{E}\|\mathcal{R}\|+\sqrt{2}c_{1}\alpha\sqrt{\frac{\mu\log d}{n}}+4c_{2}\frac{\mu\alpha^{2}\log d}{n}.$ (19)
 Note that under Assumption 3,
@@
 Then, we have
-$\|A\|_{1}\leq\sqrt{m}\|A\|_{F}\leq\sqrt{\mu d^{3}}$
+$\|A\|_{*}\leq\sqrt{m}\|A\|_{F}\leq\sqrt{\mu d^{3}}$
 and
-$1\leq\|A\|_{F}\leq\|A\|_{1}.$
-Without loss of generality, we can assume $\|A\|_{1}\in[1,\sqrt{\mu d^{3}}]$. Similarly, we have
+$1\leq\|A\|_{F}\leq\|A\|_{*}.$
+Without loss of generality, we can assume $\|A\|_{*}\in[1,\sqrt{\mu d^{3}}]$. Similarly, we have
 $\|A\|_{\infty}\leq\|A\|_{F}\leq\sqrt{\mu m_{1}m_{2}}$
@@
 Consider
-$\tilde{S}(k,l)=\{A;\|A\|_{w(F)}=1,2^{k-1}\leq\|A\|_{\infty}\leq 2^{k},\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ 2^{l-1}\leq\|A\|_{1}\leq 2^{l}\}$
+$\tilde{S}(k,l)=\{A;\|A\|_{w(F)}=1,2^{k-1}\leq\|A\|_{\infty}\leq 2^{k},\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ 2^{l-1}\leq\|A\|_{*}\leq 2^{l}\}$
 Every matrix $A$ with $\|A\|_{w(F)}=1$ belongs to some $\tilde{S}(k,l)$ with $k=\lfloor\log_{2}\frac{1}{d}\rfloor,\ldots,\lceil\frac{1}{2}\log_{2}(\mu d^{2})\rceil$ and $l=1,\ldots,\lceil\log_{2}\frac{1}{2}(\mu d^{3})\rceil$.
-Thus the total number of the pairs $(k,l)$ is less than $H^{2}=9\log_{2}^{2}(\mu d)$.
+Thus the total number of the pairs $(k,l)$ is less than $\Lambda^{2}=9\log_{2}^{2}(\mu d)$.
 Denote the events $\mathcal{E}$ that a matrix $A$ violates the following relation
-$\left|\dfrac{1}{n}\sum_{i=1}^{n}F_{A}(X_{i},\xi_{i})-f(A)\right|\leq 16c_{0}\|A\|_{\infty}\|A\|_{1}E\|R\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\frac{\mu\log d}{n}}+16c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}$
+$\left|\dfrac{1}{n}\sum_{i=1}^{n}F_{A}(X_{i},\xi_{i})-f(A)\right|\leq 16c_{0}\|A\|_{\infty}\|A\|_{*}\mathbb{E}\|\mathcal{R}\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\frac{\mu\log d}{n}}+16c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}$ (20)
 and denote $\mathcal{E}_{k,l}$ the event that there exists a matrix in $\tilde{S}(k,l)$ violating the following relation
-$\left|\dfrac{1}{n}\sum_{i=1}^{n}F_{A}(X_{i},\xi_{i})-f(A)\right|\leq 4c_{0}2^{k}2^{l}E\|R\|+\sqrt{2}c_{1}2^{k}\sqrt{\frac{\mu\log d}{n}}+4c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}.$
+$\left|\dfrac{1}{n}\sum_{i=1}^{n}F_{A}(X_{i},\xi_{i})-f(A)\right|\leq 4c_{0}2^{k}2^{l}\mathbb{E}\|\mathcal{R}\|+\sqrt{2}c_{1}2^{k}\sqrt{\frac{\mu\log d}{n}}+4c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}.$
 $\mathcal{E}$ is contained in the union $\cup_{k,l}\mathcal{E}_{k,l}$. Indeed, letting $\alpha=2^{k}$ and $\rho=2^{l}$, for a matrix $A$ in $\tilde{S}(k,l)$, we have
-$\left|\dfrac{1}{n}\sum_{i=1}^{n}F_{A}(X_{i},\xi_{i})-f(A)\right|$ $\geq 16c_{0}\|A\|_{\infty}\|A\|_{1}E\|R\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\frac{\mu\log d}{n}}+16c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}$
-$\geq 16c_{0}2^{k-1}2^{l-1}E\|R\|+2\sqrt{2}c_{1}2^{k-1}\sqrt{\frac{\mu\log d}{n}}+16c_{2}2^{2(k-1)}\frac{\mu\log d}{n}$
+$\left|\dfrac{1}{n}\sum_{i=1}^{n}F_{A}(X_{i},\xi_{i})-f(A)\right|$ $\geq 16c_{0}\|A\|_{\infty}\|A\|_{*}\mathbb{E}\|\mathcal{R}\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\frac{\mu\log d}{n}}+16c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}$
+$\geq 16c_{0}2^{k-1}2^{l-1}\mathbb{E}\|\mathcal{R}\|+2\sqrt{2}c_{1}2^{k-1}\sqrt{\frac{\mu\log d}{n}}+16c_{2}2^{2(k-1)}\frac{\mu\log d}{n}$
 $=4c_{0}\alpha\rho+\sqrt{2}\alpha\sqrt{\frac{\mu\log d}{n}}+4c_{2}\alpha^{2}\frac{\mu\log d}{n}.$
 Thus, by (19),
-$P(\mathcal{E})\leq P(\cup_{k,l}\mathcal{E}_{k,l})\leq H^{2}\exp(-2\mu\log d)\leq\frac{9\log_{2}^{2}\mu d}{d^{\mu}}\cdot\frac{1}{d^{\mu}}\leq\frac{1}{d}$
+$\mathbb{P}(\mathcal{E})\leq\mathbb{P}(\cup_{k,l}\mathcal{E}_{k,l})\leq\Lambda^{2}\exp(-2\mu\log d)\leq\frac{9\log_{2}^{2}\mu d}{d^{\mu}}\cdot\frac{1}{d^{\mu}}\leq\frac{1}{d}$
 for sufficiently large $d$. ∎
-4.8 Other technical lemmas
+4.3 Proof of Theorem 3
+From Appendix A in Klopp [2014], we know the following two inequalities. Given $\lambda\geq 3\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\|$, we have
+$\frac{1}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}-A_{0}\rangle^{2}\leq\frac{5}{3}\lambda\sqrt{r}\|\hat{A}-A_{0}\|_{F}.$ (21)
+We also have
+$\|\hat{A}-A_{0}\|_{*}\leq\sqrt{72r}\|\hat{A}-A_{0}\|_{F}.$ (22)
+Regarding the difference between $\frac{1}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}-A_{0}\rangle^{2}$ and $\|\hat{A}-A_{0}\|^{2}_{w(F)}$, we have the following lemma.
 Lemma 12.
-Recall the definitions of $P_{A_{0}}(\cdot)$ and $P_{A_{0}}^{\perp}(\cdot)$ defined in Section 1. Consider the following estimator:
-$\bar{A}=\operatorname*{arg\,min}_{A\in\Omega}\left\{L(A)+\lambda\|A\|_{1}\right\},$
-where $\Omega$ is the domain of $A$ and $L(A)$ is a convex loss function. If we take $\lambda\geq 3\|\nabla L(A_{0})\|$, then $\bar{A}-A_{0}$ satisfies
-$\|P_{A_{0}}^{\perp}(\bar{A}-A_{0})\|_{1}\leq 2\|P_{A_{0}}(\bar{A}-A_{0})\|_{1}.$
-As special cases, we have the following inequalities. Take $\lambda\geq 3\|\frac{2}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\|$ in (2), then
-$\|P_{A_{0}}^{\perp}(\hat{A}-A_{0})\|_{1}\leq 2\|P_{A_{0}}(\hat{A}-A_{0})\|_{1}.$
-Take $\lambda\geq 3\frac{\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\|}{\sqrt{\frac{1}{n}\sum_{i=1}^{n}\xi_{i}^{2}}}$ in (8) , then
-$\|P_{A_{0}}^{\perp}(\hat{A}_{S}-A_{0})\|_{1}\leq 2\|P_{A_{0}}(\hat{A}_{S}-A_{0})\|_{1}.$
-Take $\lambda\geq 3\|\frac{1}{n}\sum_{i=1}^{n}\psi_{\tau}(\xi_{i})X_{i}\|$ in (10) , then
-$\|P_{A_{0}}^{\perp}(\hat{A}_{H}-A_{0})\|_{1}\leq 2\|P_{A_{0}}(\hat{A}_{H}-A_{0})\|_{1}.$
+Assume $\|A\|_{w(F)}^{2}=1$. Then there is a constant $C$ dependent on $c_{0},c_{1},c_{2}$ such that
+$\left|\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},A\rangle^{2}-\|A\|^{2}_{w(F)}\right|\leq C\left(\|A\|_{\infty}\|A\|_{*}\mathbb{E}\|\mathcal{R}\|+\|A\|_{\infty}\sqrt{\frac{\mu\log d}{n}}+\frac{\mu\|A\|_{\infty}^{2}\log d}{n}\right)$ (23)
+for all such $A$, uniformly with probability at least $1-\frac{1}{d}$. Here $\mathcal{R}=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}$.
 Proof.
-As a solution of a convex program, by Proposition 1.3 in [Bubeck, 2014] , $\bar{A}$ satisfies:
-$\langle\nabla L(\bar{A})+\lambda Z,\bar{A}-A_{0}\rangle\leq 0.$
-where $Z\in\partial\|\bar{A}\|_{1}$. Since $\|\cdot\|_{1}$ is convex,
-$\langle Z,A_{0}-\bar{A}\rangle\leq\|A_{0}\|_{1}-\|\bar{A}\|_{1}.$
-By the convexity of $L(\cdot)$,
-$\langle\nabla L(\bar{A})-\nabla L(A_{0}),\bar{A}-A_{0}\rangle\geq 0.$
-Combining the results above, we obtain:
-$0\leq\langle\nabla L(\bar{A})-\nabla L(A_{0}),\bar{A}-A_{0}\rangle\leq\lambda(\|A_{0}\|_{1}-\|\bar{A}\|_{1})+\|\nabla L(A_{0})\|\|\bar{A}-A_{0}\|_{1}.$ (41)
+Denote
+$Z_{\tau}(A)=\left|\dfrac{1}{n}\sum_{i=1}^{n}G_{\frac{\tau}{4s}}(\langle X_{i},A\rangle)\textbf{1}_{\chi_{i}}-f(A)\right|,$
+$Z=\left|\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},A\rangle^{2}-\|A\|^{2}_{w(F)}\right|,$
+$\kappa(A)=8c_{0}\|A\|_{\infty}\|A\|_{*}\mathbb{E}\|\mathcal{R}\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\frac{\mu\log d}{n}}+8c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}.$
+Then, we consider the variables
+$Y_{\tau}=\sup_{\|A\|_{w(F)}=1}\frac{Z_{\tau}(A)}{\kappa(A)},$
+$Y=\sup_{\|A\|_{w(F)}=1}\frac{Z}{\kappa(A)}.$
+Given $\|A\|_{w(F)}=1$, we have that $\kappa(A)$ is uniformly lower bounded away from zero (see the proof of Lemma 10). And we note
+$\left|\dfrac{1}{n}\sum_{i=1}^{n}G_{\frac{\tau}{4s}}(\langle X_{i},A\rangle)\textbf{1}_{\chi_{i}}-\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},A\rangle^{2}\right|$ $\leq C_{1}sa^{2}\frac{1}{n}\sum_{i=1}^{n}\textbf{1}\left(|\xi_{i}|>\frac{\tau}{2}\right),$
+$\left|f(A)-\|A\|^{2}_{w(F)}\right|$ $\leq C_{2}sa^{2}/\tau,$
+where $C_{1},C_{2}$ are constants that are not dependent on $\tau$.
+Thus, we have that $Y_{\tau}$ converges to $Y$ in probability and thus in distribution.
+There exist constants y (dense in $\mathbb{R}$) such that
+$\lim_{\tau\rightarrow\infty}\mathbb{P}(Y_{\tau}\leq y)=\mathbb{P}(Y\leq y).$
+For $y\geq 1$, by Lemma 10, it holds that
+$\mathbb{P}(Y_{\tau}\leq y)\geq 1-\frac{1}{d}$
+for all $\tau$. There exists some $y\geq 1$ such that
+$P(Y\leq y)\geq 1-\frac{1}{d},$
+which implies Lemma 12. ∎
+Combining Lemma 12 and (21), we obtain
+$\|\hat{A}-A_{0}\|^{2}_{w(F)}$ $\leq 8c_{0}\|\hat{A}-A_{0}\|_{\infty}\|\hat{A}-A_{0}\|_{*}\mathbb{E}\|\mathcal{R}\|+2\sqrt{2}c_{1}\|\hat{A}-A_{0}\|_{w(F)}\|\hat{A}-A_{0}\|_{\infty}\sqrt{\frac{\mu\log d}{n}}$
+$+8c_{2}\frac{\mu\|\hat{A}-A_{0}\|_{\infty}\log d}{n}+\frac{5}{3}\lambda\sqrt{r}\|\hat{A}-A_{0}\|_{F}.$
+By $\|\hat{A}-A_{0}\|_{\infty}\leq 2a$ and (22), we obtain the following inequalities:
+$\lambda\sqrt{r}\|\hat{A}-A_{0}\|_{F}\leq 5u\lambda^{2}rm_{1}m_{2}+\frac{1}{5}\frac{\|\hat{A}-A_{0}\|^{2}_{F}}{\mu}\leq 5u\lambda^{2}rm_{1}m_{2}+\frac{1}{5}\|\hat{A}-A_{0}\|^{2}_{w(F)},$
+$\frac{\mu\|\hat{A}-A_{0}\|_{\infty}\log d}{n}\leq\frac{2\mu a\log d}{n},$
+$\|\hat{A}-A_{0}\|_{\infty}\|\hat{A}-A_{0}\|_{*}\mathbb{E}\|\mathcal{R}\|\leq 5\|\hat{A}-A_{0}\|_{\infty}\sqrt{r}\|\hat{A}-A_{0}\|_{F}\mathbb{E}\|\mathcal{R}\|\leq 5\times 4\mu a^{2}rm_{1}m_{2}\mathbb{E}^{2}\|\mathcal{R}\|+\frac{1}{5}\|\hat{A}-A_{0}\|^{2}_{w(F)},$
+$\|\hat{A}-A_{0}\|_{w(F)}\|\hat{A}-A_{0}\|_{\infty}\sqrt{\frac{\mu\log d}{n}}\leq\frac{1}{5}\|\hat{A}-A_{0}\|^{2}_{w(F)}+5\times 4\frac{\mu a^{2}\log d}{n}.$
+Putting together the pieces, we get
+$\|\hat{A}-A_{0}\|^{2}_{w(F)}\lesssim\mu m_{1}m_{2}\left(r\lambda^{2}+a^{2}\mathbb{E}^{2}\|\Sigma_{R}\|\right)+\frac{\mu a^{2}\log d}{n}.$ (24)
+By (29), we obtain
+$\left\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\lesssim\sigma\sqrt{\frac{1}{nm}},$
+given $n\geq m\log^{4}d\cdot(\log d+\log n)$. Then, we can take $\lambda\sim\sigma\sqrt{\frac{1}{nm}}$. By (31),
+$\mathbb{E}\left\|\sum\epsilon_{i}X_{i}\right\|\lesssim\sqrt{\frac{1}{nm}}.$
+Plugging the above two estimates into (24), we obtain
+$\frac{\|\hat{A}-A_{0}\|^{2}_{F}}{m_{1}m_{2}}\lesssim\frac{\mu^{2}\max(a^{2},\sigma^{2})rM}{n}+\frac{\mu^{2}a^{2}\log d}{n},$
+which leads to the statement of Theorem 3.
+4.4 Proof of Theorem 5
+In this section, we denote $\xi=(\xi_{1}/\sqrt{n},\ldots,\xi_{n}/\sqrt{n})$. From Appendix D in Klopp [2014], given
+$\lambda\geq\frac{3\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\|}{\|\xi\|_{2}},$
+we have that
+$\|\hat{A}_{S}-A_{0}\|_{*}\leq\sqrt{9r}\|\hat{A}_{S}-A_{0}\|_{F}.$ (25)
+Again, from Appendix D in Klopp [2014], we have
+$\frac{1}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}_{S}-A_{0}\rangle^{2}\leq\frac{14}{3}\lambda Q(A_{0})\sqrt{2r}\|\hat{A}_{S}-A_{0}\|_{F}+\frac{1}{4}\|\hat{A}_{S}-A_{0}\|^{2}_{w(F)},$
+under condition $44\mu m_{1}m_{2}\lambda^{2}r\leq 1/$ 4. By Lemma 12, we get
+$\|\hat{A}_{S}-A_{0}\|_{w(F)}^{2}$ $\leq 8c_{0}\|\hat{A}_{S}-A_{0}\|_{\infty}\|\hat{A}_{S}-A_{0}\|_{*}\mathbb{E}\|\mathcal{R}\|+2\sqrt{2}c_{1}\|\hat{A}_{S}-A_{0}\|_{w(F)}\|\hat{A}_{S}-A_{0}\|_{\infty}\sqrt{\frac{\mu\log d}{n}}$
+$+8c_{2}\frac{\mu\|\hat{A}_{S}-A_{0}\|_{\infty}\log d}{n}+\frac{14}{3}\lambda\|\xi\|_{2}\sqrt{2r}\|\hat{A}_{S}-A_{0}\|_{F}+\frac{1}{4}\|\hat{A}_{S}-A_{0}\|^{2}_{w(F)}.$
+Applying $\|\hat{A}_{S}-A_{0}\|_{\infty}\leq 2a$ and (25), repeat the same precesses as in Proof of Theorem 3 and we obtain
+$\|\hat{A}_{S}-A_{0}\|_{w(F)}^{2}\lesssim\mu m_{1}m_{2}r\left(\|\xi\|^{2}_{2}\lambda^{2}+a^{2}\mathbb{E}^{2}\|\mathcal{R}\|\right)+\frac{\mu\log d}{n},$ (26)
+where $\mathcal{R}$ is defined as in Lemma 12. By Theorem 3.1.1 in Vershynin [2018], denoting $\xi=(\xi_{1}/n,\ldots,\xi_{n}/n)$, we get
+$\mathbb{P}\left(\left|\left\|\xi\right\|_{2}-\sigma\right|\geq L_{1}\sigma t\right)\leq 2\exp(-Cnt^{2}),$
+under Assumption 2. With $t=0.1/L_{1}$, we obtain that with probability at least $1-\exp(-Cn)$ ($C$ is dependent on $L_{1}$), the following holds
+$0.9\sigma\leq\|\xi\|_{2}\leq 1.1\sigma.$
+By (29),
+$\left\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\lesssim\sigma\sqrt{\frac{1}{nm}},$
+given $n\geq m\log^{4}d\cdot(\log d+\log n)$. Thus, we can take $\lambda\sim\sqrt{\frac{1}{nm}}$. By (31),
+$\mathbb{E}\left\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|\lesssim\sqrt{\frac{1}{nm}}.$
+Plugging $\lambda$ and the upper bounds for $\mathbb{E}\left\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|$ into (26). we obtain the convergence rate
+$\frac{\|\hat{A}_{S}-A_{0}\|_{2}^{2}}{m_{1}m_{2}}\lesssim\frac{\mu^{2}\max(a^{2},\sigma^{2})rM}{n}+\frac{\mu^{2}a^{2}\log d}{n},$
+where the constants are dependent on $L_{1}\sim L_{3}$. This leads to the statement of Theorem 5 as the second term above on RHS is less that the first term.
+4.5 Proof of sharp bounds for spectral norms
+In this section, we present advanced inequalities introduced in Brailovskaya and van Handel [2024] and how to use them.
+Let $Q=\sum_{i=1}^{n}Q_{i}\in\mathbb{R}^{m_{1}\times m_{2}}$ be a random matrix where $Q_{i}$ are independent random matrices with mean zero.
+We define
+$\gamma(Q)$ $=\left(\max(\|\mathbb{E}[QQ^{T}]\|,\|\mathbb{E}[Q^{T}Q]\|)\right)^{1/2},$
+$\gamma_{*}(Q)$ $=\sup_{\|y\|_{2}=\|z\|_{2}=1}\left(\mathbb{E}(y^{T}Qz)^{2}\right)^{1/2},$
+$g(Q)$ $=\|\operatorname{Cov}(Q)\|^{1/2},$
+$R(Q)$ $=\left\|\max_{1\leq i\leq n}\|Q_{i}\|\right\|_{\infty},$
+where $\operatorname{Cov}(Q)$ is the covariance matrix of all the entries of $Q$ and its dimension is $m_{1}m_{2}\times m_{1}m_{2}$.
+Theorem 13 (Corollary 2.17 in Brailovskaya and van Handel [2024]).
+With the above notations, there exist a universal constant $C$ such that
+$\mathbb{P}(\|Q\|\geq 2\gamma(Q)+C(g(Q)^{1/2}\gamma(Q)^{1/2}\log^{3/4}d+\gamma_{*}(Q)t^{1/2}+R(Q)^{1/3}\gamma(Q)^{2/3}t^{2/3}+R(Q)t))\leq de^{-t}$ (27)
+and a universal constant $C$ such that
+$E\|Q\|\leq 2\gamma(Q)+C(g(Q)^{1/2}\gamma(Q)^{1/2}\log^{3/4}d+R(Q)^{1/3}\gamma(Q)^{2/3}\log^{2/3}d+R(Q)\log d).$ (28)
+Remark.
+The matrix inequalities in Brailovskaya and van Handel [2024] are formulated for $Q_{i}\in\mathbb{R}^{d\times d}$ (not necessarily self-adjoint), and can be extended to rectangular matrices in the above form.
+See Remark 2.1 in Brailovskaya and van Handel [2024].
+Theorem 13 requires $Q_{i}$ to have bounded spectral norms, which is too restrictive in the context of matrix completion.
+To leverage Theorem 13, we need to truncate the noise variables.
+Brailovskaya and van Handel [2024] also provides concentration inequalities for the unbounded case (Theorem 2.8) that is derived by truncation techniques.
+However, applying Theorem 2.8 in Brailovskaya and van Handel [2024] could result in worse sample size conditions as Brailovskaya and van Handel [2024] considers general cases.
+For matrix completion, we need derive specially tailored concentration inequalities from Theorem 13 (essentially from Theorem 2.6 in Brailovskaya and van Handel [2024]).
+Lemma 14.
+(1) Assume the sampling matrices $X_{i}$ satisfy Assumption 3 and $\xi_{i}$ are $\psi_{\alpha}$ random variables with $\mathbb{E}[\xi_{i}]=0$, $\mathbb{E}[\xi_{i}^{2}|X_{i}]\leq\sigma^{2}$ and uniformly bounded $\psi_{\alpha}$ norms
+$\left\|[\xi_{i}|X_{i}]\right\|_{\psi_{\alpha}}\leq L_{1}\sigma.$
+The following inequality holds with probability at least $1-\frac{1}{d}$,
+$\left\|\dfrac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\leq C\sigma\sqrt{\dfrac{1}{nm}},$ (29)
+given $n\geq m\log^{4}d\cdot(\log^{2/\alpha}n+\log^{2/\alpha}d)$.
+In particular, for sub-Gaussian noise, (29) holds given $n\geq m\log^{4}d\cdot(\log n+\log d)$.
+Here $C$ is a constant dependent on $L_{1}\sim L_{3}$.
+(2)Assume $\xi_{i}$ satisfy Assumption 1 and are symmetric and $X_{i}$ satisfy Assumption 3.
+Taking $\tau=\frac{\max(\sigma,a)}{\log^{2}d}\sqrt{\frac{n}{m}}$, the following inequality holds with probability at least $1-\frac{1}{d}$,
+$\left\|\dfrac{1}{n}\sum_{i=1}^{n}\phi_{\tau}(\xi_{i})X_{i}\right\|\leq C\max(\sigma,a)\sqrt{\dfrac{1}{nm}}.$ (30)
+where $C$ is dependent on $L_{1}$ and $L_{3}$.
+(3) If $n\geq m\log^{4}d$, we have the following result:
+$\mathbb{E}\left\|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|\leq C\sqrt{\frac{1}{nm}},$ (31)
+where $C$ is dependent on $L_{2}$ and $L_{3}$.
+(4) If $n\geq m\log^{4}d$, we have the following inequality:
+$\mathbb{E}\left\|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}\right\|\leq C\sqrt{\frac{1}{nm}},$ (32)
+where $\xi_{i}=\{|\xi_{i}|\leq\tau/2\}$.
+Proof of Lemma 14.
+Since Theorem 13 requires $\|Q_{i}\|$ to be bounded, we need to truncate $\xi_{i}$ to $\xi_{i}\textbf{1}(|\xi_{i}|\leq\tau)$ for some $\tau\geq 0$.
+Define $\bar{\xi}_{i}=\xi_{i}\textbf{1}(|\xi_{i}|\leq\tau)-\mathbb{E}[\xi_{i}\textbf{1}(|\xi_{i}|\leq\tau)|X_{i}]$.
+Note $\mathbb{E}[\bar{\xi}_{i}]=0$ and $\mathbb{E}[\bar{\xi}_{i}^{2}]\leq\sigma^{2}$.
+We now compute the four parameters for $Q=\frac{1}{n}\sum_{i=1}^{n}\bar{\xi}_{i}X_{i}$.
+Define the event
+$\mathcal{E}_{i}(j,k)=\{X_{i}=e_{j}(m_{1})e_{k}(m_{2})^{T}\}.$
+1.
+$\gamma(Q)\leq\sqrt{\frac{L_{2}\sigma^{2}}{nm}}$.
+Note $\mathbb{E}[\bar{\xi}^{2}_{i}X_{i}X_{i}^{T}]$ is a diagonal matrix.
+The elements on the diagonal line are $\sum_{k=1}^{m_{2}}P_{jk}\mathbb{E}[\bar{\xi}_{i}^{2}|\mathcal{E}_{i}(j,k)],j=1,2,\ldots,m_{1}$, which are less than $\frac{L_{2}\sigma^{2}}{m}$.
+The same bound holds for $\mathbb{E}[\bar{\xi}_{i}^{2}X_{i}^{T}X_{i}]$.
+Since
+$\mathbb{E}[QQ^{T}]=\frac{1}{n^{2}}\sum_{i=1}^{n}\mathbb{E}[\bar{\xi}_{i}^{2}X_{i}X_{i}^{T}],$
+we obtain $\|E[QQ^{T}]\|\leq\frac{L_{2}\sigma^{2}}{nm}$. Similarly, $\|\mathbb{E}[Q^{T}Q]\|\leq\frac{L_{2}\sigma^{2}}{nm}$. Thus $\gamma(Q)\leq\sqrt{\frac{L_{2}\sigma^{2}}{nm}}$.
+2. $\gamma_{*}(Q)\leq\sqrt{\frac{L_{3}\sigma^{2}}{nm\log^{3}d}}$. For $y\in\mathbb{R}^{m_{1}},z\in\mathbb{R}^{m_{2}}$ with $\|y\|_{2}=\|z\|_{2}=1$, we have the following result:
+$\mathbb{E}(y^{T}\bar{\xi}_{i}X_{i}z)^{2}$ $=\sum_{j,k}P_{jk}\mathbb{E}[\bar{\xi}_{i}^{2}|\mathcal{E}_{i}(j,k)]y^{2}_{j}z^{2}_{k}$
+$\leq\sigma^{2}\sum_{j,k}P_{jk}y^{2}_{j}z^{2}_{k}$
+$\leq\dfrac{L_{3}\sigma^{2}}{m\log^{3}d}.$
+where we use $\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}y^{2}_{j}z^{2}_{k}=\|y\|_{2}^{2}\cdot\|z\|_{2}^{2}=1$ and $\max P_{jk}\leq\frac{L_{3}}{m\log^{3}d}$.
+Since $\bar{\xi}_{i}X_{i}$ are independent and mean-zero, $\mathbb{E}(y^{T}Qz)^{2}=\dfrac{1}{n^{2}}\sum_{i=1}^{n}\mathbb{E}\left(y^{T}\bar{\xi}_{i}X_{i}z\right)^{2}\leq\frac{L_{1}L_{3}\sigma^{2}}{nm\log^{3}d}$.
+Thus $\gamma_{*}(Q)\leq\sqrt{\frac{L_{3}\sigma^{2}}{nm\log^{3}d}}$.
+3.
+$g(Q)\leq\sqrt{\frac{L_{3}\sigma^{2}}{nm\log^{3}d}}$.
+Since $\bar{\xi}_{i}X_{i}$’s are independent, $\operatorname{Cov}(Q)=\dfrac{1}{n^{2}}\sum_{i=1}^{n}\operatorname{Cov}(\bar{\xi}_{i}X_{i})$.
+Define $\textbf{1}_{(j,k)}=\textbf{1}(X_{i}=e_{j}(m_{1})e_{k}(m_{2})^{T})$ (for simplicity of notation, we do not include a subscript i) and denote $I(j,k)=\bar{\xi}_{i}\textbf{1}_{(j,k)}$.
+Note the following facts
+$\mathbb{E}[I(j,k)]=P_{jk}\mathbb{E}[\bar{\xi}_{i}|\mathcal{E}_{i}(j,k)]=0,$
+$I(j,k)I(l,h)=0\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ (j,k)\neq(l,h).$
+Thus $\text{Cov}\big{(}I(j,k),I(l,h)\big{)}=\mathbb{E}[I(j,k)I(l,h)]-\mathbb{E}[I(j,k)]E[I(l,h)]=0$ for $(j,k)\neq(l,h)$. Then we obtain
+$\text{Cov}(\bar{\xi}_{i}X_{i})\Big{(}(j,k),(l,h)\Big{)}=\begin{cases}\mathbb{E}[\bar{\xi}^{2}_{i}|\mathcal{E}_{i}(j,k)]P_{jk}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ \leavevmode\nobreak\ (j,k)=(l,h),\\ 0\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ (j,k)\neq(l,h).\end{cases}$
+Since $P_{jk}\leq\frac{L_{3}}{m\log^{3}d}$, $\|\operatorname{Cov}(\bar{\xi}_{i}X_{1})\|\leq\frac{L_{3}\sigma^{2}}{m\log^{3}d}$. Thus, $g(Q)\leq\sqrt{\frac{L_{3}\sigma^{2}}{nm\log^{3}d}}$ holds.
+4.$R(Q)\leq\frac{2\tau}{n}$. Obviously, $\|\bar{\xi}_{i}X_{i}\|=|\bar{\xi}_{i}|$, and $\max_{1\leq i\leq n}|\bar{\xi}_{i}|\leq 2\tau$.
+Now, we can plug the four estimates in (27). Taking $t=3\log d$, we get that with probability at least $1-\frac{1}{d^{2}}$, the following holds
+$\|Q\|$ $\leq 2\sqrt{\frac{\sigma^{2}}{nm}}$
+$+C\Bigg{(}\left(\frac{\sigma^{2}}{nm}\right)^{1/2}\cdot\left(\dfrac{\log^{3/4}d}{\log^{3/4}d}\right)+\left(\frac{\sigma^{2}}{nm}\right)^{1/2}\cdot\left(\dfrac{\log^{1/2}d}{\log^{3/2}d}\right)$
+$+\left(\frac{\log^{2}d\cdot\tau}{n}\right)^{1/3}\cdot\left(\frac{\sigma^{2}}{nm}\right)^{1/3}+\dfrac{\log d\cdot\tau}{n}\Bigg{)},$ (33)
+where $C$ is dependent on $L_{2}\sim L_{3}$.
+We can see the quantities in the first line and the second line of RHS of (33) are less than $C\sqrt{\frac{\sigma^{2}}{nm}}$.
+When $\xi_{i}$ are $\psi_{\alpha}$ random variables, we take $\tau=C_{\alpha}\sigma(\log^{1/\alpha}n+\log^{1/\alpha}d)$.
+By (35), with $C_{\alpha}$ properly chosen,
+$\mathbb{P}\left(\max_{1\leq i\leq n}|\xi_{i}|\geq C_{\alpha}\sigma\left(\log^{1/\alpha}n+\log^{1/\alpha}d\right)\right)\leq\frac{1}{d^{2}}.$
+To make the quantities in the last line of (33) be less than $C\sqrt{\frac{\sigma^{2}}{nm}}$, we need
+$\left(\frac{\sigma\log^{2}d\cdot(\log^{1/\alpha}n+\log^{1/\alpha}d)}{n}\right)^{1/3}\cdot\left(\frac{\sigma^{2}}{nm}\right)^{1/3}\leq\sqrt{\dfrac{\sigma^{2}}{nm}},$
+$\dfrac{\sigma\log d\cdot(\log^{1/\alpha}n+\log^{1/\alpha}d)}{n}\leq\sqrt{\dfrac{\sigma^{2}}{nm}}.$
+The above conditions imply
+$n$ $\geq m\log^{4}d\cdot(\log^{2/\alpha}n+\log^{2/\alpha}d),$
+$n$ $\geq m\log^{2}d\cdot(\log^{2/\alpha}n+\log^{2/\alpha}d).$
+On $\{\max_{1\leq i\leq n}|\xi_{i}|\leq\tau\}$,
+$\left\|\dfrac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|$ $=\left\|Q+\dfrac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\xi_{i}\textbf{1}(|\xi_{i}|\leq\tau)]X_{i}\right\|$
+$\leq\|Q\|+\left\|\dfrac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\xi_{i}\textbf{1}(|\xi_{i}|>\tau)]X_{i}\right\|$
+$\leq\|Q\|+\frac{1}{n}\sum_{i=1}^{n}|\mathbb{E}[\xi_{i}\textbf{1}(|\xi_{i}|>\tau)|X_{i}]|,$
+where we use $\mathbb{E}[\xi_{i}\textbf{1}(\xi_{i}\leq\tau)]=\mathbb{E}[\xi_{i}\textbf{1}(\xi_{i}>\tau)]$ since $\mathbb{E}[\xi_{i}]=0$.
+For $\tau=C_{\alpha}\sigma(\log^{1/\alpha}n+\log^{1/\alpha}d)$ and a random variable $Y$ with $\|Y\|_{\psi_{\alpha}}\leq L_{1}\sigma$ , we can find suitable large constant $C_{\alpha}$ dependent on $L_{1}$ such that the following holds
+$\mathbb{E}[Y\textbf{1}(|Y|>\tau)]$ $\leq\mathbb{E}[|Y|\textbf{1}(|Y|>\tau)]$
+$\leq\mathbb{E}^{1/2}[Y^{2}]P^{1/2}(|Y|>\tau)$
+$\leq\sigma\exp\left(-2\max(\log^{1/\alpha}n,\log^{1/\alpha}d)^{\alpha}\right)$
+$\leq\frac{\sigma}{n}\leq\sigma\sqrt{\frac{1}{nm}},$
+given $n\geq m$. Combining the results, we obtain
+$\mathbb{P}\left(\left\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\geq C_{\alpha}\sigma\sqrt{\frac{1}{nm}}\right)$ $\leq\mathbb{P}\left(\left\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\geq C_{\alpha}\sigma\sqrt{\frac{1}{nm}},\max_{1\leq i\leq n}|\xi_{i}|\leq\tau\right)+\mathbb{P}\left(\max_{1\leq i\leq n}|\xi_{i}|>\tau\right)$
+$\leq\frac{1}{d},$
+given $n\geq m\log^{4}d\cdot(\log^{2/\alpha}n+\log^{2/\alpha}d)$. In particular, for sub-Gaussian noise, we have that
+$\mathbb{P}\left(\left\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\leq C\sigma\sqrt{\frac{1}{nm}}\right)\geq 1-\frac{1}{d}$
+given $n\geq\log^{4}d\cdot(\log n+\log d)$. So far, we conclude (29)
+To prove the dimension free bound for $\frac{1}{n}\sum_{i=1}^{n}\phi_{\tau}(\xi_{i})X_{i}$, we denote $Q_{i}=\frac{1}{n}\phi_{\tau}(\xi_{i})X_{i}$ and $Q=\sum_{i=1}^{n}Q_{i}$.
+Since $\xi_{i}$ are symmetric, $\mathbb{E}[\phi_{\tau}(\xi_{i})|X_{i}]=0$.
+By (37), we obtain $\mathbb{E}[\psi(\xi_{i})^{2}|X_{i}]\leq\sigma^{2}$, $|\phi_{\tau}(\xi_{i})|\leq\tau$.
+Therefore, the estimates we have obtained earlier for the four parameters still apply.
+By (28), with $\tau=\frac{\max(\sigma,a)}{\log^{2}d}\sqrt{\frac{n}{m}}$, we need to impose conditions to $n$ such that in (33)
+$\left(\frac{\log^{2}d}{n}\cdot\frac{\max(\sigma,a)}{\log^{2}d}\sqrt{\frac{n}{m}}\right)^{1/3}\cdot\left(\frac{\sigma^{2}}{nm}\right)^{1/3}\leq\max(\sigma,a)\sqrt{\dfrac{1}{nm}},$
+$\dfrac{\log d}{n}\frac{\max(\sigma,a)}{\log^{2}d}\sqrt{\frac{n}{m}}\leq\max(\sigma,a)\sqrt{\dfrac{1}{nm}}.$
+Given $n\geq m\log^{4}d$, the upper bound of $\|\frac{1}{n}\sum_{i=1}^{n}\phi_{\tau}(\xi_{i})X_{i}\|$ is free of the $\log d$ factor. We have proved (30).
+Similarly, let $Q=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}$.
+Since $\epsilon_{i}\leq 1$, we have $R(Q)=\frac{1}{n}$.
+We have already estimated the four parameters above.
+By (28), it remains to ensure that
+$\left(\frac{\log^{2}d}{n}\right)^{1/3}\cdot\left(\frac{1}{nm}\right)^{1/3}\leq C\sqrt{\dfrac{1}{nm}},$
+$\dfrac{\log d}{n}\leq C\sqrt{\dfrac{1}{nm}},$
+which leads to the condition $n\geq m\log^{4}d.$ This completes the proof of (31).
+Hereafter, we use the notation $Q_{i}=\frac{1}{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}$ and $Q=\sum_{i=1}^{n}Q_{i}$.
+1.
+$\gamma(Q)\leq\sqrt{\frac{L_{2}}{nm}}$.
+Note $Q_{i}Q_{i}^{T}=\frac{1}{n^{2}}\textbf{1}_{\chi_{i}}X_{i}X_{i}^{T}$.
+$\mathbb{E}[\frac{1}{n^{2}}\textbf{1}_{\chi_{i}}X_{i}X_{i}^{T}]$ is a diagonal matrix with diagonal elements $\sum_{k=1}^{m_{2}}\mathbb{P}(\chi_{i}|\mathcal{E}_{i}(j,k))P_{jk}$.
+Thus, we have $\|Q_{i}Q_{i}^{T}\|\leq\frac{L_{2}}{n^{2}m}$ and $\|QQ^{T}\|\leq\frac{L_{2}}{nm}$.
+Similarly, we can obtain $\|QQ^{T}\|\leq\frac{L_{2}}{nm}$ and therefore $\gamma(Q)\leq\sqrt{\frac{L_{2}}{nm}}$.
+2. $\gamma_{*}(Q)\leq\sqrt{\frac{L_{3}}{nm\log^{3}d}}$. For $y\in\mathbb{R}^{m_{1}},z\in\mathbb{R}^{m_{2}}$ with $\|y\|_{2}=\|z\|_{2}=1$, we have the following result:
+$\mathbb{E}(y^{T}\bar{\xi}_{i}X_{i}z)^{2}$ $=\sum_{j,k}P_{jk}\mathbb{P}(\textbf{1}_{\chi_{i}}|\mathcal{E}_{i}(j,k))y^{2}_{j}z^{2}_{k}$
+$\leq\sum_{j,k}P_{jk}y^{2}_{j}z^{2}_{k}$
+$\leq\dfrac{L_{3}}{m\log^{3}d}.$
+where we use $\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}y^{2}_{j}z^{2}_{k}=\|y\|_{2}^{2}\cdot\|z\|_{2}^{2}=1$ and $\max P_{jk}\leq\frac{L_{3}}{m\log^{3}d}$.
+Since $Q_{i}$ are independent and mean-zero, $\mathbb{E}(y^{T}Qz)^{2}=\dfrac{1}{n^{2}}\sum_{i=1}^{n}\mathbb{E}\left(y^{T}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}z\right)^{2}\leq\frac{L_{3}}{nm\log^{3}d}$.
+Thus $\gamma_{*}(Q)\leq\sqrt{\frac{L_{3}\sigma^{2}}{nm\log^{3}d}}$.
+3.
+$g(Q)\leq\sqrt{\frac{L_{3}}{nm\log^{3}d}}$.
+Since $Q_{i}$ are independent, $\operatorname{Cov}(Q)=\dfrac{1}{n^{2}}\sum_{i=1}^{n}\operatorname{Cov}(\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}})$.
+Define $\textbf{1}_{(j,k)}=\textbf{1}(X_{i}=e_{j}(m_{1})e_{k}(m_{2})^{T})$ (for simplicity of notation, we do not include a subscript i) and denote $I(j,k)=\epsilon_{i}\textbf{1}_{(j,k)}\textbf{1}_{\chi_{i}}$.
+Note the following facts
+$\mathbb{E}[I(j,k)]=\mathbb{E}[\epsilon_{i}]\mathbb{E}[\textbf{1}_{(j,k)}\textbf{1}_{\chi_{i}}]=0,$
+$I(j,k)I(l,h)=0\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ (j,k)\neq(l,h).$
+Thus $\text{Cov}\big{(}I(j,k),I(l,h)\big{)}=\mathbb{E}[I(j,k)I(l,h)]-\mathbb{E}[I(j,k)]E[I(l,h)]=0$ for $(j,k)\neq(l,h)$. Then we obtain
+$\text{Cov}(\bar{\xi}_{i}X_{i})\Big{(}(j,k),(l,h)\Big{)}=\begin{cases}\mathbb{P}(\chi_{i}|\mathcal{E}_{i}(j,k))P_{jk}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ \leavevmode\nobreak\ (j,k)=(l,h),\\ 0\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ (j,k)\neq(l,h).\end{cases}$
+Since $P_{jk}\leq\frac{L_{3}}{m\log^{3}d}$, $\|\operatorname{Cov}(\bar{\xi}_{i}X_{1})\|\leq\frac{L_{3}}{m\log^{3}d}$. Thus, $g(Q)\leq\sqrt{\frac{L_{3}}{nm\log^{3}d}}$ holds.
+4.$R(Q)\leq\frac{2}{n}$. Obviously, $\|Q_{i}\|=|\textbf{1}_{\chi_{i}}|/n$.
+Plug the four estimates for $\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}$ in (28) and we conclude (32) given $n\geq m\log^{4}d$. ∎
+4.6 Concentration of Maximum of Independent Random Variables
+Given $\|X\|_{\psi_{\alpha}}\leq\sigma$, it holds for $t>0$
+$\mathbb{P}(|X|\geq\sigma t)\leq 2\exp(-t^{\alpha}).$
+For a metric space $(T,\rho)$, a sequence of its subsets $\{T_{m}\}_{m\geq 0}$ is called an admissible sequence if $|T_{0}|=1$ and $|T_{m}|\leq 2^{2^{m}}$.
+For $0<\alpha<\infty$, the functional $\gamma_{\alpha}(T,d)$ is defined by
+$\gamma_{\alpha}(T,\rho)=\inf_{\{T_{n}\}^{\infty}_{n=1}}\sup_{t\in T}\sum_{n=0}^{\infty}2^{n/\alpha}\rho(t,T_{n})$
+where the infimum is taken over all possible admissible sequences. If a random process $\{X_{t},t\in T\}$ satisfies
+$\mathbb{P}(|X_{t}-X_{s}|\geq t\rho(t,s))\leq 2\exp(-t^{\alpha}),$
+the following theorem holds.
+Theorem 15 (Theorem 3.2 in Dirksen [2015]).
+There exist constant $C_{1,\alpha},C_{2,\alpha}$ such that
+$\mathbb{P}\left(\sup_{t\in T}|X_{t}-X_{t_{0}}|\geq e^{1/\alpha}\left(C_{1,\alpha}\gamma_{\alpha}(T,\rho)+tC_{2,\alpha}\Delta_{\rho}(T)\right)\right)\leq\exp\left(-t^{\alpha}/\alpha\right)$ (34)
+where $\Delta_{\rho}(T)=\sup_{t,s\in T}\rho(t,s)$ and $t\geq 0$.
+Let $\{\xi_{i}\}_{i=1}^{n}$ be a sequence of independent $\psi_{\alpha}$ random variables with $\|\xi_{i}\|_{\psi_{\alpha}}\leq\sigma$.
+By Lemma A.3 in Götze et al.
+[2021], we have the following triangle inequality,
+$\|X+Y\|_{\psi_{\alpha}}\leq K_{\alpha}(\|X\|_{\psi_{\alpha}}+\|Y\|_{\psi_{\alpha}}),$
+where $K_{\alpha}=2^{1/\alpha}$ for $\alpha\in(0,1)$ and $K_{\alpha}=1$ for $\alpha\geq 1$. We define the trivial metric on $T=\{1,2,\cdots,n\}$,
+$\rho(s,t)=\begin{cases}2K_{\alpha}\sigma,t\neq s,\\ 0,t=s.\end{cases}$
+With this metric, we construct an admissible sequence $\{T_{n}\}$.
+Note that there exists $\tilde{m}$ such that $2^{2^{\tilde{m}}}\leq n\leq 2^{2^{\tilde{m}+1}}$.
+For $m\leq\tilde{m}+1$, take $T_{m}=\{1\}$.
+For $m>\tilde{m}+1$, take $T_{m}=T$.
+Then, we obtain
+$\sup_{t\in T}\sum_{m=0}^{\infty}\rho(t,T_{n})=\sup_{t\in T}\sum_{m=0}^{\tilde{m}+1}2^{m/\alpha}\rho(t,T_{m})=2K_{\alpha}\sigma\sum_{m=0}^{\tilde{m}+1}2^{m/\alpha}=\sigma C_{\alpha}2^{\tilde{m}/\alpha},$
+where $C_{\alpha}$ is a constant only dependent on $\alpha$. Noting that $2^{\tilde{m}}\apprle\log n$, we obtain
+$\gamma_{\alpha}(T,\rho)\leq C_{\alpha}\sigma(\log n)^{1/\alpha}.$
 Note
-$\|\bar{A}\|_{1}$ $=\|P_{A_{0}}(\bar{A}-A_{0})+P_{A_{0}}^{\perp}(\bar{A}-A_{0})+P_{A_{0}}(A_{0})\|_{1}$
-$\geq\|P_{A_{0}}^{\perp}(\bar{A}-A_{0})\|_{1}+\|P_{A_{0}}(A_{0})\|_{1}-\|P_{A_{0}}(\bar{A}-A_{0})\|_{1}.$
-Then, we get
-$0\leq\frac{4}{3}\|P_{A_{0}}(\bar{A}-A_{0})\|_{1}-\frac{2}{3}\|P_{A_{0}}^{\perp}(\bar{A}-A_{0})\|_{1}.$
-For (2), $L(A)=\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}$ and $\nabla L(A_{0})=\frac{2}{n}\sum_{i=1}^{n}\xi_{i}X_{i}$.
-For (8), $L(A)=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}}$ and $\nabla L(A_{0})=\frac{\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}}{L(A_{0})}=\frac{\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}}{\sqrt{\frac{1}{n}\sum_{i=1}^{n}\xi_{i}^{2}}}$.
-For (10), $L(A)=\frac{1}{n}\sum_{i=1}^{n}l_{\tau}(Y_{i}-\langle X_{i},A\rangle)$, and $\nabla L(A_{0})=\frac{1}{n}\sum_{i=1}^{n}\psi_{\tau}(\xi_{i})X_{i}$.
-∎
+$\sup_{i\in\{1,\cdots,n\}}|\xi_{i}|\leq|\xi_{1}|+\sup_{i\in\{1,\cdots,n\}}|\xi_{i}-\xi_{1}|.$
+By (34) we get
+$\mathbb{P}\left(\sup_{i\in\{1,\cdots,n\}}|\xi_{i}|\geq e^{1/\alpha}\left(C_{1,\alpha}\sigma(\log n)^{1/\alpha}+(C_{2,\alpha}+1)\sigma t\right)\right)$
+$\leq$ $\mathbb{P}\left(\sup_{i\in\{1,\cdots,n\}}|\xi_{i}-\xi_{1}|\geq e^{1/\alpha}\left(C_{1,\alpha}\sigma(\log n)^{1/\alpha}+C_{2,\alpha}\sigma t\right)\right)+\mathbb{P}\left(|\xi_{1}|\geq e^{1/\alpha}\sigma t\right)$
+$\leq$ $\exp(-t^{\alpha}/\alpha)+\frac{2}{e}\exp(-t^{\alpha}/\alpha)\leq 2\exp(-t^{\alpha}/\alpha),$
+where we use $\mathbb{P}(X+Y\geq a+b)\leq\mathbb{P}(X\geq a)+\mathbb{P}(Y\geq b)$. In a cleaner form, we present the concentration inequality
+$\mathbb{P}\left(\sup_{i\in\{1,\cdots,n\}}|\xi_{i}|\geq C_{1,\alpha}\sigma(\log n)^{1/\alpha}+C_{2,\alpha}\sigma t^{1/\alpha}\right)\leq 2\exp(-t)$ (35)
+As a consequence, if $\{\xi_{i}\}_{i=1}^{n}$ is a sequence of sub-exponential random variables satisfying $\|\xi_{i}\|_{\psi_{2}}\leq\sigma$, there exist constants $C_{1},C_{2}$ such that
+$\mathbb{P}\left(\sup_{i\in\{1,\ldots,n\}}|\xi_{i}|\geq C_{1}\sigma\sqrt{\log n}+C_{2}\sigma\sqrt{t}\right)\leq 2\exp(-t).$ (36)
+Lemma 16.
+Let $Y$ be a random variable with mean 0 and $\mathbb{E}[Y^{2}]\leq\sigma^{2}$. For any $y$,
+$\operatorname{Var}(\phi_{y}(Y))\leq\sigma^{2}.$ (37)
+Proof.
+Let $Y^{\prime}$ be an independent copy of $Y$. Since the function $\psi_{y}(\cdot)$ is Lipchitz with Lipchitz constant $1$,
+$2\operatorname{Var}(\psi_{y}(Y))\leq\mathbb{E}(\psi_{y}(Y)-\psi_{y}(Y^{\prime}))^{2}\leq\mathbb{E}(Y^{\prime}-Y)^{2}=2\mathbb{E}[Y^{2}].$
+Then, we conclude the lemma. ∎